### v1

# Deep Tree-based Retrieval for Efficient Recommendation: Theory and Method

Abstract
With the development of deep learning techniques, deep recommendation models also achieve remarkable improvements in terms of recommendation accuracy. However, due to the large number of candidate items in practice and the high cost of preference computation, these methods also suffer from low efficiency of recommendation. The recently proposed tree-based deep recommendation models alleviate the problem by directly learning tree structure and representations under the guidance of recommendation objectives. However, such models have shortcomings. The max-heap assumption in the hierarchical tree, in which the preference for a parent node should be the maximum between the preferences for its children, is difficult to satisfy in their binary classification objectives. To this end, we propose Tree-based Deep Retrieval (TDR for short) for efficient recommendation. In TDR, all the trees generated during the training process are retained to form the forest. When learning the node representation of each tree, we have to satisfy the max-heap assumption as much as possible and mimic beam search behavior over the tree in the training stage. This is achieved by TDR to regard the training task as multi-classification over tree nodes at the same level. However, the number of tree nodes grows exponentially with levels, making us train the preference model with the guidance of the sampled-softmax technique. The experiments are conducted on real-world datasets, validating the effectiveness of the proposed preference model learning method and tree learning method.
Index Terms:
Recommender system; Tree-based Index; Multi-classification; Sampled Softmax; Efficient Recommendation
1 Introduction
Due to the excessive quantity of daily information, people severely suffer from information overload in the information era. The recommendation is an important means to address information overload by providing a personalized ranking list on a set of information. With the development of deep learning techniques, recommendation techniques also achieve remarkable development and improvements in ranking performance. The widespread application of these techniques has created great economic benefits for various kinds of content providers in industrial companies.
Through the use of deep learning, we not only learn better representations for users, items, and contexts but also provides a more generalized expression for users’ preference scores via neural networks than the widely-used inner product in matrix factorization. Both lead to stronger recommendation performance, as shown in models like DIN [1], DIEN [2], NCF [3], CDL [4], and CKE [5]. However, the use of the neural preference function brings online-serving challenges for recommender systems due to the high cost of preference computation. Generally speaking, immediate responses to adaptive recommendations are prerequisites for excellent customer experiences and custom retention. Existing popular and well-performed graph-based indexes, e.g. HNSW [6], quantization-based indexes, e.g. PQ [7], AQ [8] and SCANN [9], and Hash-based indexes, e.g. SGDH [10] are usually built based on inner product or Euclidean distance, such that they are not suitable for speeding up the recommendation of deep models. In particular, neural networks based preference functions are not usually a valid metric, and difficult to be compatible with Euclidean distance and inner product. Therefore, two items being close in Euclidean or inner-product space may have divergent neural preference scores.
To guarantee the compatibility between search indexes with the neural preference functions, it should be a good solution to learn the search indexes together with the recommendation model under the guidance of recommendation objectives. The representative work is the tree-based deep model (i.e. TDM [11] and its improved version JTM [12]). These models use the balanced tree index, which is constructed by hierarchically clustering item representations from top to down. Given the tree, the top-k ranked items are obtained by layer-wise beam search, which selects the k-largest tree-nodes based on neural preference scores in each level from top to down. In this way, beam search achieves logarithmic computation complexity w.r.t. the number of items. To guarantee the accuracy of beam search, these tree-based deep models rely on the following max-heap assumption: the preference scores of query for a parent node should be the maximum between the preference scores of its children node. For the sake of satisfying the assumption, these tree-based models cast the overall problem as a binary classification problem by treating nodes in the path from the root to positive samples as positive and randomly sampled nodes as negative. However, these tree models suffer from the following two drawbacks. First, the max-heap assumption is not well satisfied by the used binary classification objectives due to the lack of horizontal competition among all tree nodes at the same level. Second, the learned index only includes a single tree, which is different from the widely-used multiple tree index in industrial libraries, such as Annoy. A single tree may easily misuse data around the boundary, which can be corrected by other trees of different structures. Therefore, this drawback also provides an opportunity to improve the accuracy of efficient recommendations and strike a better balance between efficiency and accuracy.
We proposed Deep Forest-based Recommender (TDR for short) for compositing multiple trees. To guarantee that a set of trees can bring improvements over a single tree, these trees should be diversified. In TDR, we keep all the trees and the corresponding preference models generated during the training process goes on to form the forest model. To satisfy the max-heap assumption as much as possible, TDR regards the training task as multi-classification over tree nodes at the same level, which enables horizontal competition among them and mimics beam search behavior in the training stage. However, the number of tree nodes in each level grows exponentially with levels, making the softmax loss suffer from computational challenges. To promote the training efficiency, we resort to sampled softmax for approximation, which only requires a small number of sampled nodes to use. Additionally, we propose a tree learning method that allow us to learn the preference model and tree alternately.
As an extension of the preliminary paper [13], which contributed to the hierarchical multi-classification training mode a, we further make the following contributions:
  • We find the issue of multi-classification training mode under beam search from the aspect of Bayes optimality and propose a rectification method.
  • We propose a negative sampling method, i.e., top-down sampling, for the multi-classification problem with the tree index. This sampling method is guided by sampled softmax theory, which can lead to more accurate estimations of loss gradient.
  • We provide a generalization analysis of the tree-based deep retrieval model, demonstrating the great generalization capability of our proposed methods.
  • We evaluate the proposed methods on four real-world datasets and validate the superiority of TDR to the baselines and the effectiveness of the tree learning method and sampling method.
2 Related work
This study aims to enhance recommendation efficiency through the construction of a hierarchical multi-classification tree, focusing on both the methodologies and theoretical foundations. We begin by reviewing recent advancements in efficient recommendation. Subsequently, we survey recent significant techniques that are highly beneficial in developing such recommender systems and then delve into the theories closely connected to hierarchical multi-classification within tree structures.
2.1 Efficient Recommendation
Efficient recommendation relies on building a search index, including LSH [14], inverted index [15, 16], tree index [17, 18] [^1] It is used in an industrial library ANNOY (https://github.com/spotify/annoy) and graph index [6] for all items. The recommender system usually uses the inner product for computing preference scores, and the top-k recommendation can be cast into the maximum inner product search (MIPS) problem. The search index is usually constructed based on Euclidean distance and has been extended to the inner product. This extension can be achieved by establishing the relationship between nearest neighbor search and MIPS [19, 20, 21], or learning from either item representations [22, 23, 24, 25] or the raw data directly [26, 27, 28, 29]. With the introduction of deep learning into the recommender systems, the preference score function becomes complicated, such that it is challenging to transform from neural ranking to NNS or MIPS. Existing work either directly used metric-based index[30], or learns search index from raw data directly together with recommendation models [11, 12, 31].
2.2 Methods for Developing Recommender Systems
2.2.1 Negative Sampling in RecSys
Negative sampling is an important method to address the negative missing and exposure bias problems and to speed up the convergence of recommender training  [32, 33, 34]. It includes static sampling, such as uniform sampling [35] and popularity sampling, and adaptive sampling. The adaptive sampling is context-dependent, whose representation work includes adaptive oversampling [35], rejection sampling [36, 37], clustering-based sampling [32], and dynamic negative sampling [38]. The core idea of these adaptive samplers is that items with larger preference scores should be sampled with higher probability.
2.2.2 Techniques of Speedup Softmax Computation
In natural language applications, it is very computationally expensive to represent an output distribution over the choice of a word based on the softmax function. To address the efficiency, many approximate algorithms were proposed. For example, hierarchical softmax [39] and lightRNN [40] decomposed the probabilities, and Contrastive Divergence [41] approximated the gradient-based on MCMC. As an alternative, negative sampling is also widely used in reducing the computational cost of training the models. The representative work includes Noise-Contrastive Estimation [42] with the unigram distribution as a sampler, Generative Adversarial Networks [43, 44] with a neural-network empowered sampler, Self-Contrast Estimator [45] by the model in the immediately preceding epoch as a sampler, self-adversarial negative sampling [46] and Kernel-based sampling [47] with the tree index.
2.3 Theoretical Work
2.3.1 Bayes Optimality in Multi-Classification and Hierarchical Classification
Bayes optimality aims for the learned classifier to achieve the highest possible performance on previously unseen data distributions. In the realm of multi-classification, Bayes optimality has been extensively studied[48, 49]. Top-k Bayesian optimality was introduced to address ambiguities in ground truth categories by focusing on the k most likely predicted categories[50, 51, 52]. Much research on Bayes optimality initially centered around flat models. In recent years, researchers have begun to extend it to hierarchical classification. [53] uses Bayes optimality to measure the performance of probability estimation using tree models in hierarchical classification. [54] examines the performance degradation when using beam search in hierarchical classification to recall relevant items with tree models and proposes a method to achieve Bayes optimality under beam search conditions. Our work takes significant inspiration from the insights on Bayes optimality under beam search provided in [54]. While [54] primarily focuses on binary classification, we extend this analysis to a multi-class classification context and provide a comprehensive theoretical account by leveraging the concept of Bregman divergence.
2.3.2 Data-dependent Generalization Bounds for Multi-Classification
Generalization error bounds are used to provide formal guarantees for the performance of the learned classifier. [55] provides data-dependent generalization bounds for multi-classification problems based on margins. [56] further developed these bounds for multi-classification with multiple kernels. To accommodate multi-classification with a large number of label classes, data-dependent generalization error bounds with a mild dependency on the number of classes have also been studied [57]. The aforementioned bounds primarily focus on flat structures. In the context of hierarchical structures, [58] proposed a multi-class, hierarchical data-dependent generalization error bound for kernel classifiers used in large-scale taxonomies, explaining several empirical results related to hierarchical structures. We extend this work by incorporating beam search and considering more complex neural networks to derive the generalization error bound for hierarchical multi-classification tree models.
3 Preliminaries
In this section, we first review the tree-based recommendation models, i.e., TDM and JTM. We then elaborate on the training process of the preference model and the updating process of the tree index in both TDM and JTM. Finally, we provide a formal description of the tree model and formulate the retrieval process using beam search
3.1 Tree-based index and the deep model
To tackle the top-k retrieval of the most preferred items, Zhu et al.  [11, 12] develop a max-heap-like tree index which is compatible with any advanced preference models, i.e. TDM and JTM. TDM and JTM utilize a variation of DIN [1] as the preference model. As for the tree-based index, the binary tree is chosen but any multi-way tree such that each item of corpus is mapped to a leaf node is compatible in fact. The rigorous max-heap tree satisfies the formulation that
$p^{j}(n|u)=\frac{\max_{n_{c}\in node\ n^{\prime}s\ children}p^{(j+1)}(n_{c}|u)}{\alpha^{j}}$ (1)
where $p^{j}(n|u)$ denotes user $u$’s real preference to node $n$ at level $j$ and $\alpha^{j}$ is a layer-specific normalization term to make the sum of preference scores in level $j$ equals to one. E.q. (1) illustrates that the user’s preference for a node is proportional to the largest one of its children’s preference scores.
Input: User $u$, tree, beam size $B$ and solution size $K$, the preference model $p$
Output: $K$ items
1:  Result set $A=\emptyset$, candidate set $Q=\{root\ node\ n_{1}\}$;
2:  while Q is not empty do
3:     Remove all the leaf nodes from $Q$ and add these leaf nodes into result set $A$ if Q contains leaf nodes;
4:     Compute preference $p(n|u)$ for each left node in set $Q$;
5:     Parents={the top-B nodes of Q according to $p(n|u)$};
6:     Q={children of node $n\ |\ n$in Parents};
7:  end while
8:  return The top-K items w.r.t. the top-K leaf nodes according to $p(n|u)$, $n\in A$.
Algorithm 1 Beam Search: Layer-wise Retrieval
Top-k retrieval in recommendation is to find the k leaf nodes with the largest preference scores in the tree. If the real preference $p^{j}(n|u)$ can be available for any node at each level $j$ for user $u$, we can find the top-k leaf nodes by only retrieving all the top-k nodes at each layer from top layer to bottom layer. This retrieval process guarantees to finding the top-k leaf nodes because the top-k nodes with the largest preference scores at level $j+1$ must belong to the children of top-k nodes with the largest preference scores at layer $j$. The max-heap tree index allows users can obtain the top-k recommendation from coarse clusters to fine clusters until each cluster only contains one leaf node. Furthermore, the time complexity of the retrieval process is logarithmic w.r.t. the corpus size. The layer-wise retrieval process is presented in Algorithm 1, where $p(n|u)$ is exactly user $u$’s preference to node $n$ and we omit the superscript (i.e. layer).
However, the real preference $p^{j}(n|u)$ is unavailable and various kinds of prediction models can be used to estimate the preference. The prediction model used in TDM and JTM is a variation of DIN [1]. The inputs of the prediction model are the history of user behaviors (i.e. the items that the user interacted with sorted by interaction time) and the node. The user profile (e.g. age, gender, height, and so on) can be used to enhance the model if the information is available. The user behaviors are partitioned into several time windows and zero padding is used to make up the time windows if some items are missing. Each time window outputs the weighted average embedding where the weight is calculated by an activation unit. In the activation unit, the concatenation (of the item embedding, the point-wise product between item embedding and node embedding, node embedding) is fed into the MLP and the output is the activation weight. The concatenation (embedding of user profile, each output of time windows, node embedding) is fed into MLP to generate the like and dislike probabilities in the preference model. PReLU is used as the activation function and batch normalization is applied to the MLP in the preference model.
3.2 Learn the preference model and the tree
In tree-based recommendation systems, we need to learn the preference model (i.e. the neural network model to measure the scores between users and nodes) and the tree (i.e. the bijective mapping relationships between items and the leaf nodes of the tree). We use $\pi(\cdot)$ to denote the bijective mapping between items and leaf nodes in the latter content, i.e. $\pi(item_{i})=n_{leaf}$ means that the leaf node $n_{leaf}$ represents the item $item_{i}$. The preference model and the tree can be learned alternately. Concretely, fix the bijective mapping $\pi(\cdot)$ when learning the preference model and fix the preference model when learning the bijective mapping $\pi(\cdot)$. Repeat the alternate learning process until both the two parts converge.
To learn the preference model, TDM and JTM regard the training task as a binary classification problem essentially. Supposing user $u$ has an interaction with an item that corresponds to a leaf node is $n_{leaf}$, $n_{leaf}$ and its ancestors are positive nodes with label 1 but all the left nodes are negative nodes with label 0. Then the training loss for user $u$ is
$\mathcal{L}(u,Y_{u}^{+},Y_{u}^{-})=$ $-\sum_{n\in Y_{u}^{+}}y_{u}(n)\log p(\hat{y}_{u}(n)=1|u)$ (2)
$-\sum_{n\in Y_{u}^{-}}(1-y_{u}(n))\log p(\hat{y}_{u}(n)=0|u).$
$Y_{u}^{+}$ and $Y_{u}^{-}$ denote the positive node set and negative node set for user $u$ respectively. $y_{u}(n)$ denotes the label of node $n$ for user $u$. $p(\hat{y}_{u}(n)=1|u)$ and $p(\hat{y}_{u}(n)=0|u)$ denote the like and dislike probabilities for user $u$ to node $n$, which are exactly corresponding to the two outputs of preference model. Using all negative nodes to train the model is unacceptable both in time-consuming and memory consumption. Both TDM and JTM draw a negative node set (denoted as $S_{u}^{-}$) uniformly from all the negative nodes. Then $\mathcal{L}(u,Y_{u}^{+},S_{u}^{-})$ is used as training loss.
Treating the training task as a binary classification problem leads to three shortcomings. 1). The max-heap assumption is not well satisfied due to the lack of horizontal competition among all tree nodes at the same level. 2). The retrieval is conducted by a layer-wise stream (e.g. Algorithm 1) but the training process doesn’t include layer-wise mode which can lead to a gap between training and prediction. 3). There are much more negative nodes than positive nodes which can make the negative nodes dominate the training loss. At the same time, the imbalance between negative and positive nodes leads to an extra computation burden to find the proper size of sampled negative nodes. To overcome these shortcomings, we regard the preference model learning task as a multi-class classification problem which will be elaborated in later sections.
Both TDM and JTM utilize hierarchical clustering to update the tree structure. TDM clusters all the items recursively by k-means using the item embedding until each cluster only contains one item. The recursive clustering process can form a tree structure and the item in each final cluster is exactly mapped to a leaf node. JTM assigns all the items to each node from the top layer to the bottom layer such that the sum of the log-likelihoods for all users on each layer can be maximized. Interested readers can refer to the content of TDM  [11] and JTM  [12] for more details about how to learn the trees respectively. In our work, we propose a new tree learning method that is compatible with our proposed preference model learning mode.
3.3 Formalization of tree model and retrieval process
Consider a user space $\mathcal{U}$ and an item set $\mathcal{Y}$. It is assumed that the data is generated i.i.d. from some distribution $\mathbb{P}$ over $\mathcal{U}\times\mathcal{Y}$. An instance in this context is represented as $(u,y)\in\mathcal{U}\times\mathcal{Y}$, with the corresponding joint probability denoted as $p(u,y)$. Given a specific user $u$, the conditional probability of observing item $y$ is denoted by $\eta_{y}(u)=p(y\mid u)$. Since we regard the recommendation task as a multi-class classification problem, a common practice in this domain [59], the conditional probabilities $\eta(u)$ for a given user $u$ satisfy the normalization condition $\sum_{y\in\mathcal{Y}}\eta_{y}(u)=1$.
The tree model comprises a $b$-ary tree $\mathcal{T}$ and a node preference function $f$. The tree $\mathcal{T}$ consists of multiple layers of nodes, with nodes at the $j$-th level denoted by $\mathcal{N}^{j}$. The entire set of nodes can be expressed as $\mathcal{N}=\bigcup_{j=0}^{H}\mathcal{N}^{j}$, where $H=\log_{b}|\mathcal{N}|$ is the height of the tree. $\mathcal{N}^{0}$ contains only the root node, while $\mathcal{N}^{H}$ represents the set of leaf nodes. The number of nodes in the $j$-th layer is denoted as $c(j)=\left|\mathcal{N}^{j}\right|$. For each level $j$, we assume the nodes are listed in a fixed order, denoted as $(n_{1}^{j},n_{2}^{j},\dots,n_{c(j)}^{j})$. We introduce the function $\ell$ to return the layer of a node and the function $\delta$ to return the index of a node within its layer, i.e., $\ell(n_{i}^{j})=j$ and $\delta(n_{i}^{j})=i$.
For each node $n\in\mathcal{N}$, let $\rho^{j}(n)\in\mathcal{N}^{j}$ denote its ancestor in the $j$-th layer $(0\leq j\leq\ell(n))$. Specifically, $\rho^{\ell(n)}(n)$ is the node $n$ itself. The set $\mathcal{D}(n)\subseteq\mathcal{N}$ represents $n$’s child nodes, while $\mathfrak{L}(n)\subseteq\mathcal{N}^{H}$ represents the leaf nodes in the subtree rooted at $n$. To connect the tree $\mathcal{T}$ with the item set $\mathcal{Y}$, the mapping $\pi:\mathcal{N}^{H}\rightarrow\mathcal{Y}$ is introduced. $\pi$ maps each leaf of the tree to a unique item, with the inverse mapping denoted by $\varphi:\mathcal{Y}\rightarrow\mathcal{N}^{H}$. An item $y$ is mapped to a leaf and corresponds to a path in the tree from the root to the leaf, the $j$-th node in the path is $\rho^{j}(\varphi(y))$ and its index in this layer is denoted as $\delta^{j}(y)=\delta(\rho^{j}(\varphi(y)))$.
The node preference model $f:\mathcal{U}\times\mathcal{N}\rightarrow\mathbb{R}$ outputs preference scores for tree nodes, which are used to retrieve the top-$k$ nodes at each layer. The preference score for the $i$-th node in the $j$-th layer, given a user $u$, is denoted as $o_{i}^{j}(u)=f(u,n_{i}^{j})$. For simplicity, we will omit the user $u$ and use the notation $o_{i}^{j}$ for the remainder of this paper.
During retrieval, a layer-wise beam search with beam size $k$ is conducted along the tree $\mathcal{T}$. This process is guided by the preference between the user $u$ and the node $n$, and can be formulated as follows:
$\mathcal{B}^{j}(u)\in\operatornamewithlimits{argTopk}\limits_{i\in\widetilde{\mathcal{B}}^{j}(u)}\ o_{i}^{j},$ (3)
where $\mathcal{B}^{j}(u)$ denotes the set of indices of selected nodes in the $j$-th layer for the user $u$, and ~$\widetilde{\mathcal{B}}^{j}(u)=\left\{\delta(n)\mid n\in\bigcup_{i\in\mathcal{B}^{j-1}(u)}\mathcal{D}(n^{j-1}_{i})\right\}$ represents the index set of nodes to be expanded in the $j$-th layer. It’s important to note that the equation employs the symbol $\in$” instead of “$=$” because ties in probabilities may occur, leading to multiple potential sets of top-$k$ nodes. If |~$\left|\widetilde{\mathcal{B}}^{j}(u)\right|<k$, then all nodes in $j$-th layer will be selected. By recursively applying E.q. (3) until the $H$-th layer, the top-$k$ leaf nodes $\mathcal{B}^{H}(u)$ is retrieved, and the final retrieved item set for the user $u$ is
$\widehat{\mathcal{Y}}(u)=\{\pi(n_{i}^{H})\mid i\in\mathcal{B}^{H}(u)\}.$ (4)
4 Multi-class classification training mode
To address the shortcomings of the training mode in TDM and JTM, we develop a layer-wise training mode in this section. We regard the training task as a multi-classification problem at each layer, employing the classic multi-class cross-entropy loss to optimize the tree model. Upon examining the Bayes optimality of this training scheme, we found that the traditional multi-class cross-entropy loss leads to suboptimality under beam search. To mitigate this issue, we introduce a rectification method aimed at enhancing the retrieval performance of the tree model under beam search.
4.1 Multi-class cross-entropy loss
TDM and JTM regard the training task as a binary classification problem and binary cross-entropy is used as the training loss. In this way, each node contributes to the training loss individually. However, the layer-wise retrieval is conducted from top layer to bottom layer and we need to compare the candidates with each other node at the same layer. There is a gap between training and prediction in TDM and JTM. We propose the layer-wise training mode which regards the training task as a multi-class classification problem.
Concretely speaking, at layer $j$, which consists of $c(j)$ nodes, each node corresponds to a class, and multi-class classification with $c(j)$ classes is performed at that layer. For any user $u\in\mathcal{U}$, we attach a label $z_{i}^{j}(u)$ to each node $n_{i}^{j}$. For simplicity, we will omit $u$ and use the notation $z_{i}^{j}$ in the rest of the paper. Given item $y$, the positive node can be identified by backtracking from the item’s mapping leaf node to the root node, as demonstrated in TMD and JTM, and the corresponding label will be assigned as $1$, i.e., $z^{j}_{\delta^{j}(y)}=1$. The remaining $c(j)-1$ nodes are the negative nodes with the label 0 w.r.t. to the positive node. We made a small modification to the original DIN model so that the final layer consists of a single neuron, which outputs a value representing the preference score between the user and the node. Thus, for the $j$-th layer, the model outputs $c(j)$ preference scores, which are precisely suited for performing multi-class classification with $c(j)$ classes.
For user $u$, $o_{i}^{j}$ represents user $u$’s preference to the $i$-th node at layer $j$ and $p_{i}^{j}$ denotes the probability that the $i$-th node is positive node at layer $j$, where
$p_{i}^{j}=\frac{\exp{o_{i}^{j}}}{\sum_{k=1}^{c(j)}\exp{o_{k}^{j}}}.$ (5)
Given instance $(u,y)$, the multi-class cross entropy is used as the training loss at layer $j$, i.e.,
$\mathcal{L}_{j}(u,y)=-\log p^{j}_{\delta^{j}(y)}=-\log\frac{\exp o_{\delta^{j}(y)}^{j}}{\sum_{k=1}^{c(j)}\exp o_{k}^{j}}$ (6)
From E.q. (5) and E.q. (6), we can know that all the nodes contribute to the training loss simultaneously at layer $j$ which may relieve the gap between training and prediction. The training loss w.r.t. the whole tree can be written as
$\mathcal{L}(u,y)=\sum_{j=1}^{H}\mathcal{L}_{j}(u,y).$ (7)
It is obvious that we no longer suffer from the imbalance of positive node size and negative node size that happens in binary-class classification problem.
4.2 Suboptimality of multi-class cross-entropy loss under beam search
Treating the training task as a multi-class classification problem at each layer can alleviate some of the limitations associated with binary classification. However, the use of multi-class cross-entropy loss still leads to suboptimality under beam search. In this subsection, we will demonstrate how multi-class cross-entropy loss results in suboptimal outcomes when combined with beam search.
Similar to Bayes optimality defined in [52, 54], we define retrieval optimality under beam search to suit our hierarchical multi-class setting:
Definition 1 (Top-$k$ Retrieval Bayes Optimal).
Given the beam size $k$, the tree model consist of tree $\mathcal{T}$ and node preference function $f$ is top-$k$ retrieval Bayes optimal, if for any $u\in\mathcal{U}$ the following equation holds
$\widehat{\mathcal{Y}}(u)\in\operatornamewithlimits{argTopk}\limits_{y\in\mathcal{Y}}\eta_{y}(u),$ (8)
The underlying idea of Definition 8 is straightforward: for each user $u$, a top-$k$ retrieval Bayes optimal tree model should return the $k$ items with the highest conditional probabilities.
Given the user $u$, the conditional risk of multi-class cross-entropy loss in layer $j$ is
$\mathcal{R}_{j}(u)=\mathbb{E}_{y\sim\eta(u)}\left[\mathcal{L}_{j}(u,y)\right]=\mathbb{E}_{y\sim\eta(u)}\left[-\log\frac{\exp o_{\delta^{j}(y)}^{j}}{\sum_{k=1}^{c(j)}\exp o_{k}^{j}}\right]$ (9)
As previously mentioned, a label $z_{i}^{j}$ is attached for the node $n_{i}^{j}$. In fact, $z_{i}^{j}$ can be viewed as a conditional random variable and the $j$-th layer’s variables composite the random vector $Z^{j}$. Let $o^{j}$ be the preference vector for the $j$-th layer, whose $i$-th component is $o^{j}_{i}$. Previous literature [52] has proven that multi-class cross-entropy loss is rank-consistent, i.e., optimizing the multi-class cross-entropy loss (i.e. E.q. (9)) would lead to each element in $o^{j}$ matches the rank of the corresponding element in $\mathbb{E}\left[Z^{j}\right]$. We state their theorem and make some modifications to be compatible with our problem.
Building on the aforementioned definition, we will demonstrate that optimizing a tree model using multi-class cross-entropy loss can result in suboptimal outcomes. The demonstration leverages the concept of Bregman divergence and the fact the multi-class cross-entropy loss is a Bregman divergence. Let $M$ denote the number of classes, for a strictly convex and differentiable function $\psi:\mathbb{R}^{M}\to\mathbb{R}$, the Bregman divergence $D_{\psi}:\mathbb{R}^{M}\times\mathbb{R}^{M}\to\mathbb{R}$ induced by $\psi$ is defined as follows:
$D_{\psi}(z,o)=\psi(z)-\psi(o)-\nabla\psi(o)^{T}(z-o).$ (10)
By taking $\psi(s)=\sum_{m=1}^{M}s_{m}\log s_{m}$ and $g(o)_{m}=\exp(o_{m})/\sum_{k=1}^{M}\exp(o_{k})$, the multi-class cross-entropy loss
$\mathcal{L}(e_{i},o)=-\log\left(\frac{\exp(o_{i})}{\sum_{m=1}^{M}\exp(o_{m})}\right)$ (11)
can be expressed as the form of Bregman divergence:
$\mathcal{L}(e_{i},o)=D_{\psi}(e_{i},g(o))$ (12)
Here, $z,o,e_{i}\in\mathbb{R}^{M}$, and the vector $e_{i}$ is a one-hot vector where the $i$-th component is $1$ and other components are 00.
To make the demonstration more concise and clear, we define the concept of strictly rank preserving as follows:
Definition 2 (Rank Consistent).
Given $x,y\in\mathbb{R}^{M}$, we say that $y$ is rank consistent w.r.t. $x$, denoted $R(y,x)$, if for all $i,j\in\left[M\right]$,
$(x_{i}>x_{j}\Rightarrow y_{i}>y_{j})\wedge(x_{i}=x_{j}\Rightarrow y_{i}=y_{j})$
With the fact that multi-class cross-entropy loss is a Bregman divergence and the softmax function $g$ satisfies $R(s,g(s))$ for any $s\in\mathbb{R}^{M}$, the following theorem shows that optimizing the multi-class cross-entropy loss for layer $j$ will lead to $R(o^{j},\mathbb{E}[Z^{j}])$:
Theorem 1 (Theorem 3.1 of [52]).
Let $Z$ be an arbitrary random variable taking values in $\mathbb{R}^{M}$ for which both $\mathbb{E}[Z]$ and $\mathbb{E}[\psi(Z)]$ are finite. Given convex, differentiable function $\psi:\mathbb{R}^{M}\mapsto\mathbb{R}$, if continuous function $g:\mathbb{R}^{M}\mapsto\mathbb{R}^{M}$ satisfies that $R(s,g(s))$ holds for $\forall s\in\operatorname{domain}(g)$, and $\mathbb{E}[Z]\subseteq\operatorname{range}(g)$, then
$\operatornamewithlimits{argmin}\limits_{s\in\mathbb{R}^{M}}\mathbb{E}_{Z}\left[D_{\psi}(Z,g(s)\right]\subseteq\left\{s\in\mathbb{R}^{M}\mid R(s,\mathbb{E}[Z])\right\}$
In the multi-class training mode, for a given instance $(u,y)$, we only set the label $z_{\delta^{j}(y)}^{j}=1$ and other labels are set to 00 in the $j$-th layer. Then the expectation of $Z^{j}$ is as follows:
$\mathbb{E}[Z^{j}]=\left(\sum\limits_{n\in\mathfrak{L}(n_{1}^{j})}\eta_{\pi(n)},\sum\limits_{n\in\mathfrak{L}(n_{2}^{j})}\eta_{\pi(n)},\ \cdots,\sum\limits_{n\in\mathfrak{L}\left(n_{c(j)}^{j}\right)}\eta_{\pi(n)}\right).$ (13)
Notice that optimizing the multi-class cross-entropy loss (i.e. E.q. (9)) will lead to $R(o^{j},\mathbb{E}[Z^{j}])$ from layer $1$ to $H$, we can find it is sub-optimal under beam search.
Proposition 1.
The multi-class cross-entropy loss will make the tree model not top-$k$ retrieval Bayes optimal.
Proof.
Since optimizing the multi-class cross-entropy loss leads to $R(o^{j},\mathbb{E}[Z^{j}])$ for $1\leq j\leq H$, when performing the beam search along the tree, the result of $\operatornamewithlimits{argTopk}\limits_{i\in[c(j)]}o^{j}_{i}$ is identical to the result of $\operatornamewithlimits{argTopk}\limits_{i\in[c(j)]}\mathbb{E}[z^{j}_{i}]$.
[…]
4.3 Rectification under beam search
Through the above analysis of how multi-class cross-entropy loss leads to suboptimal outcomes under beam search, we can identify that the root cause is the incorrect assignment of labels to non-leaf nodes — labels of all ancestors of the leaf node $\varphi(y)$ are directly set to $1$ for any instance $(u,y)$. To mitigate this issue, the assignment of ancestors’ labels needs to be rectified. Fix user $u$, only when the given item $y$ is the one with the largest conditional probability among the items which are mapped from leaves $\mathfrak{L}(n_{i}^{j})$, can label $z_{i}^{j}$ of node $n_{i}^{j}$ can be assigned as $1$, which will make the tree close to the max-heap from the aspect of expectation. Denote the rectified label as $\bar{z}^{j}_{i}$, the formalization of assignment is as follows:
$\bar{z}_{i}^{j}(u)=\begin{cases}1,&n_{i}^{j}=\rho^{j}\left(\varphi(y)\right)\wedge y=\operatornamewithlimits{argmax}\limits_{y^{\prime}\in\pi\left(\mathfrak{L}(n_{i}^{j})\right)}\eta_{y^{\prime}}\\ 0,&\text{else}\end{cases}$
With such a rectified assignment of the label of nodes, for each node $n_{i}^{j}$, we can calculate its expectation of the corresponding random variable:
$\mathbb{E}[\bar{z}_{i}^{j}(u)]=\sum_{y\in\mathcal{Y}}\bar{z}_{i}^{j}(u)*\eta_{y}(u)=\max\limits_{n\in\mathfrak{L}(n_{i}^{j})}\eta_{\pi(n)},$
so the expectation of $j$-th layer’s random vector is
$\mathbb{E}[\bar{Z}^{j}]=\left(\max\limits_{n\in\mathfrak{L}(n_{1}^{j})}\eta_{\pi(n)},\max\limits_{n\in\mathfrak{L}(n_{2}^{j})}\eta_{\pi(n)},\dots,\max\limits_{n\in\mathfrak{L}(n_{c(j)}^{j})}\eta_{\pi(n)}\right).$
Notice that currently each layer’s expectation isn’t normalized, so a normalized term $\alpha^{j}(u)=\sum_{i=1}^{c(j)}\tilde{z}_{i}^{j}(u)$ should be attached for the unnormalized rectified label, and we obtain the assignment of the label:
Definition 3 (Label Rectification).
Fix user $u$, given an item $y$, the normalized rectified label $\tilde{z}_{i}^{j}(u)$ of node $n_{i}^{j}$ is defined as follows:
$\tilde{z}_{i}^{j}(u)=\begin{cases}\frac{1}{\alpha^{j}(u)},&n_{i}^{j}=\rho^{j}\left(\varphi(y)\right)\wedge y=\operatornamewithlimits{argmax}\limits_{y^{\prime}\in\varphi\left(\mathfrak{L}(n_{i}^{j})\right)}\eta_{y^{\prime}}\\ 0,&\text{else}\end{cases}$
We have shown that optimizing the multi-class cross-entropy loss will make each element of the preference vector match the rank of the corresponding element of the expectation vector. However, the whole analysis is based on the fact that the multi-class cross-entropy loss is a Bregman divergence and each layer’s labels corresponds to a one-hot vector. A subsequent question is raised — if the label is rectified as Definition 3, can we still perform an analysis with the Bregman divergence? The answer is yes. All we need is to modify the original multi-class cross-entropy loss.
Proposition 2.
For any given $w\in[0,+\infty)$, construct a vector $e_{i}(w)=(0,\dots,0,w,0,\dots,0)$ where the $k$-th component is $i$ and all other components are 00. For $o\in\mathbb{R}^{M}$, the loss
~$\widetilde{\mathcal{L}}(e_{i}(w),o)=-w\log\left(\frac{\exp(o_{i})}{\sum_{m=1}^{M}\exp(o_{m})}\right)+w\log w+w-1$ (14)
can be expressed in the form of Bregman divergence.
Proof.
See Appendix. ∎
Compared to the original multi-class cross-entropy loss (E.q. (6)), the loss of E.q. 14 is with a slight modification that multiplies a weight and adds with a constant when given $w$, and is perfectly compatible with the proposed rectified label. Replace the original multi-class cross-entropy loss with the modified loss and set the label of nodes as Definition 3, given the instance $(u,y)$, for the $j$-th layer, now the following loss is optimized:
~$\widetilde{\mathcal{L}}_{j}(u,y)=\widetilde{L}(e_{\delta^{j}(y)}(\tilde{z}^{j}_{\delta^{j}(y)}),o^{j})$
$=-\tilde{z}^{j}_{\delta^{j}(y)}\log\left(\frac{\exp(o^{j}_{\delta^{j}(y)})}{\sum_{k=1}^{c(j)}\exp(o_{k})}\right)+\tilde{z}^{j}_{\delta^{j}(y)}\log\tilde{z}^{j}_{\delta^{j}(y)}+\tilde{z}^{j}_{\delta^{j}(y)}-1$
and corresponding conditional risk for user $u$ of $j$-th layer is
~$\widetilde{\mathcal{R}}_{j}(u)=\mathbb{E}_{y\sim\eta_{y}(u)}\left[\widetilde{\mathcal{L}}^{j}(u,y)\right].$
As indicated in Proposition 2, the modified loss $\widetilde{\mathcal{L}}$ still owns the property of Bregman divergence. Notice with the normalization, $\mathbb{E}[\tilde{Z}^{j}]\subseteq\operatornamewithlimits{range}(g)$ holds for $1\leq j\leq H$, by the Theorem 1, we have
$\operatornamewithlimits{argmin}\limits_{o^{j}\in\mathbb{R}^{c(j)}}\widetilde{\mathcal{R}}_{j}(u)\subseteq\left\{o^{j}\in\mathbb{R}^{c(j)}|R(o^{j},\mathbb{E}[\tilde{Z}^{j}])\right\}$
for any user $u\in\mathcal{U}$, which means optimizing the modified loss can lead the tree model to be top-$k$ retrieval Bayes optimal under beam search.
Lemma 1.
Given a user $u\in\mathcal{U}$, if the tree model, which consists of a tree $\mathcal{T}$ and a preference model $f$, satisfies $R(o^{j},\mathbb{E}[\tilde{Z}^{j}])$ for all $1\leq j\leq H$, then
$\mathcal{B}^{j}(u)\in\operatornamewithlimits{argTopk}\limits_{i\in\left[c(j)\right]}\mathbb{E}[\tilde{z}^{j}_{i}]\Longrightarrow\mathcal{B}^{j+1}(u)\in\operatornamewithlimits{argTopk}\limits_{i\in\left[c(j+1)\right]}\mathbb{E}[\tilde{z}^{j+1}_{i}]$
holds for any beam size $k$ within the range $11\leq j\leq H-$ 1.
Proposition 3.
For any user $u\in\mathcal{U}$, if the tree model, which consists of a tree $\mathcal{T}$ and a preference model $f$, satisfies that $R(o^{j},\mathbb{E}[\tilde{Z}^{j}])$ for all $1\leq j\leq H$, then the tree model is top-$k$ retrieval Bayes optimal for any beam size $k$.
Proof.
See Appendix. ∎
The above analysis has shown that optimizing the modified loss with the normalized rectified label can mitigate the issue of sub-optimality under beam search of original multi-class cross-entropy loss from a theoretical aspect. In practice, we optimize the loss by way of stochastic gradient descent, when given the $w$, the gradient of term $w\log w+w-1$ with respect to preference score $o$ is 00, so it can be ignored in the practical optimization and the loss is the original multi-class cross-entropy loss multiplied with a weight $w$. Furtherly, let’s look at the weight $w$. The $w$ is set as the normalized label $\tilde{z}$ where the role of normalization is to make the sum of expectation of random variable corresponding to the rectified label in the same layer be $1$ to satisfy $\mathbb{E}[\widetilde{Z}^{j}]\subseteq\operatorname{range}(g)$. But we have to point out that in practical optimization (i.e. stochastic gradient descent), the operation of dividing the rectified label by the expectation actually influences the learning rate. And it is easy to understand that the sum of expectations for different layers given an instance varies dramatically but the sum of expectations for the same layer of different instances are similar, but we aggregate the loss of different layers together for a given instance $(u,y)$, so the normalization will make the optimization difficult in practice due to the difficulty of setting a suitable learning rate. Therefore, we optimize the modified loss with the original rectified label $\bar{z}^{j}_{i}$ without normalization, which can be viewed as the adaptive learning rate adjustment, and the loss for $j$-th layer is
~$\widetilde{\mathcal{L}}_{j}(u,y)=-\bar{z}^{j}_{\delta^{j}(y)}\log\left(\frac{\exp o^{j}_{\delta^{j}(y)}}{\sum_{k=1}^{c(j)}\exp(o^{j}_{k})}\right),$ (15)
and the loss of the whole tree is
~$\widetilde{\mathcal{L}}(u,y)=\sum_{j=1}^{H}\widetilde{\mathcal{L}}_{j}(u,y).$ (16)
However, there remains an issue that cannot be ignored. All of our analysis is based on the conditional probability given user $u$, which is unknown in practice. A direct way is estimating the conditional probability using the preference model $f$. However, as the $f$ is a complicated neural network, the calculation for the estimation of conditional probability by $f$ is very time-consuming. What’s more, since the preference model is initialized randomly, the estimation from it is very inaccurate for a long time of the training, which will make the training more difficult to converge or even fall into local optimum finally. Considering the above shortcomings of estimating the conditional probability by the preference model $f$, we make use of a trained small model to perform the probability estimation in our experiment. The detailed process will be introduced in the experimental section.
5 Negative Sampling to Estimate Softmax Loss Gradient
5.1 Unbiased estimation of softmax loss gradient
For the training instance $(u,v)$, calculating its training loss (i.e., E.q. (7)) requires calculating the probability of the positive sample at each layer $j$. In practical application scenarios, where the size of the product set ranges from millions to billions or even tens of billions, calculating softmax probabilities on such a large scale is extremely inefficient. This necessitates some techniques to estimate the gradient of the loss function. Hierarchical softmax[39] employs a tree-building approach to compute softmax probabilities through hierarchical clustering instead of flat softmax calculations. However, in real tasks that require learning model parameters, estimating the gradient of the loss function is preferred over estimating the loss function itself. Moreover, constructing a parameterized tree based on Hierarchical softmax introduces a large number of additional parameters, exacerbating the complexity of the problem and reducing trainability. The loss based on NCE (Noise Contrastive Estimation) requires sampling a large number of negative samples to approximate the gradient of the original loss function. Sampled softmax is proposed to approximate full softmax during model training  [60, 61], which offers advantages in recommendation scenarios, such as alleviating popularity bias, discovering and utilizing difficult negative samples for gradient updates, and maximizing ranking metrics[62]. Therefore, sampled softmax is selected to alleviate the high computational complexity issue associated with softmax based loss functions in our proposed method.
Instead of calculating the training loss E.q. (6) over all classes, only the positive class and $m$ negative classes are considered where the $m$ negative classes are sampled from all the negative classes according to certain probability distribution $\bm{q}$ with replacement. In the rest of the paper, we use $s^{j}$ to denote the training samples at layer $j$ and use $q_{s^{j}_{i}}$ ($1\leq i\leq c(j)$) to denote the probability that the $i$-th term of $s^{j}$ can be sampled from the negative nodes of layer $j$. We suppose $s^{j}_{1}$ is always the positive class and the rest terms of $s^{j}$ are negative classes of layer $j$ without loss of generality. For example, $m=5$ and the sample set is $s^{j}=\{2,3,4,8,7,4\}$. The $s_{j}$ indicates that the $2$-rd node of layer $j$ is the positive node and the $4$-th node is sampled twice while other nodes (i.e. indexed at $3,8,7$) are sampled once each.
However, we don’t use the outputs w.r.t. sampled nodes of $s^{j}$ to approximate the loss directly. The slight adjustment is conducted for each output by
$\hat{o}_{s_{i}^{j}}^{j}=\left\{\begin{aligned} &o^{j}_{s^{j}_{i}}-\ln(mq^{j}_{s^{j}_{i}})\ \ \ \ if\ y_{s_{i}^{j}}=0\\ &o^{j}_{s^{j}_{i}}-\ln(1)\ \ \ \ \ \ \ \ \ if\ y_{s_{i}^{j}}=1\end{aligned}\right.$ (17)
We subtract the natural logarithm of the sampling probability times $m$ from their original matching scores for the sampled negative sample nodes, while the matching scores of the positive samples remain unchanged. This adjustment can guarantee the sampled softmax is unbiased when $m$ is infinite (i.e. $m\to\infty$) [61]. The training loss is calculated over the adjusted outputs and the original training loss E.q. (6) at layer $j$ can be adjusted to
^$\hat{\mathcal{L}}_{j}(u)=-\sum_{i=1}^{m+1}y_{s_{i}^{j}}^{j}\log\hat{p}_{i}^{j}=(\log\sum_{i=1}^{m+1}\exp(\hat{o}_{s_{i}^{j}}^{j}))-(\sum_{i=1}^{m+1}y_{s_{i}^{j}}^{j}\hat{o}_{s_{i}^{j}}^{j}).$ (18)
where
$\hat{p}_{i}^{j}=\frac{\exp{\hat{o}_{s_{i}^{j}}^{j}}}{\sum_{k=1}^{m+1}\exp{\hat{o}_{s_{k}^{j}}^{j}}}.$ (19)
The training loss w.r.t. the whole tree for user $u$ becomes
^$\hat{\mathcal{L}}(u)=\sum_{j\in all\ layers}\hat{\mathcal{L}}_{j}(u).$ (20)
Through simple calculations, the gradient of the multi-class cross-entropy loss function (E.q. 6) with respect to $o_{i}^{j}$ can be obtained as
$\frac{\partial\mathcal{L}^{j}(u,v)}{\partial o_{i}^{j}}=\frac{\partial\left(\log\sum_{i=1}^{c(j)}\exp(o_{i}^{j})-\sum_{i=1}^{c(j)}y_{i}^{j}o_{i}^{j}\right)}{\partial o_{i}^{j}}=p_{i}^{j}-y_{i}^{j}$ (21)
Furthermore, the gradient of the approximate loss (E.q. 18) with respect to $o_{i}^{j}$ is as follows:
^$\frac{\partial\hat{\mathcal{L}}^{j}(u,v)}{\partial o_{i}^{j}}=\sum_{k=1}^{m+1}\mathbb{I}(s_{k}^{j}=i)(\hat{p}_{k}^{j}-y_{s_{k}^{j}}^{j})=\sum_{k=1}^{m+1}\mathbb{I}(s_{k}^{j}=i)\hat{p}_{k}^{j}-y_{i}^{j}$ (22)
The previous literature  [61, 47] has proved that proper specified sampling distribution $\bm{q}$ used in sampled softmax can lead to unbiased estimator of gradient for the original loss (i.e. E.q. (6) and E.q. (7)). We state their theorem in a different way so that the theorem can be compatible with our tree index.
Theorem 2 (Theorem 2.1 of  [47]).
The gradient of loss E.q. (18) (or E.q. (20)) w.r.t. sampled softmax is an unbiased estimator of the gradient of the loss E.q. (6) (or E.q. (7)) w.r.t. full softmax if and only if $q^{j}_{i}\propto\exp{o_{i}^{j}}$ holds where $1\leq i\leq c(j)$ the $i$-th node isn’t a positive node at layer $j$.
This theorem indicates that if the sampling probability is proportional to the exponential of preference w.r.t. the negative node then we can get the training loss’s unbiased estimator of gradient. However, it still leads to linear time complexity w.r.t. the corpus size to make $q^{j}_{i}\propto\exp{o_{i}^{j}}$ as calculating the full softmax. To address this problem, the previous literature  [47, 63] develops the kernel-based methods to approximate the full softmax for inner product models. However, their kernel-based methods aren’t suitable to our model as our preference model is a complicated neural network.
5.2 Top-down Sampling
Uniform sampling is easy to understand and can be implemented quickly, which has been applied in many research and practical work. In our experiments, we can sample the negative nodes at each layer through uniform sampling, i.e. $q_{i}^{j}=\frac{1}{c(j)-1}$ at E.q. (17). Concretely, at layer $j$, each negative node is sampled uniformly with replacement until $m$ negative nodes are sampled. The $i$-th node of layer $j$ is sampled with probability $q_{i}^{j}=\frac{1}{c(j)-1}$ probability at each sampling operation. Besides simpleness, uniform sampling deserves another import advantage that it can lead to good exploration ability. As every negative node can be chosen with equal chance, the sampled nodes deserve good diversity among each other. Furthermore, uniform sampling doesn’t introduce any extra computational burden (e.g kernel-based sampling methods require to compute the kernel to get the distribution $q$.), so that the training process can be quick.
Although uniform sampling is easy to implement and has good exploration capabilities, the uniform distribution may be far from the true softmax distribution. However, we hope that the sampling distribution $q$ can approximate the true softmax, so as to better estimate the gradient of the loss function. Therefore, we propose a new sampling method, by adjusting the retrieval process of beam search, that directly utilizes the index tree and preference model for sampling, making the sampling distribution $q$ closer to the true softmax distribution compared to uniform sampling. Notice that hierarchical beam search retrieval on the tree (i.e., Algorithm 1) from the top layer to the bottom layer can be completed with a logarithmic complexity of the product set size (i.e., $O(\log|I|)$); the additional training time introduced by our proposed sampling method is acceptable.
We will elaborate on the proposed sampling method, which is referred to as Top-down sampling. Firstly, only one node can be retained at each layer when sampling along the tree(i.e., we set $B=1$ in Algorithm 1). Secondly, rather than expanding the child node with the highest preference score, we sample one child node for expansion. This sampling is done with softmax probabilities, which are calculated based on the preference scores of all the child nodes belonging to the current node. By expanding a single node at each layer, we obtain a sample path through the tree from the root to a leaf. The sampled nodes along this path serve as the negative nodes for their respective layers. By repeating this sampling process $m$ times, we are able to obtain $m$ negative nodes for each layer.
Figure 1: Top-down sampling along the tree
Figure 1 illustrates the process of Top-down sampling. The sampling process start from the root, i.e., node $1$. The node $1$’s child nodes are node $2$ and node $3$, with an expanding probability of $\frac{\exp(o_{2})}{\exp(o_{2})+\exp(o_{3})}=0.3$ and $\frac{\exp(o_{3})}{\exp(o_{2})+\exp(o_{3})}=0.7$ respectively. The node $2$ is sampled to expand. The sampling process restart from node $2$ and is repeated until one leaf is sampled. Finally the path $1\rightarrow 2\rightarrow 4\rightarrow 9$ is sampled from the tree, and sampled negative nodes are $\{2,4,9\}$. Since there is only one parent node (we have set $B=1$ in Algorithm 1) and each parent node has only a few child nodes, it is possible to quickly calculate local softmax probabilities at each layer. For any node on the tree, the probability of it being sampled by Top-down sampling is exactly the product of the probabilities on the edges expanded from the root node to that node. For example, in Figure 1, the sampling probability for the node $4$ is $0.18=0.3\times 0.6$, and for the node $9$ is $0.144=0.3\times 0.6\times 0.8$. It’s easy to verify that the sum of the probabilities of each node being sampled is exactly $1$ at any layer.
To obtain an unbiased estimation of the gradient for the training loss (i.e., E.q. (6) or E.q. (7)), according to Theorem 1, the sampling probability for any node should be proportional to the exponential of its corresponding preference score. It can be observed that if the preferences for a parent node and its child nodes satisfy certain rules, this objective can be achieved through top-down sampling. This is formalized in the following theorem:
Theorem 3.
Let $o_{n}$ denote the match score of a user for node $n$, and $Children(n)$ denote the set of child nodes of node $n$. If for any node $n$, the following condition is satisfied:
$\exp(o_{n})\propto\sum_{c\in Children(n)}\exp(o_{c}),$ (23)
and the proportionality coefficients within the same layer are equal, then the probability of sampling the $i$-th node at the $j$-th layer through top-down sampling precisely equals the softmax probability of that node at that layer, that is,
$q_{i}^{j}=\frac{\exp(o_{i}^{j})}{\sum_{k=1}^{C(j)}\exp(o_{k}^{j})}$
where $C(j)$ denotes the number of nodes at layer $j$, and $o_{i}^{j}$ denotes the match score of the $i$-th node at the $j$-th layer.
The condition in Theorem 3.2 (i.e., E.q. (23)) is quite stringent and may not be fully met in practical applications. However, this theorem provides a direction for the design of sampling schemes, guiding us to devise reasonable sampling strategies.
6 Tree learning
Here we introduce how to update the tree (i.e. update the bijective mapping $\pi(\cdot)$) when the preference model is fixed. We regard the level of root node as level 0 and the level of leaf nodes as level $l_{max}$. $\mathcal{A}_{i}=\{(u^{\prime},i^{\prime})|i^{\prime}=item_{i}\}$ means $\mathcal{A}_{i}$ consists of all the user-item pairs where the item is exactly the $i$-th item. $b_{j}(\pi(item_{i}))$ denotes leaf node $\pi(item_{i})$’s ancestor node at level $j$. At first, we use the way of TDM and JTM to construct the initial tree. The tree is used to represent the user interests’ hierarchical information so that similar items should organized in close positions on the level $l_{max}$ of the tree. It’s natural to rely on the category information of items to build the initial tree. Firstly, we shuffle all the categories and put the items together if they share the same category id. If the item belongs to more than one category, a random one is assigned to the item for uniqueness. Secondly, all the items locating in one category are partitioned into two equal parts recursively until each part only contains one item. In this way, we can construct a near-complete binary tree from root node to leaf nodes and some simple balancing strategy can be used to adjust the tree. We use the binary tree in our experiments. In fact, any multi-way tree can be constructed by such way if we partition each cluster into more nearly equal parts at each iteration.
Given the old mapping $\pi$ and each interaction pair set $\mathcal{A}_{i}$, once we fix the preference model, we assign each item to each node of tree step by step from root node to leaf nodes to get the new mapping $\pi$. Firstly, all items are assigned to the root node, i.e. current level $l=0$. Then, we try to assign all the items to the nodes of level $min(l_{max},l+d)$ (where $d$ is a small number and we use $d=7$ in our experiments). The number of assigned items for node $n$ at level $min(l_{max},l+d)$ should be the number of the leaf nodes of the sub-tree who regards the node $n$ as root node. The assigned item set w.r.t. each node at level $min(l_{max},l+d)$ has no intersection and the union is exactly the whole item set (i.e. all the items will be assigned to the nodes at level $min(l_{max},l+d)$ disjointly.). For the $i$-th item $item_{i}$ which is assigned to root node, it need to be assigned to one of the node at level $min(l_{max},l+d)$ and it has $c\leq 2^{d}$ candidate nodes as we use the binary tree. Without loss of generality, we denote these candidate nodes as $n_{1},n_{2},\dots,n_{c}$ and the user $u$’s ($u\in\mathcal{A}_{i}$) preference to node $n_{j}$ ($1\leq j\leq c$) is denoted as $o_{u}^{j}$ computed by the fixed preference model. We define $item_{i}$’s matching score to node $j$ ($1\leq j\leq c$) as follows
$Score(item_{i},n_{j})=\sum_{(u,item_{i})\in\mathcal{A}_{i}}\log\hat{p}(n_{j}|u)$ (24)
where $\hat{p}(n_{j}|u)=\frac{\exp{o_{u}^{j}}}{\sum_{k=1}^{c}\exp{o_{u}^{k}}}$. We rank the $c$ scores and assign $item_{i}$ to the node with the maximum matching score with $item_{i}$. If the number of assigned items breaks the limitation (i.e. the number of assigned items to each node is exactly the number of leaf nodes who regard the current node as root), we assigned $item_{i}$ to the node with second largest matching score with $item_{i}$ or even successive nodes if the number of assigned items still breaks the limitation. By this way, we can assign each item belonging to the root to a node at level $min(l_{max},l+d)$. Now, we regard $n_{1},\dots,n_{c}$ as root nodes of $c$ sub-trees and repeat the process recursively to assign the items to higher layers until each leaf node is assigned a item. Finally, we can get the new mapping $\pi^{\prime}$. To compute the matching score (i.e. E.q. (24)), we need to compute the denominator $\sum_{k=1}^{c}\exp{o_{u}^{k}}$, so the number $d$ need to be small to fit the time complexity and computational resource.
Figure 2: Updating the mapping relationship between items and leaf nodes, where d=2.
To better understand the tree update process, Fig 2 provides an illustrative diagram. Assume there are $8$ items represented by the letters $a$ to $h$, and the step $d$ is set to $2$. Initially, all products are located in the root node. We need to distribute the products from the root node to the layer which is at a distance of $2$ from the root, which contains $4$ nodes, and each node is allocated $2$ items. According to the matching score, item $f$ should be assigned to node $4$, which has the highest score. However, since node $4$ already contains two items, product $f$ is adjusted to node $5$, which has the next highest score. After allocating items to the second layer, we continue to distribute items to a deeper layer. Since the leaves are on the third layer, which is at a distance of $1$ from the second layer, less than the step $d$, we directly assign items to the leaf layer. Each leaf node is allocated one item, and the tree updating is completed.
In fact, our tree learning strategy also carries on the hierarchical clustering from coarse grain to fine grain, and $d$ controls the grain size. The main difference between ours and the one in JTM is the calculation of matching scores. The matching score in JTM can be calculated by summing up the log-likelihood probability as the preference model outputs the probability directly but our strategy relies on the softmax probability when calculating the matching scores.
7 Generalization Analysis
We provide a generalization analysis of the tree-based deep retrieval model. Our analysis employs Rademacher complexity to elucidate the generalization error of the tree model optimized with the rectified loss in conjunction with the sampled softmax technique, which aligns seamlessly with our previously proposed approach, demonstrating its great generalization capability.
The tree model utilizes the DIN [1] to predict preference scores, which has achieved significant success in both the research field and the industry. We formulate it first for the generalization analysis. The user’s historical interaction sequence is used to characterize the user and serves as the input to the model. In this context, we denote the matrix of item embedding vectors that the user $u$ has interacted with as $\mathbf{A}^{(u)}$,
$\mathbf{A}^{(u)}=\left[\boldsymbol{a}_{1}^{(u)},\boldsymbol{a}_{2}^{(u)},\dots,\boldsymbol{a}_{K}^{(u)}\right]\in\mathbb{R}^{d\times K},$ (25)
where we consider the last $K$ recorded item interaction histories and $\boldsymbol{a}_{k}^{(u)}$ represents the embedding of $k$-th item in the user $u$’s historical interaction sequence. The model uses a two-layer fully connected network to compute a weight between the node and the user-history item, which can be expressed as:
$w_{k}^{(u)}(n)=\sigma\left(\boldsymbol{W}_{w}^{(2)}\sigma\left(\boldsymbol{W}_{w}^{(1)}\left[\boldsymbol{a}_{k}^{(u)};\boldsymbol{a}_{k}^{(u)}\odot\boldsymbol{w}_{n}\right]\right)\right)\in\mathbb{R},$ (26)
where $\boldsymbol{W}_{w}^{(1)}\in\mathbb{R}^{h\times 3d},\boldsymbol{W}_{w}^{(2)}\in\mathbb{R}^{1\times h}$, $\boldsymbol{w}_{n}$ is the embedding of the node $n$, and $\sigma$ is the activation function, which is $\operatorname{PReLU}$ in our model setting.
The DIN model splits user interaction sequences into $K^{\prime}$ time windows, enabling a dynamic representation of user interests to enhance the accuracy of personalized recommendations. The $t$-th time window $T_{t}$ of length $k$ is a set of continuous indices, i.e., $T_{t}=\{i_{0},i_{0}+1,\dots,i_{0}+k\}$. The time windows satisfy the following conditions:
$T_{t_{1}}\cap T_{t_{2}}$ $=\ \emptyset,\ \forall t_{1}\neq t_{2}\in\left[K^{\prime}\right]$
$\bigcup_{t=1}^{K^{\prime}}T_{t}$ =$=\{1,2,\dots,K\}$
The interacted item embeddings belonging to the $t$-th time window will be aggregated with weights calculated in E.q. (26):
$\boldsymbol{z}_{t}^{(u)}=\sum_{k\in T_{t}}w_{k}^{(u)}\boldsymbol{a}_{k}^{(u)}$ (27)
Then, the concatenation of $K^{\prime}$ aggregated embeddings and node $n_{i}^{j}$’s embedding will be fed into a multilayer perceptron with $L$ hidden layers to output the user’s preference score for node $n_{i}^{j}$, which can be expressed as follows:
$\begin{split}&o_{i}^{j}(u)=f_{Din}(u,n_{i}^{j})\\ &=\boldsymbol{W}_{L}\cdot\phi_{L-1}\circ\phi_{L-2}\circ\dots\circ\phi_{1}\Big{(}\big{[}\boldsymbol{z}_{1}^{(u)};\boldsymbol{z}_{2}^{(u)};\dots;\boldsymbol{z}_{K^{\prime}}^{(u)};\boldsymbol{w}_{n_{i}^{j}}\big{]}\Big{)}\end{split}$ (28)
where $\boldsymbol{W}_{L}\in\mathbb{R}^{1\times d_{L-1}}$, and the function $\phi_{k}(\mathbf{x})$ is defined:
$\phi_{k}(\mathbf{x})=\phi(\boldsymbol{W}_{k}\mathbf{x})\in\mathbb{R}^{d_{k}\times 1},\forall k\in[L-1].$ (29)
The function $\sigma$ is an element-wise Lipshitz continuous activation function with a Lipshitz constant $c_{\sigma}$, has the property $\sigma(0)=0$ and $\boldsymbol{W}_{k}\in\mathbb{R}^{d_{k}\times d_{k-1}}$ represents the weight matrix. Suppose the norm of the weight matrix is bounded, we can get such function space for the DIN model:
$\begin{split}\mathcal{F}_{Din}=\Big{\{}f:(u,n)\mapsto f_{Din}(u,n)\mid\|\boldsymbol{w}_{n}\|_{2}\leq B_{0},\forall n\in\mathcal{N};\\ \|\boldsymbol{W}_{k}\|_{1}\leq B_{1},\forall k\in[L];\ \|\boldsymbol{W}_{w}^{(j)}\|\leq B_{2},\forall j\in{1,2}\Big{\}}\end{split}$ (30)
Lemma 2 (Rademacher Complexity of DIN).
Suppose $\forall u\in\mathcal{U},\ \forall k\in[K],\|\boldsymbol{a}_{k}^{(u)}\|_{2}\leq B_{a}$, then the Rademacher complexity of ${\mathcal{F}_{Din}}$ can be bounded by
^$\hat{\mathcal{R}}_{m}({\mathcal{F}_{Din}})\leq\frac{2c_{\phi}^{L-1}B_{1}^{L}(B_{0}+KB_{w_{1}}+KB_{w_{2}}\mathcal{\tau})}{\sqrt{m}},$ (31)
where $B_{w_{1}}=B_{2}^{2}c_{\phi}^{2}B^{2}_{a},B_{w_{2}}=B_{2}^{2}c_{\phi}^{2}\left(B_{0}B_{a}^{2}+B_{0}B_{a}\right),\tau=\sqrt{(2B-1)N}/\sqrt{B-1}$.
With the Rademacher complexity of DIN model, we now show the generalization bound of rectified loss (i.e., E.q. (16)) in section 4.3. As the rectified loss for the whole tree is composed of the sum of the loss of each layer, let us consider the generalization bound for the rectified loss for the $j$-th layer, the result is shown as follows:
Theorem 4.
Suppose the function in $\mathcal{F}_{Din}$ is bounded by a constant $B_{\mathcal{F}}$, the following inequality holds with a probability of at least $1-\delta$:
$\begin{split}\mathbb{E}_{(u,y)\sim\mathcal{D}}\left[\widetilde{\mathcal{L}}_{j}(u,y)\right]&\leq\frac{1}{m}\sum_{i=1}^{m}\widetilde{\mathcal{L}}_{j}(u_{i},y_{i})\\ &+8c(j)\frac{c_{\phi}^{L-1}B_{1}^{L}(B_{0}+KB_{w_{1}}+KB_{w_{2}}\mathcal{\tau})}{\sqrt{m}}\\ &+\Big{(}4\log c(j)+8B_{\mathcal{F}}\Big{)}\sqrt{\frac{2\log\left(4/\delta\right)}{m}}.\end{split}$ (32)
As outlined in Section 5, we employ the sampled softmax technique to avoid the extensive computation for the normalization term in the full softmax. Consider the logit vector is $o\in\mathbb{R}^{N}$ for $N$ labels, suppose the $i$-th label is positive and others are negative. A subset of $M$ indices of negative labels $\mathcal{I}^{\prime}_{M}$ is sampled according to the sample distribution $Q$. Then adjusted logit vector $\hat{o}$ can be calculated as E.q. (17). The full softmax loss and sampled softmax loss are calculated on the $o$ and $\hat{o}$ respectively. To incorporate the sampled softmax loss into our theoretical analysis, we show that the full softmax loss can be bounded by the expectation of the expectation of the sampled softmax loss w.r.t. $Q$:
Lemma 3.
Suppose the $i$-th label is positive of $N$ labels, the full softmax loss over logit $o$ will be bounded by the sum of the expectation of the sampled softmax loss over adjusted logit $\hat{o}$ w.r.t. $Q$ and the KL divergence of $Q$ w.r.t. $P$, which can be expressed as follows:
$\begin{split}\log\frac{\sum\limits_{k=1}\limits^{N}\exp o_{k}}{\exp o_{i}}\leq{\mathbb{E}}_{\mathcal{I}^{\prime}_{M}}\left[\log\frac{\sum\limits_{i^{\prime}\in\mathcal{I}^{\prime}_{M}\cup\{i\}}\exp\hat{o}_{i^{\prime}}}{\exp\hat{o}_{i}}\right]+D_{\text{KL}}(Q\|P),\end{split}$
where $Q$ is the negative sample distribution, and $p_{j}\triangleq\exp{o_{j}}/\sum_{k=1}^{N}\exp{o_{k}}$.
Combine Theorem 4 and Lemma 3, and aggregate from the $1$-th layer of the tree to the $H$-th layer, the expectation risk of the rectified loss will be finally bounded as follows:
$\begin{split}\mathbb{E}_{(u,y)\sim\mathcal{D}}\left[\widetilde{\mathcal{L}}(u,y)\right]&\leq\frac{1}{m}\sum_{i=1}^{m}\sum_{j=1}^{H}\mathbb{E}_{\mathcal{I}^{\prime}_{M}}\left[\widehat{\widetilde{\mathcal{L}}}_{j}(u_{i},y_{i})\right]\\ &+\sum_{j=1}^{H}D_{KL}(Q^{j}||P^{j})\\ &+8\tau\frac{c_{\phi}^{L-1}B_{1}^{L}(B_{0}+KB_{w_{1}}+KB_{w_{2}}\mathcal{\tau})}{\sqrt{m}}\\ &+\Big{(}4\log(\tau)+8HB_{\mathcal{F}}\Big{)}\sqrt{\frac{2\log\left(4/\delta\right)}{m}}.\end{split}$ (33)
The above equation shows the factors affecting the expectation risk of the tree model optimized with rectified loss and sampled softmax technique. The first term on the right side of the inequality is the expectation of the loss w.r.t the sampling distribution, which can be regarded as the empirical loss we are optimizing in practice. The second term is the KL divergence of the negative sampling distribution relative to the distribution of the softmax preference scores at each layer. It can be observed that the closer the negative sampling distribution is to the softmax preference scores, the lower the generalization risk. This precisely aligns with the philosophy of our proposed top-down sampling method, explaining why the top-down sampling method works from the perspective of generalization. The third term reflects the impact of model parameters and tree structure on generalization risk. From it, we can observe that When the number of items (i.e., $N$) is fixed, a larger branching factor $B$ leads to a smaller $\tau$, which allows us to obtain a tighter generalization bound.
8 Experiments
We conduct a series of experiments designed to answer the following research questions related to the joint optimization framework of preference models and tree indexes: RQ-1: Is the multi-class training mode more suitable than the binary classification training mode for the tree model? RQ-2: Is the previously proposed method, which uses sampled-softmax theory for negative sampling to accelerate training, effective? RQ-3: Can the TDR model learn a reasonable mapping relationship between items and leaf nodes? RQ-4: Does the performance of the proposed sampling method fluctuate with the number of negative samples selected? RQ-5: Is the proposed loss with rectified labels effective in mitigating the sub-optimality of the multi-class cross-entropy loss in practice?
8.1 Datasets
The experiments are conducted on the four real-world datasets [^2] https://drive.google.com/drive/folders/1ahiLmzU7cGRPXf5qGMqtAChte2eYp9gI: Movie Lens 10M (abbreviated as Movie), MIND Small Dev(abbreviated as MIND), Amazon Books(abbreviated as Amazon), Tmall Click (abbreviated as Tmall). Movie is a public film rating dataset, MIND is a news recommendation dataset collected by Microsoft, Amazon is a dataset documenting book purchases and ratings on the Amazon e-commerce site, Tmall records the shopping behaviors of users on Alibaba’s Tmall marketplace. We processed all datasets into a form of implicit feedback, where interactions between users and items are marked as 1, and non-interactions as 0. Additionally, users with fewer than 15 interactions were filtered out, with the post-filtering data statistics presented in Table I. In each dataset, 10% of users are randomly selected as validation users, 10% as test users, and all remaining users as training users. For validation and test users, the first half of the behavior is considered the interaction history, and the latter half is considered as the labels to predict.
TABLE I: Statistics of datasets
Dataset #User #Item #Interaction Density
Movie 69,878 10,677 10,000,054 1.34%
MIND 36,281 7,129 5,610,960 2.16%
Amazon 29,980 67,402 2,218,926 0.11%
Tmall 139,234 135,293 10,487,585 0.05%
8.2 Baselines
We compared the proposed TDR model with other related algorithms, which fall into two broad categories: 1) models that rely on brute-force search without indexing, and 2) tree-based models similar to our TDR approach. It should be noted that the former requires a significant amount of time to perform retrieval, comparing them with index-based recall models is not fair, and their results are only used as a reference. The introductions of these algorithms are as follows:
  • Item-CF [64]: It is a basic collaborative filtering algorithm and is widely used in recommendation tasks. We implement it by ourselves and the cosine correlation coefficient is used to measure the similarity between items.
  • DIN [1]: DIN utilizes deep models with an attention mechanism to capture the user preference for items. Since it cannot be used for efficient recommendation, the performance based on brute-force search is reported as a reference metric.
  • YoutubeDNN [59]: YouTubeDNN is a recommendation system that uses the inner products as the preference scores between users and items. After all the user embedding and item embedding are learned, the recommendation task is transformed into a Maximum Inner Product Search (MIPS). In experiments, the performance of YouTubeDNN in MIPS using brute-force search is reported as a reference.
  • PLT [65]: PLT can be used to organize and index a large number of candidate items to facilitate the rapid retrieval of items most relevant to user interests in recommendation.
  • TDM [11]: TDM is a tree-based recommendation framework proposed by Alibaba Group, which simultaneously trains the preference model of users for products as well as the tree index structure to perform efficient recommendations through beam search during inference.
  • JTM [12]: JTM is an enhanced version of TDM. It introduces a method for jointly learning the index structure and the user preference prediction model, aiming to achieve better performance.
  • OTM [54]: OTM corrects the node labels using the concept of Bayesian Optimality under beam search and reduces the training-testing discrepancy by sampling negative nodes through beam search, resulting in improved performance.
8.3 Metric
To evaluate the performance of the retrieval model, this work uses $Precision@K$, $Recall@K$, and $F\text{-}measure@K$ as the evaluation criteria. Suppose $\mathcal{P}(u)$ ($\left|\mathcal{P}(u)\right|=K$) denotes the set of items recalled by the model for user $u$, and $\mathcal{G}(u)$ represents the set of items that the user has actually interacted with, i.e., the tags associated with user $u$. Then the evaluation metrics for user $u$, $Precision@K$ and $Recall@K$, are calculated as follows:
$\begin{split}Precision@K(u)=\frac{|\mathcal{P}(u)\cap\mathcal{G}(u)|}{K}\ ,\\ Recall@K(u)=\frac{|\mathcal{P}(u)\cap\mathcal{G}(u)|}{|\mathcal{G}(u)|}\ ,\ \ \ \end{split}$
and $F\text{-}measure@K$ is calculated as follows:
$F\text{-}measure@K(u)=\frac{2\cdot Precision@K(u)\cdot Recall@K(u)}{Precision@K(u)+Recall@K(u)}.$
8.4 Settings
In the experiments, TDM, JTM, and TDR use the DIN model or its variance as the preference score model. Following the settings in the original papers of TDM[11] and JTM[12], each user retains only the most recent 70 interactions with items. When the number of items a user has interacted with is less than 70, the sequence is padded with zeros. The 70-th item in the interaction sequence is considered the predicted label item. Furthermore, for the TDM, JTM, and TDR algorithms, the first 69 items in the interaction sequence are segmented using 10 sliding windows, with the number of items in each window being $\left[1,1,1,2,2,2,10,10,20,20\right]$ respectively. TDM, JTM, and TDR are all set to include 12 trees. TDR utilizes two sampling methods: Uniform Sampling and Top-down Sampling, which are denoted as TDR(U) and TDR(T), respectively. Additionally, TDR incorporates a modification that includes a loss with the rectified label in the Top-down sampling method, which is referred to as TDR(T-RL).
For TDR(T-RL), as mentioned in the previous section, we need to use a trained small model to estimate the conditional probability. In our experiment, we choose the SASRec as the estimator. SASRec is one of the most representative sequential models with the self-attention mechanism. Concretely, for any user $u$, a historical sequence as the $u$’s representation in the experiment will be fed into the SASRec model and the output semantic vector will be used to compute an inner product with every item embedding, and finally, the inner products over the entire item set will be fed into a softmax function and the conditional probability can be obtained. We train the SASRec model with the same trainset as other algorithms. We tune the number of encoder layers, the number of heads, and the model dimension for each dataset on the validation dataset respectively. Finally, we SASRec model on the testset and it performs at most 0.8% lower than TDR(T) on metric $F\text{-}measure@20$ across four datasets, which means it’s a well-behaved estimator. And notice SASRec is a two-tower model, so the additional computation due to the probability estimating is relatively small and acceptable. In fact, in our experiment, the proportion of time spent on estimating probabilities compared to the total time is approximately 17.5% to 22.6% across four datasets, which is completely acceptable.
The number of negative samples per layer for both TDR(U) and TDR(T) is set to 70, while for TDM and JTM, the number of negative samples on the Movie, MIND, Amazon, and Tmall datasets are set to 6, 3, 5, and 6, respectively. DIN includes a representation layer and a multilayer perceptron of sizes $[128,64,2]$, outputting the probabilities of the probability of the user liking the corresponding item. YouTubeDNN [59] includes a representation and a multilayer perceptron of sizes $\left[128,64,24\right]$, with the number of negative samples set to 1,000. For TDR and the four comparison algorithms, during the training phase, the first 69 user-item interactions in the interaction history are considered as the sequence representing the user, with the 70-th item being the label to predict. In all methods, the item representation dimension is set to 24, Adam[66] is used as the optimizer, and the learning rate is set to $1.0e$-$3$ with exponential decay. For tree-based beam search, the candidate set size (Beam size) is set to 150. All experiments were conducted on a Linux server equipped with a $3.00$ GHz Intel CPU, $300$ GB of main memory, and NVIDIA $20/30$ series GPUs
8.5 Comparation results
TABLE II: Experimental results for different algorithms on different datasets.
Dataset Algorithm top-K=20 top-K=40 top-K=60
Precision Recall F-measure Precision Recall F-measure Precision Recall F-measure
Movie Item-CF 0.1058 0.0508 0.0600 0.1211 0.1218 0.1037 0.1266 0.1886 0.1292
DIN 0.2123 0.1134 0.1303 0.1948 0.1993 0.1694 0.1798 0.2667 0.1838
YoutubeDNN 0.2043 0.1113 0.1270 0.1874 0.1924 0.1630 0.1746 0.2560 0.1770
PLT 0.1606 0.0859 0.0981 0.1478 0.1489 0.1264 0.1385 0.2006 0.1389
TDM 0.1901 0.1012 0.1159 0.1741 0.1770 0.1499 0.1624 0.2383 0.1641
JTM 0.2141 0.1169 0.1335 0.1922 0.1993 0.1682 0.1770 0.2650 0.1815
OTM 0.2321 0.1290 0.1470 0.2009 0.2086 0.1767 0.1822 0.2687 0.1862
TDR(U) 0.2367 0.1231 0.1431 0.2054 0.2032 0.1759 0.1842 0.2616 0.1850
TDR(T) 0.2511 0.1343 0.1546 0.2135 0.2169 0.1853 0.1893 0.2772 0.1924
TDR(T-RL) 0.2545 0.1374 0.1580 0.2176 0.2240 0.1905 0.1936 0.2876 0.1985
Mind Item-CF 0.2243 0.1218 0.1416 0.1962 0.1977 0.1740 0.1769 0.2518 0.1833
DIN 0.4145 0.1889 0.2295 0.3466 0.2910 0.2738 0.3040 0.3620 0.2855
YoutubeDNN 0.4025 0.1850 0.2243 0.3357 0.2851 0.2671 0.2936 0.3541 0.2775
PLT 0.2341 0.0927 0.1168 0.2024 0.1532 0.1494 0.1818 0.2011 0.1626
TDM 0.3156 0.1365 0.1687 0.2706 0.2236 0.2116 0.2375 0.2832 0.2225
JTM 0.3457 0.1547 0.1892 0.2806 0.2354 0.2213 0.2432 0.2925 0.2288
OTM 0.3724 0.1609 0.1978 0.3144 0.2523 0.2404 0.2785 0.3188 0.2544
TDR(U) 0.3904 0.1804 0.2188 0.3063 0.2621 0.2448 0.2612 0.3191 0.2482
TDR(T) 0.4044 0.1853 0.2250 0.3318 0.2808 0.2634 0.2889 0.3472 0.2726
TDR(T-RL) 0.4210 0.1938 0.2350 0.3401 0.2869 0.2695 0.2957 0.3544 0.2787
Amazon Item-CF 0.0175 0.0178 0.0169 0.0284 0.0554 0.0360 0.0314 0.0860 0.0440
DIN 0.0665 0.0476 0.0506 0.0583 0.0803 0.0605 0.0521 0.1060 0.0626
YoutubeDNN 0.0606 0.0399 0.0436 0.0522 0.0682 0.0527 0.0470 0.0910 0.0553
PLT 0.0238 0.0147 0.0162 0.0201 0.0245 0.0193 0.0181 0.0321 0.0202
TDM 0.0514 0.0333 0.0365 0.0419 0.0530 0.0416 0.0361 0.0674 0.0418
JTM 0.0606 0.0411 0.0444 0.0507 0.0674 0.0517 0.0452 0.0892 0.0537
OTM 0.0704 0.0440 0.0486 0.0584 0.0710 0.0566 0.0514 0.0923 0.0583
TDR(U) 0.0709 0.0502 0.0535 0.0594 0.0820 0.0619 0.0521 0.1067 0.0630
TDR(T) 0.0757 0.0517 0.0557 0.0613 0.0829 0.0631 0.0533 0.1073 0.0639
TDR(T-RL) 0.0777 0.0542 0.0580 0.0626 0.0847 0.0644 0.0544 0.1093 0.0651
Tmall Item-CF 0.0093 0.0084 0.0083 0.0181 0.0314 0.0216 0.0205 0.0498 0.0273
DIN 0.0463 0.0332 0.0357 0.0369 0.0516 0.0394 0.0317 0.0654 0.0391
YoutubeDNN 0.0331 0.0211 0.0236 0.0276 0.0349 0.0278 0.0243 0.0457 0.0287
PLT 0.0167 0.0111 0.0122 0.0133 0.0176 0.0137 0.0114 0.0224 0.0137
TDM 0.0310 0.0209 0.0229 0.0253 0.0335 0.0262 0.0220 0.0433 0.0266
JTM 0.0394 0.0275 0.0298 0.0315 0.0427 0.0331 0.0269 0.0538 0.0327
OTM 0.0459 0.0283 0.0318 0.0361 0.0433 0.0353 0.0307 0.0546 0.0353
TDR(U) 0.0445 0.0315 0.0340 0.0361 0.0496 0.0382 0.0313 0.0634 0.0383
TDR(T) 0.0464 0.0329 0.0355 0.0364 0.0503 0.0385 0.0312 0.0634 0.0382
TDR(T-RL) 0.0482 0.0342 0.0369 0.0373 0.0514 0.0395 0.0317 0.0644 0.0388
The experiment results are shown in Table II. In the table, the best results are highlighted in bold, while the second-best results are indicated in italics. Based on the results, we can make the following findings:
[…]
  • The results of DFR(T) are always better than those of DFR(U). For instance, with top-K set to 20, in terms of Precision, DFR(T) achieved improvements over DFR(U) of 1.3%, 16.8%, 125%, and 2.2% across four datasets, respectively. This phenomenon indicates that the proposed top-down sampling better satisfies the sampling probability requirements of Sampled-softmax, meaning the sampling probabilities corresponding to top-down sampling are closer to the softmax probabilities. This phenomenon validates the theoretical analysis presented in this chapter (i.e., Theorem 3.2), and also adequately addresses RQ-2.
  • DFR(T) outperforms the brute-force retrieval of YouTubeDNN across all datasets and even surpasses brute-force retrieval of DIN on the Movie and Amazon datasets. This phenomenon demonstrates that joint optimization of the index and matching model can exhibit stronger performance than optimizing them independently, echoing one of the research motivations of this chapter. Moreover, with the help of the index structure, the matching model can exclude some distracting items in certain scenarios, enhancing the precision of the recall results.
8.6 Effectiveness of index updating
We demonstrate the effectiveness of the tree index update strategy proposed in TDR, specifically the effectiveness of updating the mapping between the tree’s leaf nodes and items. We illustrate the variations in Precision, Recall, and F-measure in response to updates made to the tree index, with the top-K parameter set at 20, across four datasets: Movie, MIND, Amazon, and Tmall. Since both TDM and JTM were proposed by Alibaba Group, and JTM[12] is an improved version of TDM[11], with JTM’s updating strategy being more aligned with the model than TDM’s, we mainly compare TDR with JTM under Top-down sampling and uniform random sampling. The experimental results are shown in Figure 3. Observing the experimental results, the following findings can be made:
  • TDR, when based on Top-down sampling, consistently exhibits superior performance, significantly outshining JTM. Even TDR with uniform random sampling tends to outperform JTM, with the only exception being its performance on the Amazon dataset. This result further illustrates the superiority of the training mode based on multi-classification over the binary classification training mode used by JTM, addressing RQ-1.
  • The TDR method based on Top-down sampling consistently outperforms the one based on uniform random sampling, indicating that Top-down sampling can estimate the Softmax gradient more accurately than uniform random sampling. This validates the proposed sampling method’s rationale and echoes the theoretical conclusions of Top-down sampling, also addressing RQ-2.
  • TDR, both when based on Top-down sampling and when based on uniform random sampling, exhibits an initial increase and then a gradual convergence in all three metrics across various datasets. This phenomenon indicates that the tree update strategy employed by TDR can effectively acquire a reasonable mapping between items and the tree’s leaf nodes. The index structure is able to co-evolve with the matching model, achieving improved recall performance. This result satisfactorily addresses RQ-3.
(a) Movie
(b) MIND
(c) Amazon
(d) Tmall
Figure 3: TDR Tree Updating
8.7 Sensitivity w.r.t. number of negative samples
We analyze the impact of the number of negative samples and the size of the candidate set during Beam Search (i.e., Beam size) on model performance. All experiments in this subsection are conducted on the initial tree constructed based on the Movie and MIND datasets respectively. The difference between TDM and JTM lies in the way the tree is updated; therefore, we only need to test the performance of TDR based on Top-down sampling, TDR based on uniform random sampling, and JTM. Additionally, we set the top-K to 20 in this section of the experiments.
(a) Movie
(b) MIND
(c) Amazon
(d) Tmall
Figure 4: The experimental results of TDR varying the number of negative sampling samples
The experimental results are shown in Figure 4. As the number of negative samples for sampling increases, the recall rates of the TDR models based on both sampling methods gradually improve and eventually converge. Among them, TDR based on Top-down sampling outperforms TDR based on uniform random sampling, indicating that Top-down sampling is closer to the true softmax sampling strategy, and can more accurately estimate the multinomial cross-entropy loss function and model gradients. The performance of the JTM model fluctuates significantly with the change in the number of samples and generally declines, suggesting that the training model based on binary classification is very sensitive to the number of samples. As the number of negative samples increases, the loss function gradually becomes dominated by negative samples, severely affecting the model’s performance. This experiment answers RQ-4 effectively.
9 Conclusion
The existing tree-base recommendation systems (e.g. TDM and JTM) regard the preference model training task as a binary-class classification problem. Their training mode suffers from the gap between training and prediction due to the lack of competition among nodes at the same layer. To address this problem, We develop a layer-wise training mode that regrades the training task as a multi-class classification problem and we provide the tree learning strategy to update the tree. Through further inspection, we found the suboptimality of the multi-class classification training mode under beam search and proposed a rectified loss to correct it. Given the significant time cost of calculating Softmax probabilities with large-scale datasets, we developed a sampling strategy based on the Sampled-Softmax theory to estimate the Softmax gradient, thus accelerating the training process. Finally, we provide a generalization analysis of the tree-based deep retrieval model, demonstrating the great generalization capability of our proposed methods. The experimental results validate the effectiveness of our proposed TDR.
Appendix A Proofs of Lemma, proposition, and Theorem
A.1 Proof of Lemma 1
Proof.
We complete the proof by contradiction. When $c(j+1)\leq k$, $\operatornamewithlimits{argTopk}\limits_{i\in[c(j+1)]}\mathbb{E}[\tilde{z}^{j+1}_{i}]=\{[c(j+1)]\}=\mathcal{B}^{j+1}(u)$, so we only need to consider the case where $c(j+1)>k$.
Denote $\operatornamewithlimits{argTopk}\limits_{i\in[c(j+1)]}\mathbb{E}[\tilde{z}^{j+1}_{i}]$ as $\{\mathcal{C}^{j+1}_{1},\mathcal{C}^{j+1}_{2},\dots,\mathcal{C}^{j+1}_{m}\}$, and suppose $\mathcal{B}^{j+1}(u)\notin\operatornamewithlimits{argTopk}\limits_{i\in[c(j+1)]}\mathbb{E}[\tilde{z}^{j+1}_{i}]$, we can make the following derivation:
$\mathcal{B}$ (${}^{j+1}(u)\notin\operatornamewithlimits{argTopk}\limits_{i\in[c(j+1)]}\mathbb{E}[\tilde{z}^{j+1}_{i}]$
$\Longrightarrow\mathcal{B}^{j+1}(u)\neq\mathcal{C}^{j+1}_{i}(1\leq i\leq m)$
$\Longrightarrow\exists i_{1},i_{2}\in\left[c(j+1)\right]\ s.t.\ i_{1}\in\mathcal{B}^{j+1}(u)\wedge i_{1}\notin\mathcal{C}^{j+1}_{i}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\wedge i_{2}\notin\mathcal{B}^{j+1}(u)\wedge i_{2}\in\mathcal{C}_{i_{0}}^{j+1}$
$\Longrightarrow\mathbb{E}[\tilde{z}^{j+1}_{i_{1}}]<\mathbb{E}[\tilde{z}^{j+1}_{i_{2}}]\wedge o^{j+1}_{i_{1}}\geq o^{j+1}_{i_{2}}$
$\Longrightarrow i_{1}\in\widetilde{\mathcal{B}}^{j+1}(u)\wedge i_{2}\notin\widetilde{\mathcal{B}}^{j+1}(u)$
$\Longrightarrow\delta(\rho^{j}(n_{i_{1}}^{j+1}))\in\mathcal{B}^{j}(u)\wedge\delta(\rho^{j}(n_{i_{2}}^{j+1}))\notin\mathcal{B}^{j}(u).$
Let’s denote $\mathcal{B}^{j}(u)$ as $\{n^{j}_{s_{1}},n^{j}_{s_{2}},\dots,n^{j}_{s_{k}}\}$, where the indices satisfy $o^{j}_{s_{1}}\geq o^{j}_{s_{2}}\geq\dots\geq o^{j}_{s_{k}}$. Notice that $R(o^{j},\mathbb{E}[\widetilde{Z}^{j}])$ and $\mathbb{E}[\widetilde{z}^{j}_{s_{i}}]=\max\limits_{n^{\prime}\in\mathfrak{L}(n_{s_{1}}^{j})}\eta_{\pi(n^{\prime})}$, we have:
$\max\limits_{n^{\prime}\in\mathfrak{L}(n_{s_{1}}^{j})}\eta_{\pi(n^{\prime})}\geq\max\limits_{n^{\prime}\in\mathfrak{L}(n_{s_{2}}^{j})}\eta_{\pi(n^{\prime})}\geq\dots\geq\max\limits_{n^{\prime}\in\mathfrak{L}(n_{s_{k}}^{j})}\eta_{\pi(n^{\prime})}.$ (34)
For each node $n_{s_{i}}^{j}$, we then consider its node child $n^{j+1}_{c_{i}}\in\mathcal{D}(n_{s_{i}}^{j})$ where $c_{i}=\operatornamewithlimits{argmax}\limits_{i\in\{\delta(n)\mid n\in\mathcal{D}(n^{j}_{s_{i}})\}}\mathbb{E}[\tilde{z}^{j+1}_{i}]$, and the following equation
$\max\limits_{n^{\prime}\in\mathfrak{L}(n^{j+1}_{c_{i}})}\eta_{\pi(n^{\prime})}=\max\limits_{n^{\prime}\in\mathfrak{L}(n^{j}_{s_{i}})}\eta_{\pi(n^{\prime})}$ (35)
holds for $1\leq i\leq k$. Combine the E.q. (34) and E.q. (35), we can get:
$\max\limits_{n^{\prime}\in\mathfrak{L}(n^{j+1}_{c_{1}})}\eta_{\pi(n^{\prime})}\geq\max\limits_{n^{\prime}\in\mathfrak{L}(n^{j+1}_{c_{2}})}\eta_{\pi(n^{\prime})}\geq\cdots\geq\max\limits_{n^{\prime}\in\mathfrak{L}(n^{j+1}_{c_{k}})}\eta_{\pi(n^{\prime})}$
Then, with $R(o^{j+1},\mathbb{E}[\widetilde{Z}^{j+1}])$ and $\mathbb{E}[\widetilde{z}^{j+1}_{c_{i}}]=\max\limits_{n^{\prime}\in\mathfrak{L}(n_{c_{i}}^{j+1})}\eta_{\pi(n^{\prime})}$, we have $o^{j+1}_{c_{1}}\geq o^{j+1}_{c_{2}}\geq\dots\geq o^{j+1}_{c_{k}}$, which means that the rank of $o^{j+1}_{c_{k}}$ in the elements of $o^{j+1}_{i}$ is not lower than $k$. And because $i_{1}\in\mathcal{B}^{j+1}(u)$, we have
$o^{j+1}_{i_{1}}\geq o^{j+1}_{c_{k}}$ $\Longrightarrow\mathbb{E}[\tilde{z}^{j+1}_{i_{1}}]\geq\mathbb{E}[\tilde{z}^{j+1}_{c_{k}}]\Longrightarrow\mathbb{E}[\tilde{z}^{j+1}_{i_{2}}]>\mathbb{E}[\tilde{z}^{j+1}_{c_{k}}]$
$\Longrightarrow\max\limits_{n^{\prime}\in\mathfrak{L}(n_{i_{2}}^{j+1})}\eta_{\varphi(n^{\prime})}>\max\limits_{n^{\prime}\in\mathfrak{L}(c_{k})}\eta_{\varphi(n^{\prime})}.$ (36)
Notice $i_{\rho}\triangleq\delta(\rho^{j}(n_{i_{2}}^{j+1}))\notin\mathcal{B}^{j}(u)\wedge s_{k}\in\mathcal{B}^{j}(u)$ and $\mathcal{B}^{j}(u)\in\operatornamewithlimits{argTopk}\limits_{i\in[c(j)]}\mathbb{E}[\tilde{z}^{j}_{i}]$, we can have $\mathbb{E}[\tilde{z}^{j}_{s_{k}}]\geq\mathbb{E}[\tilde{z}^{j}_{i_{\rho}}]$. Therefore, we have:
$\begin{split}\max\limits_{n^{\prime}\in\mathfrak{L}(n^{j}_{s_{k}})}\eta_{\pi(n^{\prime})}\geq\max\limits_{n^{\prime}\in\mathfrak{L}(n^{j}_{i_{\rho}})}\eta_{\pi(n^{\prime})}\geq\max\limits_{n^{\prime}\in\mathfrak{L}(n^{j+1}_{i_{2}})}\eta_{\pi(n^{\prime})},\end{split}$ (37)
where the last inequality holds as the $\mathfrak{L}(n^{j+1}_{i_{2}})\subseteq\mathfrak{L}(n^{j}_{i_{\rho}})$. Combine the E.q. (35) and E.q. (37), we can get:
$\max\limits_{n^{\prime}\in\mathfrak{L}(n^{j+1}_{c_{k}})}\eta_{\pi(n^{\prime})}\geq\max\limits_{n^{\prime}\in\mathfrak{L}(n^{j+1}_{i_{2}})}\eta_{\pi(n^{\prime})}.$ (38)
E.q. (36) and E.q. (37) induces the contradiction, so the proposition holds. ∎
A.2 Proof of Proposition 3
Proof.
Let’s consider the retrieval process along the tree for any user $u$ and any beam size $k$. It is obvious that when $c(j)\leq k$, $\mathcal{B}^{j}(u)=\{[c(j)]\}\in\operatornamewithlimits{argTopk}\limits_{i\in[c(j)]}\mathbb{E}[\tilde{z}^{j}_{i}]$.
Let $j^{*}=\operatornamewithlimits{argmin}\limits_{\{j:c(j)>k\}}c(j)$. When $11\leq j\leq j^{*}-$ 1, $\mathcal{B}^{j}(u)\in\operatornamewithlimits{argTopk}\limits_{j\in[c(j)]}\mathbb{E}[\tilde{z}^{j}_{i}]$ holds; when $j\geq j^{*}$, by recursively applying Lemma 1 from $j^{*}$ to $H-1$, we finally get $\mathcal{B}^{H}(u)\in\operatornamewithlimits{argTopk}\limits_{i\in[c(H)]}\mathbb{E}[\tilde{z}^{H}_{i}]$. Notice that $\mathbb{E}[\tilde{z}_{i}^{H}]=\eta_{\pi(n_{i}^{H})}$, that is to say $\mathcal{B}^{H}(u)\in\operatornamewithlimits{argTopk}\limits_{i\in[c(H)]}\eta_{\pi(n_{i}^{H})}$, so the final retrieved item set $\widehat{\mathcal{Y}}=\{\pi(n_{i}^{H})|i\in\mathcal{B}^{H}(u)\}\in\operatornamewithlimits{argTopk}\limits_{y\in\mathcal{Y}}\eta_{y}(u)$, i.e., the tree model is top-$k$ retrieval Bayes optimal. ∎
A.3 Proof of Proposition 2
Proof.
By taking $\psi(x)=\sum_{m=1}^{M}x_{m}\log x_{m}$ and $g(o)_{m}=\exp(o_{m})/\sum_{k=1}^{M}\exp(o_{k})$, we have
$D_{\psi}$ $\left(e_{i}(w),g(o)\right)$
$=\psi\left(e_{i}(w)\right)-\psi\left(g(o)\right)-\nabla\psi(g(o))^{T}\left(e_{i}(w)-g(o)\right)$
$=w\log w-\sum_{m=1}^{M}g(o)_{m}\log g(o)_{m}$
$\quad-\Big{<}\left(\log g(o)_{1}+1,\dots,\log g(o)_{i}+1,\dots,\log g(o)_{m}+1\right),$
$\quad\quad\quad\left(-g(o)_{1},\dots,w-g(o)_{i},\dots,-g(o)_{m}\right)\Big{>}$
$=w\log w-\sum_{m=1}^{M}g(o)_{m}\log g(o)_{m}$
$\quad-(-\sum_{m=1}^{M}g(o)_{m}\log g(o)_{m}-1+w\log g(o)_{i}+w)$
$=-w\log g(o)_{i}+w\log w+w-1=\widetilde{\mathcal{L}}(e_{i}(w),o).$
A.4 Proof of Theorem 2
Proof.
We use induction to prove this theorem. Let $nodes(j)$ denote all the nodes on the $j$-th level of the tree. Assuming the sampling probability of node $n$ on the $j$-th level according to top-down sampling is $p(n)=\frac{\exp{o_{n}}}{\sum\limits_{n^{\prime}\in nodes(j)}\exp{o_{n^{\prime}}}}$. When $j=0$, we sample the root node with probability $1$, equaling to the root node’s softmax probability; when $j=1$, notice that $Children(root)=nodes(1)$, so the assumption holds due to the definition of top-down sampling. Suppose that $n_{c}$ is any child node of node $n$ in layer $j+1$. The sampling probability of node $n_{c}$ is calculated as:
$p(n_{c})$ $=p(n)\cdot\frac{\exp{o_{n_{c}}}}{\sum_{n^{\prime}_{c}\in\text{Children}(n)}\exp{o_{n^{\prime}_{c}}}}$ (19)
$=\frac{\exp{o_{n}}}{\sum_{n^{\prime}\in\text{nodes}(j)}\exp{o_{n^{\prime}}}}\cdot\frac{\exp{o_{n_{c}}}}{\sum_{n^{\prime}_{c}\in\text{Children}(n)}\exp{o_{n^{\prime}_{c}}}}$
According to E.q. (23), and with the assumption that the proportionality coefficients within the same layer are equal, we have
$\sum_{n^{\prime}\in\text{nodes}(j)}\exp{o_{n^{\prime}}}$ $\propto\sum_{n^{\prime}\in\text{nodes}(j)}\sum_{n^{\prime}_{c}\in\text{Children}(n^{\prime})}\exp{o_{n^{\prime}_{c}}}$ (20)
$=\sum_{n^{\prime\prime}\in\text{nodes}(j+1)}\exp{n^{\prime\prime}}.$
Therefore, by substituting equations E.q. (23) and E.q. (20) into E.q. (19), we can obtain:
$p(n_{c})=\frac{\exp{o_{n_{c}}}}{\sum_{n^{\prime\prime}\in\text{nodes}(j+1)}\exp{n^{\prime\prime}}}.$
That is, when sampling to the $j+1$-th level in the top-down sampling, the probability of sampling the corresponding node is equivalent to the probability of performing softmax on the $j+1$-th level. As $j$ increases, when reaching the layer where the leaf nodes are located, this conclusion also holds. ∎
A.5 Proof of Lemma 2
Proof.
For any vectors $\boldsymbol{v},\boldsymbol{u}_{i}\in\mathbb{R}^{d},\|\boldsymbol{v}\|_{1}\leq B_{v}$, notice the following inequality:
$\sup_{\boldsymbol{v}}\sum_{i}\boldsymbol{v}^{\mathsf{T}}\boldsymbol{u}_{i}\leq B_{v}\max_{j\in[d]}\left|\sum_{i}\boldsymbol{e}_{j}\boldsymbol{u}_{i}\right|\leq\sum_{i}B_{v}\max_{\begin{subarray}{c}j\in[d]\\ s\in\{-1,1\}\end{subarray}}s\boldsymbol{e}_{j}\boldsymbol{u}_{i}.$ (39)
By applying the E.q. (39), we can get:
^$\hat{\mathcal{R}}_{m}$ $(\mathcal{F}_{Din})=\mathbb{E}_{\sigma}\left[\sup\limits_{f\in\mathcal{F}_{Din}}\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}f(u_{i},n_{i})\right]$
$=\mathbb{E}_{\sigma}\Bigg{[}\sup\limits_{\boldsymbol{z},\boldsymbol{w}_{n},\{\boldsymbol{W}_{k}\}^{L}_{k=1}}\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\Big{<}\boldsymbol{W}_{L},\phi_{L-1}\circ$
$\quad\quad\phi_{L-2}\circ\dots\circ\phi_{1}\left(\boldsymbol{z}^{(u_{i})}_{1};\boldsymbol{z}^{(u_{i})}_{2};\dots;\boldsymbol{z}^{(u_{i})}_{K^{\prime}};\boldsymbol{w}_{n_{i}}\right)\Big{>}\Bigg{]}$
$\underset{(a)}{\leq}\|\boldsymbol{W}_{L}\|_{1}\mathbb{E}_{\sigma}\Bigg{[}\sup\limits_{\begin{subarray}{c}s\in\{-1,1\},j\in[d_{L-1}]\\ \boldsymbol{z},\boldsymbol{w}_{n},\{\boldsymbol{W}_{k}\}_{k=1}^{L-1}\end{subarray}}\frac{1}{m}\sum_{i=1}^{m}s\sigma_{i}\Big{<}\boldsymbol{e}_{j},\phi_{L-1}\circ$
$\quad\quad\phi_{L-2}\circ\dots\circ\phi_{1}\left(\boldsymbol{z}^{(u_{i})}_{1};\boldsymbol{z}^{(u_{i})}_{2};\dots;\boldsymbol{z}^{(u_{i})}_{K^{\prime}};\boldsymbol{w}_{n_{i}}\right)\Big{>}\Bigg{]}$
$\underset{(b)}{\leq}c_{\phi}\|\boldsymbol{W}_{L}\|_{1}\mathbb{E}_{\sigma}\Bigg{[}\sup\limits_{\begin{subarray}{c}s\in\{-1,1\},j\in[d_{L-1}]\\ \boldsymbol{z},\boldsymbol{w}_{n},\{\boldsymbol{W}_{k}\}_{k=1}^{L-1}\end{subarray}}\frac{1}{m}\sum_{i=1}^{m}s\sigma_{i}\Big{<}\boldsymbol{e}_{j},\boldsymbol{W}_{L-1}\cdot$
$\quad\quad\phi_{L-2}\circ\dots\circ\phi_{1}\left(\boldsymbol{z}^{(u_{i})}_{1};\boldsymbol{z}^{(u_{i})}_{2};\dots;\boldsymbol{z}^{(u_{i})}_{K^{\prime}};\boldsymbol{w}_{n_{i}}\right)\Big{>}\Bigg{]}$
$\underset{(c)}{\leq}c_{\phi}\|\boldsymbol{W}_{L}\|_{1}\|\boldsymbol{W}_{L-2}\|_{1}\mathbb{E}_{\sigma}\Bigg{[}\sup\limits_{\begin{subarray}{c}s\in\{-1,1\},j\in[d_{L-2}]\\ \boldsymbol{z},\boldsymbol{w}_{n},\{\boldsymbol{W}_{k}\}_{k=1}^{L-1}\end{subarray}}\frac{1}{m}\sum_{i=1}^{m}s\sigma_{i}$
$\quad\quad\Big{<}\boldsymbol{e}_{j},\phi_{L-2}\circ\dots\circ\phi_{1}\left(\boldsymbol{z}^{(u_{i})}_{1};\boldsymbol{z}^{(u_{i})}_{2};\dots;\boldsymbol{z}^{(u_{i})}_{K^{\prime}};\boldsymbol{w}_{n_{i}}\right)\Big{>}\Bigg{]}$
$\underset{(d)}{\leq}2c_{\phi}^{L-1}\prod_{k=1}^{L}\|\boldsymbol{W}_{k}\|_{1}\mathbb{E}_{\sigma}\Bigg{[}\sup\limits_{\begin{subarray}{c}\boldsymbol{z},\boldsymbol{w}_{n},\\ j\in[(K^{\prime}+1)d]\end{subarray}}\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\Big{<}\boldsymbol{e}_{j},$
$\quad\quad\left(\boldsymbol{z}^{(u_{i})}_{1};\boldsymbol{z}^{(u_{i})}_{2};\dots;\boldsymbol{z}^{(u_{i})}_{K^{\prime}};\boldsymbol{w}_{n_{i}}\right)\Big{>}\Bigg{]}$
$\leq 2c_{\phi}^{L-1}\prod_{k=1}^{L}\|\boldsymbol{W}_{k}\|_{1}\Bigg{(}\underbrace{\mathbb{E}_{\sigma}\Big{[}\sup_{j\in[d],\boldsymbol{w}_{n}}\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\left<\boldsymbol{e}_{j},\boldsymbol{w}_{n_{i}}\right>\Big{]}}_{I_{1}}$
$\quad+\sum_{t=1}^{K^{\prime}}\ \underbrace{\mathbb{E}_{\sigma}\Big{[}\sup_{j\in[d],\boldsymbol{z}_{t}}\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\left<\boldsymbol{e}_{j},\boldsymbol{z}_{t}^{(u_{i})}\right>\Big{]}}_{I_{2}}\Bigg{)}$ (40)
where (a) use E.q. (39) as $\boldsymbol{W}_{L}\in\mathbb{R}^{1\times d_{L-1}}$ is actually a vector; (b) holds since $\sigma$ is applied element wise, we can bring $\boldsymbol{e}^{T}_{j}$ inside the function and the use of contraction inequality [67]; (c) use E.q (39) again as $\boldsymbol{e}_{j}^{\mathsf{T}}\boldsymbol{W}_{L-1}$ is still a vector; (d) holds by applying E.q. (39) recursively and utilizing the fact that ^$\hat{\mathcal{R}}_{m}(\mathcal{F}\cup-\mathcal{F})\leq 2\hat{\mathcal{R}}_{m}(\mathcal{F})$.
As term of $I_{1}$, using Cauchy-Schwartz inequality and Jensen inequality, we have:
$I_{1}$ $\leq\frac{1}{m}\mathbb{E}_{\sigma}\left[\left\|\sum_{i=1}^{m}\sigma_{i}\boldsymbol{w}_{n_{i}}\right\|_{2}\right]\leq\frac{1}{m}\left(\mathbb{E}_{\sigma}\left[\left\|\sum_{i=1}^{m}\sigma_{i}\boldsymbol{w}_{n_{i}}\right\|_{2}^{2}\right]\right)^{1/2}$
$=\frac{1}{m}\left(\sum_{i=1}^{m}\left\|\boldsymbol{w}_{n_{i}}\right\|^{2}\right)^{1/2}\leq\frac{B_{0}}{\sqrt{m}}.$ (41)
As term of $I_{2}$, we have:
$I_{2}$ $=\frac{1}{m}\mathbb{E}_{\sigma}\left[\sup_{j\in[d],\boldsymbol{w}}\sum_{i=1}^{m}\sigma_{i}\left\langle\boldsymbol{e}_{j},\sum_{k\in T_{t}}w_{k}^{(u_{i})}\boldsymbol{a}_{k}^{(u_{i})}\right\rangle\right]$
$\leq\frac{1}{m}\sum_{k\in T_{t}}\mathbb{E}_{\sigma}\left[\sup_{j\in[d],\boldsymbol{w}}\sum_{i=1}^{m}\sigma_{i}\left\langle\boldsymbol{e}_{j},w_{k}^{(u_{i})}\boldsymbol{a}_{k}^{(u_{i})}\right\rangle\right]$
$=\frac{1}{m}\sum_{k\in T_{t}}\mathbb{E}_{\sigma}\Bigg{[}\sup_{j\in[d],\boldsymbol{w}}\sum_{i=1}^{m}\sigma_{i}\phi\Big{(}\boldsymbol{W}_{w}^{(2)}\phi\Big{(}\boldsymbol{W}_{w}^{(1)}$
$\quad\quad\quad\left[\boldsymbol{a}_{k}^{(u_{i})};\boldsymbol{a}_{k}^{(u_{i})}\odot\boldsymbol{w}_{n_{i}};\boldsymbol{w}_{n_{i}}\right]\Big{)}\Big{)}\left\langle\boldsymbol{e}_{j},\boldsymbol{a}_{k}^{(u_{i})}\right\rangle\Bigg{]}$
$\leq\frac{1}{m}c_{\phi}^{2}\left\|\boldsymbol{W}_{w}^{(1)}\right\|_{1}\left\|\boldsymbol{W}_{w}^{(2)}\right\|_{1}\sum_{k\in T_{t}}\mathbb{E}_{\sigma}\Bigg{[}\sup_{\begin{subarray}{c}\boldsymbol{w},j\in[d]\\ j^{\prime}\in[3d]\end{subarray}}\sum_{i=1}^{m}\sigma_{i}$
$\quad\quad\left\langle\boldsymbol{e}_{j^{\prime}},\left[\boldsymbol{a}_{k}^{(u_{i})};\boldsymbol{a}_{k}^{(u_{i})}\odot\boldsymbol{w}_{n_{i}};\boldsymbol{w}_{n_{i}}\right]\right\rangle\left\langle\boldsymbol{e}_{j},\boldsymbol{a}_{k}^{(u_{i})}\right\rangle\Bigg{]}.$ (42)
Furthermore, we can get
$\mathbb{E}_{\sigma}\Bigg{[}\sup_{\begin{subarray}{c}\boldsymbol{w},j\in[d]\\ j^{\prime}\in[3d]\end{subarray}}\sum_{i=1}^{m}\sigma_{i}\left\langle\boldsymbol{e}_{j^{\prime}},\left[\boldsymbol{a}_{k}^{(u_{i})};\boldsymbol{a}_{k}^{(u_{i})}\odot\boldsymbol{w}_{n_{i}};\boldsymbol{w}_{n_{i}}\right]\right\rangle\left\langle\boldsymbol{e}_{j},\boldsymbol{a}_{k}^{(u_{i})}\right\rangle\Bigg{]}$ (43)
$\leq\mathbb{E}_{\sigma}\left[\sup_{j\in[d],j^{\prime}\in[d]}\sum_{i=1}^{m}\sigma_{i}\left\langle\boldsymbol{e}_{j^{\prime}},\boldsymbol{a}_{k}^{(u_{i})}\right\rangle\left\langle\boldsymbol{e}_{j},\boldsymbol{a}_{k}^{(u_{i})}\right\rangle\right]$
$\quad+\mathbb{E}_{\sigma}\left[\sup_{j\in[d],j^{\prime}\in[d],\boldsymbol{w}}\sum_{i=1}^{m}\sigma_{i}\left\langle\boldsymbol{e}_{j^{\prime}},\boldsymbol{a}_{k}^{(u_{i})}\odot\boldsymbol{w}_{n_{i}}\right\rangle\left\langle\boldsymbol{e}_{j},\boldsymbol{a}_{k}^{(u_{i})}\right\rangle\right]$
$\quad+\mathbb{E}_{\sigma}\left[\sup_{j\in[d],j^{\prime}\in[d],\boldsymbol{w}}\sum_{i=1}^{m}\sigma_{i}\left<\boldsymbol{e}_{j^{\prime}},\boldsymbol{w}_{n_{i}}\right>\left<\boldsymbol{e}_{j},\boldsymbol{a}_{k}^{(u_{i})}\right>\right]$
$=I_{3}+I_{4}+I_{5}.$
As term of $I_{3}$, notice that
$\sum_{i=1}^{m}\sigma_{i}\left<\boldsymbol{e}_{j^{\prime}},\boldsymbol{a}_{k}^{(u_{i})}\right>\left<\boldsymbol{e}_{j},\boldsymbol{a}_{k}^{(u_{i})}\right>=\sum_{i=1}^{m}\sigma_{i}\boldsymbol{e}_{j^{\prime}}^{\mathsf{T}}\boldsymbol{P}_{a}^{(u_{i})}\boldsymbol{e}_{j}$
$=\sum_{i=1}^{m}\sigma_{i}\mathrm{Tr}\left(\boldsymbol{e}_{j}\boldsymbol{e}_{j^{\prime}}^{\mathsf{T}}\boldsymbol{P}_{a}^{(u_{i})}\right)=\mathrm{Tr}\left(\boldsymbol{e}_{j}\boldsymbol{e}_{j^{\prime}}^{\mathsf{T}}\left(\sum_{i=1}^{m}\sigma_{i}\boldsymbol{P}_{a}^{(u_{i})}\right)\right)$
$=\left\langle\boldsymbol{e}_{j}\boldsymbol{e}_{j^{\prime}}^{\mathsf{T}},\sum_{i=1}^{m}\sigma_{i}\boldsymbol{P}_{a}^{(u_{i})}\right\rangle_{F},$
where $\boldsymbol{P}_{a}^{(u_{i})}=\boldsymbol{a}_{k}^{(u_{i})}{\boldsymbol{a}_{k}^{(u_{i})}}^{\mathsf{T}}$. Then, we can get
$I_{3}$ $=\mathbb{E}_{\sigma}\left[\sup_{j\in[d],j^{\prime}\in[d]}\sum_{i=1}^{m}\sigma_{i}\left\langle\boldsymbol{e}_{j^{\prime}},\boldsymbol{a}_{k}^{(u_{i})}\right\rangle\left\langle\boldsymbol{e}_{j},\boldsymbol{a}_{k}^{(i)}\right\rangle\right]$ (44)
$=\mathbb{E}_{\sigma}\left[\sup_{j\in[d],j^{\prime}\in[d]}\left\langle\boldsymbol{e}_{j}\boldsymbol{e}_{j^{\prime}}^{\mathsf{T}},\sum_{i=1}^{m}\sigma_{i}\boldsymbol{P}_{a}^{(u_{i})}\right\rangle_{F}\right]$
$\leq\mathbb{E}_{\sigma}\left[\left\|\sum_{i=1}^{m}\sigma_{i}\boldsymbol{P}_{a}^{(u_{i})}\right\|_{F}\right]=\sqrt{\sum_{i=1}^{m}\left\|\boldsymbol{P}_{a}^{(u_{i})}\right\|_{F}^{2}}\leq\sqrt{mB_{a}^{4}}.$
As terms of $I_{4}$, use the same analysis technique as for $I_{3}$, we can get
$I_{4}=\mathbb{E}_{\sigma}\left[\sup_{j\in[d],j^{\prime}\in[d],\boldsymbol{w}}\sum_{i=1}^{m}\sigma_{i}\left\langle\boldsymbol{e}_{j^{\prime}},\boldsymbol{a}_{k}^{(u_{i})}\odot\boldsymbol{w}_{n_{i}}\right\rangle\left\langle\boldsymbol{e}_{j},\boldsymbol{a}_{k}^{(u_{i})}\right\rangle\right]$ (45)
$=\mathbb{E}_{\sigma}\left[\sup_{j\in[d],j^{\prime}\in[d],\boldsymbol{w}}\sum_{i=1}^{m}\sigma_{i}\left\langle\boldsymbol{e}_{j^{\prime}}\odot\boldsymbol{w}_{n_{i}},\boldsymbol{a}_{k}^{(u_{i})}\right\rangle\left\langle\boldsymbol{e}_{j},\boldsymbol{a}_{k}^{(u_{i})}\right\rangle\right]$
$=\mathbb{E}_{\sigma}\left[\sup_{j\in[d],j^{\prime}\in[d],\boldsymbol{w}}\sum_{i=1}^{m}\sigma_{i}\left\langle\boldsymbol{e}_{j}\boldsymbol{e}_{j^{\prime}}^{\mathsf{T}}\odot\boldsymbol{w}_{n_{i}}^{\mathsf{T}},\boldsymbol{P}_{a}^{(u_{i})}\right\rangle_{F}\right]$
$=\mathbb{E}_{\sigma}\left[\sup_{j\in[d],j^{\prime}\in[d],\boldsymbol{w}}\sum_{n\in\mathcal{N}}\sum_{i:n_{i}=n}\sigma_{i}\left\langle\boldsymbol{e}_{j}\boldsymbol{e}_{j^{\prime}}^{\mathsf{T}}\odot\boldsymbol{w}_{n}^{\mathsf{T}},\boldsymbol{P}_{a}^{(u_{i})}\right\rangle_{F}\right]$
$\leq\sum_{n\in\mathcal{N}}\mathbb{E}_{\sigma}\left[\sup_{j\in[d],j^{\prime}\in[d]}\left\langle\boldsymbol{e}_{j}\boldsymbol{e}_{j^{\prime}}^{\mathsf{T}}\odot\boldsymbol{w}_{n}^{\mathsf{T}},\sum_{i:n_{i}=n}\sigma_{i}\boldsymbol{P}_{a}^{(u_{i})}\right\rangle_{F}\right]$
$\leq B_{0}\sum_{n\in\mathcal{N}}\mathbb{E}_{\sigma}\left[\left\|\sum_{i:n_{i}=n}\sigma_{i}\boldsymbol{P}_{a}^{(u_{i})}\right\|_{F}\right]$
$\leq B_{0}B_{a}^{2}\sum_{n\in\mathcal{N}}\sqrt{\left|\{i:n_{i}=n\}\right|}\leq B_{0}B_{a}^{2}\sqrt{\frac{2B-1}{B-1}Nm},$
where the last inequality holds as $\sum_{n\in\mathcal{N}}\left|\{i:n_{i}=n\}\right|=m$, $|\mathcal{N}|=\sum_{h=0}^{\lceil\log_{B}N\rceil-1}B^{h}+N\leq\frac{2B-1}{B-1}N$ for the $B$-ary tree for the $N$ item, and use the Cauchy-Schwartz inequality.
As terms of $I_{5}$, use the same technique as for $I_{4}$, we have
$I_{5}$ $=\mathbb{E}_{\sigma}\left[\sup_{j\in[d],j^{\prime}\in[d],\boldsymbol{w}}\sum_{i=1}^{m}\sigma_{i}\left<\boldsymbol{e}_{j^{\prime}},\boldsymbol{w}_{n_{i}}\right>\left<\boldsymbol{e}_{j},\boldsymbol{a}_{k}^{(u_{i})}\right>\right]$
$=\mathbb{E}_{\sigma}\left[\sup_{j\in[d],j^{\prime}\in[d],\boldsymbol{w}}\sum_{n\in\mathcal{N}}\sum_{i:n_{i}=n}\sigma_{i}\left<\boldsymbol{e}_{j^{\prime}},\boldsymbol{w}_{n_{i}}\right>\left<\boldsymbol{e}_{j},\boldsymbol{a}_{k}^{(u_{i})}\right>\right]$
$\leq\sum_{n\in\mathcal{N}}\mathbb{E}_{\sigma}\left[\sup_{j\in[d],j^{\prime}\in[d]}\left<\boldsymbol{e}_{j^{\prime}},\boldsymbol{w}_{n}\right>\left<\boldsymbol{e}_{j},\sum_{i:n_{i}=n}\sigma_{i}\boldsymbol{a}_{k}^{(u_{i})}\right>\right]$
$\leq B_{0}\sum_{n\in\mathcal{N}}\mathbb{E}_{\sigma}\left[\left\|\sum_{i:n_{i}=n}\sigma_{i}\boldsymbol{a}_{k}^{(u_{i})}\right\|_{2}\right]\leq B_{0}B_{a}\sqrt{\frac{2B-1}{B-1}Nm}.$ (46)
[…]
^$\begin{split}&\hat{\mathcal{R}}_{m}(\mathcal{F}_{Din})=\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{F}_{Din}}\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}f\left(u_{i},n_{i}\right)\right]\\ &\leq 2c_{\phi}^{L-1}\prod_{k=1}^{L}\|\boldsymbol{W}_{k}\|_{1}\Bigg{[}\frac{B_{0}}{\sqrt{m}}+c_{\phi}^{2}\left\|\boldsymbol{W}_{w}^{(1)}\right\|_{1}\left\|\boldsymbol{W}_{w}^{(2)}\right\|_{1}K\\ &\quad\left(\frac{B_{a}^{2}}{\sqrt{m}}+\frac{B_{0}B^{2}_{a}}{\sqrt{m}}\sqrt{\frac{2B-1}{B-1}N}+\frac{B_{0}B_{a}}{\sqrt{m}}\sqrt{\frac{2B-1}{B-1}N}\right)\Bigg{]}\\ &\leq\frac{2c_{\phi}^{L-1}B_{1}^{L}(B_{0}+KB_{w_{1}}+KB_{w_{2}}\mathcal{\tau})}{\sqrt{m}},\end{split}$
where $B_{w_{1}}=B_{2}^{2}c_{\phi}^{2}B^{2}_{a},B_{w_{2}}=B_{2}^{2}c_{\phi}^{2}\left(B_{0}B_{a}^{2}+B_{0}B_{a}\right),\tau=\sqrt{(2B-1)N}/\sqrt{B-1}$. ∎
A.6 Proof of Theorem 4
Proof.
Given sample set $S=\{u_{i},y_{i}\}_{i=1}^{m}$, consider the function space $\mathcal{F}_{\ell}^{j}=\{(u,y)\rightarrow f_{\ell}^{j}(u,y)\}$, where function $f_{\ell}^{j}(u,y)=\widetilde{\mathcal{L}}_{j}(u,y)=-\bar{z}^{j}_{\delta^{j}(y)}\frac{\exp o^{j}_{\delta^{j}(y)}(u)}{\sum_{k=1}^{c(j)}\exp o^{j}_{k}(u)}$ and $o_{k}^{j}(u)=f_{Din}(u,n_{k}^{j})$. Due to $\left|f_{Din}\right|\leq B_{\mathcal{F}}$, we can get $\left|f_{\ell}^{j}\right|\leq 2B_{\mathcal{F}}+\log c(j)$, so by the Lemma 4, we can get:
$\begin{split}\mathbb{E}_{(u,y)\sim\mathcal{D}}\left[\widetilde{\mathcal{L}}_{j}(u,y)\right]&\leq\frac{1}{m}\sum_{i=1}^{m}\widetilde{\mathcal{L}}_{j}(u_{i},y_{i})+2\hat{\mathcal{R}}_{m}(\mathcal{F}_{\ell}^{j},S)\\ &+\Big{(}4\log c(j)+8B_{\mathcal{F}}\Big{)}\sqrt{\frac{2\log\left(4/\delta\right)}{m}}.\end{split}$ (47)
For the empirical Rachemader complexity of $\mathcal{F}_{\ell}^{j}$, we have
^$\begin{split}&\hat{\mathcal{R}}_{m}(\mathcal{F}_{\ell}^{j},S)=\mathbb{E}_{\sigma}\left[\sup_{f_{\ell}^{j}\in\mathcal{F}_{\ell}^{j}}\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}f_{\ell}^{j}(u_{i},y_{i})\right]\\ &=\mathbb{E}_{\sigma}\left[\sup_{\begin{subarray}{c}f_{Din}\in\mathcal{F}_{Din}\\ \bar{z}\in\{0,1\}\end{subarray}}\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\bar{z}^{j}_{\delta^{j}(y_{i})}\log\left(\frac{\sum_{k=1}^{c(j)}\exp o^{j}_{k}(u_{i})}{\exp o_{\delta^{j}(y_{i})}^{j}(u_{i})}\right)\right]\\ &=\mathbb{E}_{\sigma}\left[\sup_{\begin{subarray}{c}f_{Din}\in\mathcal{F}_{Din}\\ z^{\prime}\triangleq 2\bar{z}-1\in\{-1,1\}\end{subarray}}\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\frac{z^{\prime}+1}{2}\log\left(\frac{\sum_{k=1}^{c(j)}\exp o^{j}_{k}(u_{i})}{\exp o_{\delta^{j}(y_{i})}^{j}(u_{i})}\right)\right]\\ &\leq\mathbb{E}_{\sigma}\left[\sup_{\begin{subarray}{c}f_{Din}\in\mathcal{F}_{Din}\end{subarray}}\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}*-\log\left(\frac{\exp o_{\delta^{j}(y_{i})}^{j}(u_{i})}{\sum_{k=1}^{c(j)}\exp o^{j}_{k}(u_{i})}\right)\right]\\ &=\hat{\mathcal{R}}_{m}(\ell\circ\mathcal{F}_{Din},S)\end{split}$ (48)
where the last equality holds as $-\log\left(\frac{\exp o_{\delta^{j}(y_{i})}^{j}(u_{i})}{\sum_{k=1}^{c(j)}\exp o^{j}_{k}(u_{i})}\right)$ is a composite of logistic loss $\ell_{y}(o)=-\sum_{i}y_{i}\frac{\exp o_{i}}{\sum_{j}\exp o_{j}}$ and $f_{Din}$. The logistic loss function $\ell$ is $1$ Lipschitz[68], and for the $j$-th layer, the label num is $c(j)$, by the Lemma 5, we have
^$\hat{\mathcal{R}}_{m}(\ell\circ\mathcal{F}_{Din},S)\leq 2c(j)\hat{\mathcal{R}}_{m}(\mathcal{F}_{Din},S)$ (49)
Combine the E.q. (47), E.q. (48) and E.q. (49), and utilize the Lemma 2, we can obtain the desired result. ∎
A.7 Proof of Lemma 3
Proof.
The $i$-th label is positive of $N$ labels, and the sample distribution is $Q$, so according to E.q. (17), the adjusted logit can be calculated as follows:
$\hat{o}_{i^{\prime}}=\left\{\begin{aligned} &o_{i}-\ln(Mq_{i^{\prime}})\ \ \ \ if\ i^{\prime}\neq i\\ &o_{i}-\ln(1)\quad\quad\ \ \ if\ i^{\prime}=I\end{aligned}\right.$ (50)
Then, we can have:
$\begin{split}-&\log\frac{\exp o_{i}}{\sum_{k=1}^{N}\exp o_{k}}-{\mathbb{E}}_{\mathcal{I}^{\prime}_{M}}\Big{[}-\log\frac{\exp\hat{o}_{i}}{\sum\limits_{i^{\prime}\in\mathcal{I}^{\prime}_{M}\cup\{i\}}\exp\hat{o}_{i^{\prime}}}\Big{]}\\ &=\mathbb{E}_{\mathcal{I}^{\prime}_{M}}\left[\log\Big{(}\sum_{k=1}^{N}\exp o_{k}\Big{)}-\log\left(\exp o_{i}+\sum_{i^{\prime}\in\mathcal{I}^{\prime}_{M}}\frac{\exp o_{i^{\prime}}}{Mq_{i^{\prime}}}\right)\right]\\ &=\mathbb{E}_{\mathcal{I}^{\prime}_{M}}\left[-\log\left(p_{i}+\frac{1}{M}\sum_{i^{\prime}\in\mathcal{I}^{\prime}_{M}}\frac{p_{i^{\prime}}}{q_{i^{\prime}}}\right)\right]\\ &=\mathbb{E}_{\mathcal{I}^{\prime}_{M}}\left[\log\frac{M}{\sum_{i^{\prime}\in\mathcal{I}^{\prime}_{M}}\Big{(}\frac{q_{i^{\prime}}}{p_{i^{\prime}}+p_{i}q_{i^{\prime}}}\Big{)}^{-1}}\right]\\ &\underset{(a)}{\leq}\mathbb{E}_{\mathcal{I}^{\prime}_{M}}\left[\log\left({\prod_{i^{\prime}\in\mathcal{I}^{\prime}_{M}}\frac{q_{i^{\prime}}}{p_{i^{\prime}}+p_{i}q_{i^{\prime}}}}\right)^{\frac{1}{M}}\right]\\ &=\frac{1}{M}\sum_{i^{\prime}\in\mathcal{I}^{\prime}_{M}}\mathbb{E}_{i^{\prime}\sim Q}\left[\log\frac{q_{i^{\prime}}}{p_{i^{\prime}}+p_{i}q_{i^{\prime}}}\right]\\ &\leq\frac{1}{M}\sum_{i^{\prime}\in\mathcal{I}^{\prime}_{M}}\mathbb{E}_{i^{\prime}\sim Q}\left[\log\left(\frac{q_{i^{\prime}}}{p_{i^{\prime}}}\right)\right]=D_{\text{KL}}(Q\|P),\end{split}$
where (a) uses Harmonic-Geometric inequality. ∎
Appendix B Auxiliary Lemmas
Lemma 4 (Theorem 26.5(2) of [69]).
If the magnitude of loss function $l$ is bounded above by $c$, with probability greater than $1-\delta$ for all $h\in\mathcal{H}$, we have
$\begin{split}\mathbb{E}_{(x,y)\sim\mathcal{D}}[\ell(h(x),y)]&\leq\frac{1}{m}\sum_{i=1}^{m}\ell\left(h\left(x_{i}\right),y_{i}\right)\\ &+2\hat{\mathcal{R}}_{m}(\ell\circ\mathcal{H},S)+4c\sqrt{\frac{2\ln(4/\delta)}{m}}\end{split}$
where =$\ell\circ\mathcal{H}=\{\ell(h(x),y)\mid(x,y)\in\mathcal{X}\times\mathcal{Y},h\in\mathcal{H}\}$.
Lemma 5 (Lemma 2 of [68]).
Let $\mathcal{F}$ be class of real functions and $\mathcal{H}=[\mathcal{F}_{j}]^{k}_{j=1}$ be a k-dimensional function class. If $\mathcal{A}:\mathbb{R}^{k}\rightarrow\mathbb{R}$ is a Lipschitz function with constant $L$ and satisfies $\mathcal{A}(0)=0$, then
^$\hat{\mathcal{R}}_{m}(\mathcal{A}\circ\mathcal{H})\leq 2kL\hat{\mathcal{R}}_{m}(\mathcal{F}).$
Lemma 6 (Theorem 1 of [70]).
Let $\psi:\mathbb{R}^{M}\mapsto\mathbb{R}$ be a strictly convex differentiable function, and $D_{\psi}:\mathbb{R}^{M}\times\mathbb{R}^{M}\mapsto\mathbb{R}$ is the Bregman divergence induced by $\psi$. Let $Y$ be an arbitrary random variable taking values in $\mathbb{R}^{M}$ for which both $\mathbb{E}[Y]$ and $\mathbb{E}[\psi(Y)]$ are finite, we have
$\operatornamewithlimits{argmin}\limits_{s\in\mathbb{R}^{M}}\mathbb{E}_{Y}[D_{\psi}(Y,s)]=\mathbb{E}[Y]$
Acknowledgments
This should be a simple paragraph before the References to thank those individuals and institutions who have supported your work on this article.