### v2

# Learning Deep Tree-based Retriever for Efficient Recommendation: Theory and Method

Abstract
With the advancement of deep learning techniques, deep recommendation models have achieved remarkable improvements in recommendation accuracy. However, due to the large number of candidate items in practice and the high cost of preference computation, these methods also suffer from low efficiency of recommendation. The recently proposed tree-based deep recommendation models alleviate the problem by directly learning tree structure and representations under the guidance of recommendation objectives. To guarantee the effectiveness of beam search for recommendation accuracy, these models strive to ensure that the tree adheres to the max-heap assumption, where a parent node’s preference should be the maximum among its children’s preferences. However, the binary classification objectives of these models inherently limit their ability to fully satisfy this assumption. To this end, we propose Deep Tree-based Retriever (DTR for short) for efficient recommendation. When learning the node representation of the tree, we have to satisfy the max-heap assumption as much as possible and mimic beam search behavior over the tree. DTR accomplishes this by framing the training task as a multi-class classification over tree nodes at the same level. To further enhance model performance, we propose a rectification method to mitigate the suboptimality inherent in this training mode under beam search. As the number of tree nodes grows exponentially with the levels, we employ sampled softmax to approximate optimization and thereby enhance efficiency. Furthermore, we propose a tree-based sampling method to reduce the bias inherent in sampled softmax. The experiments are conducted on four real-world datasets, validating the effectiveness of the proposed preference model learning method and tree learning method.
Index Terms:
Recommender system; Tree-based Index; Multi-class Classification; Sampled Softmax; Efficient Recommendation
1 Introduction
In the information era, the overwhelming volume of daily information leads to significant information overload. The recommendation plays a crucial role in mitigating information overload by providing a personalized ranking list tailored to individual preferences. With the advancements of deep learning techniques, recommendation techniques also achieve remarkable improvements in ranking performance. The widespread application of these techniques has generated considerable economic benefits for various kinds of content providers in industrial companies.
Through the use of deep learning, we not only learn better representations for users, items, and contexts but also provide a more generalized expression for users’ preference scores via neural networks than the widely-used inner product in matrix factorization. Both lead to stronger recommendation performance, as demonstrated by models such as DIN [1], DIEN [2], NCF [3], CDL [4], and CKE [5]. However, the use of neural network-based preference functions brings online-serving challenges for recommender systems due to the high cost of preference computation. Generally speaking, immediate responses to adaptive recommendations are prerequisites for excellent customer experiences and custom retention. Existing popular and well-performed graph-based indexes, e.g. HNSW [6], quantization-based indexes, e.g. PQ [7], AQ [8] and SCANN [9], and Hash-based indexes, e.g. SGDH [10] are typically constructed based on inner product or Euclidean distance, such that they are not suitable for speeding up the recommendation of deep models. In particular, neural network-based preference functions do not form a valid metric and are incompatible with Euclidean distance and inner product. As a result, two items that are close in Euclidean or inner-product space may have significantly different neural preference scores.
To guarantee the compatibility between search indexes with the neural network-based preference functions, it is a good solution to learn the search indexes together with the recommendation model under the guidance of recommendation objectives. The representative work is the tree-based deep model (i.e. TDM [11] and its improved version JTM [12]). These models use the balanced tree index, which is constructed by hierarchically clustering item representations from top to down. Given the tree, the top-k ranked items are obtained by layer-wise beam search, which selects the k-largest tree nodes based on neural preference scores in each level from top to down. In this way, beam search achieves logarithmic computation complexity w.r.t. the number of items. To guarantee the accuracy of beam search, these tree-based deep models rely on the following max-heap assumption: the preference scores of query for a parent node should be the maximum between the preference scores of its children node. For the sake of satisfying the assumption, these tree-based models cast the overall problem as a binary classification problem by treating nodes in the path from the root to positive samples as positive and randomly sampled nodes as negative. However, these tree models suffer from the following drawback: the max-heap assumption is not well satisfied by the used binary classification objectives due to the lack of explicit horizontal competition among all tree nodes at the same level. This drawback provides an opportunity to improve the accuracy of the efficient recommendation.
We proposed Deep Tree-based Retriever (DTR for short) for the efficient recommendation. To satisfy the max-heap assumption as much as possible, DTR regards the training task as multi-class classification over tree nodes at the same layer, which enables horizontal competition among them and mimics beam search behavior in the training stage. Within this multi-class classification training mode, DTR utilizes a multi-class cross-entropy loss to optimize the deep model; however, theoretical analysis from the aspect of Bayes optimality indicates that this loss function still results in suboptimal performance under beam search, prompting us to propose a rectification method. Additionally, as the number of tree nodes increases exponentially with levels, the softmax loss becomes computationally expensive. Therefore, we resort to sampled softmax for approximation to promote training efficiency and propose a tree-based negative sampling method guided by the sampled softmax theory to estimate the gradient of the original softmax loss more accurately. Moreover, we propose a tree learning method compatible with this training mode, enabling the alternate learning of the preference model and the tree. Furthermore, we provide a generalization analysis of our proposed method, showing that it has good generalization capabilities and that the generalization could be enhanced through an appropriate negative sampling distribution and an increased number of branches in the tree structure.
As an extension of the preliminary paper [13], which contributed to the layer-wise multi-class classification training mode and a tree learning method compatible with this training mode, we further make the following contributions:
  • We identify the issue of suboptimality of multi-class classification training mode under beam search from the aspect of Bayes optimality and propose a rectification method to address this issue.
  • We propose a tree-based negative sampling method for the multi-class classification problem in the tree. This sampling method can lead to more accurate estimations of loss gradient to reduce the inherent bias in sampled softmax.
  • We provide a generalization analysis of the deep tree-based retriever, demonstrating the great generalization capability of our proposed methods.
  • We evaluate the proposed methods on four real-world datasets and validate the superiority of DTR to the baselines and the effectiveness of the rectification method and the sampling method.
We organize the rest of the paper as follows. Section 2 provides an overview of related work. Section 3 presents notations and formalizes the tree-based model. Our proposed method is mainly presented in Section 4 and Section 5; the former introduces the multi-class classification training mode and the rectification method to mitigate suboptimality, while the latter details the proposed tree-based negative sampling method. Subsequently, Section 6 introduces how to update the tree index, and Section 7 provides a generalization analysis of the proposed method. The experiment details and results are presented in Section 8, and the conclusion of this paper is provided in Section 9. Finally, the missing proofs are included in the appendix.
2 Related work
This study aims to enhance the accuracy of efficient recommendation through the deep tree-based retriever, focusing on both the methodologies and theoretical foundations. We begin by reviewing recent advancements in efficient recommendation. Subsequently, we survey recent significant techniques that are highly beneficial in developing such recommender systems and then delve into the theories closely connected to our study.
2.1 Efficient Recommendation
Efficient recommendation relies on building a search index, including LSH [14], inverted index [15, 16], tree index [17, 18] and graph index [6] for all items. The recommender system usually uses the inner product for computing preference scores, and the top-k recommendation can be cast into the maximum inner product search (MIPS) problem. The search index is usually constructed based on Euclidean distance and has been extended to the inner product. This extension can be achieved by establishing the relationship between nearest neighbor search and MIPS [19, 20, 21], or learning from either item representations [22, 23, 24, 25] or the raw data directly [26, 27, 28, 29]. With the introduction of deep learning into the recommender systems, the preference score function becomes complicated, such that it is challenging to transform from neural ranking to NNS or MIPS. Existing work either directly used metric-based index[30], or learns search index from raw data directly together with recommendation models [11, 12, 31].
2.2 Methods for Developing Recommender Systems
2.2.1 Negative Sampling in RecSys
Negative sampling is an important method to address the negative missing and exposure bias problems and to speed up the convergence of recommender training  [32, 33, 34]. It includes static sampling, such as uniform sampling [35], popularity sampling, and adaptive sampling. The adaptive sampling is context-dependent, whose representation work includes adaptive oversampling [35], rejection sampling [36, 37], clustering-based sampling [32], and dynamic negative sampling [38]. The core idea of these adaptive samplers is that items with larger preference scores should be sampled with higher probability.
2.2.2 Techniques of Speedup Softmax Computation
In natural language applications, it is very computationally expensive to represent an output distribution over the choice of a word based on the softmax function. To address the efficiency, many approximate algorithms were proposed. For example, hierarchical softmax [39] and lightRNN [40] decomposed the probabilities, and Contrastive Divergence [41] approximated the gradient-based on MCMC. As an alternative, negative sampling is also widely used in reducing the computational cost of training the models. The representative work includes Noise-Contrastive Estimation [42] with the unigram distribution as a sampler, Generative Adversarial Networks [43, 44] with a neural-network empowered sampler, Self-Contrast Estimator [45] by the model in the immediately preceding epoch as a sampler, self-adversarial negative sampling [46] and Kernel-based sampling [47] with the tree index.
2.3 Theoretical Work
2.3.1 Bayes Optimality in Multi-class Classification and Hierarchical Classification
Bayes optimality aims for the learned classifier to achieve the highest possible performance on previously unseen data distributions. In the realm of multi-class classification, Bayes optimality has been extensively studied [48, 49]. Top-$k$ Bayes optimality is introduced to address ambiguities in ground truth classes by focusing on the $k$ most likely predicted classes [50, 51, 52]. In recent years, researchers have begun to extend Bayes optimality to hierarchical classification. [53] utilizes it to measure the performance of probability estimation using tree models in hierarchical classification. [54] examines performance degradation of tree models under beam search in hierarchical classification and proposes a method for achieving Bayes optimality under beam search.
2.3.2 Generalization Bounds for Multi-class Classification
Generalization error bounds are used to provide guarantees for the performance of the learned classifier. [55] provides data-dependent generalization bounds for multi-class classification based on margins. [56] further developed these bounds with multiple kernels. To accommodate the scenario with a large number of classes, data-dependent generalization error bounds with a mild dependency on the number of classes have also been studied [57]. The aforementioned bounds primarily focus on flat structures. In the context of hierarchical structures, [58] proposed a multi-class, hierarchical data-dependent generalization error bound for kernel classifiers used in large-scale taxonomies.
3 Preliminaries
3.1 Problem Definition and Notation
Consider a user space $\mathcal{U}$ and an item set $\mathcal{Y}$, and assume that the data is generated i.i.d. from some distribution $\mathbb{P}$ over $\mathcal{U}\times\mathcal{Y}$. An instance is represented as $(u,y)\in\mathcal{U}\times\mathcal{Y}$, with the corresponding joint probability $p(u,y)$. For a given user $u$, the conditional probability of observing item $y$ is denoted by $\eta_{y}(u)=p(y|u)$. Given the user $u$, the objective of the recommendation is to return the top-$k$ items with the highest conditional probabilities $\eta_{y}(u)$.
In this paper, we use the following notational conventions: bold lowercase and uppercase letters for vectors and matrices respectively, such as $\boldsymbol{a}$ and $\boldsymbol{A}$, and non-bold letters for scalars or constants, such as $k$ and $C$. For vectors, $\boldsymbol{a}_{i}$ denotes its $i$-th component and $\|\boldsymbol{a}\|_{p}$ deontes the $\ell_{p}$ norm. For matrix, $\|\boldsymbol{A}\|_{p}=\sup\limits_{\|\boldsymbol{x}\|_{p}=1}\|\boldsymbol{A}\boldsymbol{x}\|_{p}$ denotes matrix norms induced by vector $p$-norms. We denote the set $\{1,2,\dots,m\}$ by $\left[m\right]$ for any natural number $m$. Other notations with specific meanings will be explicitly defined or explained in the corresponding section.
3.2 Tree Models
Modern recommender systems need to retrieve top-$k$ items from a large-scale corpus (i.e., $\left|\mathcal{Y}\right|$ ranges from millions to billions or even tens of billions). To tackle the top-$k$ retrieval for such a large corpus efficiently and effectively, TDM  [11] and JTM [12] develop a max-heap-like tree index where the user’s preference to a node should be proportional to the largest one of its children’s preference scores ideally, which is compatible with any advanced preference model. In this subsection, we will formalize the tree index and the preference model, introduce the top-$k$ retrieval process in detail, and subsequently elaborate on the training process of the tree model.
3.2.1 Tree Index
The tree index is constructed over the item set to enable efficient recommendation. Once items are mapped to leaf nodes, item recommendation is transformed to retrieve corresponding leaf nodes which can be performed efficiently within the tree structure. The $B$-ary tree $\mathcal{T}$ is used to form the tree index. $\mathcal{T}$ consists of a set of nodes $\mathcal{N}$, with the nodes at the $j$-th level denoted by $\mathcal{N}^{j}$. The entire node set $\mathcal{N}=\bigcup_{j=0}^{H}\mathcal{N}^{j}$, where the tree height $H=\lceil\log_{B}|\mathcal{Y}|\rceil$ to accommodate all items.
For each level $j$, we denote the number of nodes in that level as $c(j)=\left|\mathcal{N}^{j}\right|$ and assume the nodes are listed in a fixed order, denoted as $(n_{1}^{j},n_{2}^{j},\dots,n_{c(j)}^{j})$. For the $i$-th node in the $j$-th level, denoted as $n_{i}^{j}$, let $\rho^{l}(n_{i}^{j})\in\mathcal{N}^{l}$ represent its ancestor in level $l$ ($0\leq l\leq j$), where $\rho^{j}(n_{i}^{j})$ refers to the node itself. The set $\mathcal{D}(n_{i}^{j})\subseteq\mathcal{N}$ represents child nodes of $n_{i}^{j}$ in the $(j\text{+}1)$-th level, while $\mathfrak{L}(n_{i}^{j})\subseteq\mathcal{N}^{H}$ denotes the leaf nodes within the subtree rooted at $n_{i}^{j}$. We introduce the function $\delta$ to return the index of a node within its level, i.e., $\delta(n_{i}^{j})=i$.
The mapping $\pi:\mathcal{N}^{H}\rightarrow\mathcal{Y}$ is introduced to map each leaf node of the tree to a unique item, with the inverse mapping denoted by $\varphi:\mathcal{Y}\rightarrow\mathcal{N}^{H}$. An item $y$ is mapped to a leaf and corresponds to a path in the tree from the root to the leaf. The $j$-th node in this path is $\rho^{j}(\varphi(y))$ and its index in this level is denoted as $\delta^{j}(y)=\delta(\rho^{j}(\varphi(y)))$. Given the user $u\in\mathcal{U}$, we attach a label $z_{i}^{j}(u)$ to each node $n_{i}^{j}$. The assignment of $z_{i}^{j}(u)$ is based on the observed instance $(u,y)$ and will be elaborated in later sections. For simplicity and unless otherwise specified, we will omit $u$ and use the notation $z_{i}^{j}$ in the rest of the paper.
3.2.2 Preference Model
The retrieval process along the tree is guided by users’ preferences for tree nodes. Since the real preference of user $u$ for node $n$ is unavailable, a preference model $\mathcal{M}:\mathcal{U}\times\mathcal{N}\rightarrow\mathbb{R}$ is utilized to predict the user’s preference scores for tree nodes. In the tree-based model, a variation of DIN [1] is employed as the preference model. In this subsection, we provide a detailed formalization of the DIN model.
In the DIN, the historical interaction sequence of a user $u$ serves as the input. Specifically, the model takes as input a matrix of item embedding vectors, corresponding to the items with which user $u$ has interacted, denoted as $\boldsymbol{A}^{(u)}$:
$\boldsymbol{A}^{(u)}=\left[\boldsymbol{a}_{1}^{(u)},\boldsymbol{a}_{2}^{(u)},\dots,\boldsymbol{a}_{K}^{(u)}\right]\in\mathbb{R}^{d\times K},$ (1)
where the last $K$ items of the historical interaction sequence are considered, and $\boldsymbol{a}_{k}^{(u)}$ represents the embedding of $k$-th item. The model employs a two-layer fully connected network to compute the weight between the node $n$ and the $k$-th interacted item, which can be expressed as:
$w_{k}^{(u)}=\phi\left(\boldsymbol{W}_{w}^{(2)}\phi\left(\boldsymbol{W}_{w}^{(1)}\left[\boldsymbol{a}_{k}^{(u)};\boldsymbol{a}_{k}^{(u)}\odot\boldsymbol{w}_{n};\boldsymbol{w}_{n}\right]\right)\right)\in\mathbb{R},$ (2)
where $\boldsymbol{W}_{w}^{(1)}\in\mathbb{R}^{h\times 3d},\boldsymbol{W}_{w}^{(2)}\in\mathbb{R}^{1\times h}$, $\boldsymbol{w}_{n}$ is the embedding of the node $n$, and $\phi$ is the activation function, which is $\operatorname{PReLU}$ in the practical model setting.
To enable a dynamic representation of user interests, the user interaction sequence is partitioned into $K^{\prime}$ time windows, where the $t$-th time window $T_{t}$ of length $k$ is a set of continuous indices, i.e., $T_{t}=\{i_{0},i_{0}+1,\dots,i_{0}+k\}$. These time windows are mutually exclusive and their union is exactly $\left[K^{\prime}\right]$. In the $t$-th time window, the interacted item embeddings will be aggregated using weights calculated in E.q. (2), as shown below:
$\boldsymbol{z}_{t}^{(u)}=\sum_{k\in T_{t}}w_{k}^{(u)}\boldsymbol{a}_{k}^{(u)}\in\mathbb{R}^{d}.$ (3)
Then, the concatenation of $K^{\prime}$ aggregated embeddings and the embedding of node $n_{i}^{j}$ will be fed into an MLP with $L$ hidden layers to output the user’s preference score for node $n_{i}^{j}$, which can be expressed as follows:
$f_{Din}(u,n_{i}^{j})=\boldsymbol{W}_{L}\cdot\phi_{L-1}\circ\dots\circ\phi_{1}\Big{(}\big{[}\boldsymbol{z}_{1}^{(u)};\boldsymbol{z}_{2}^{(u)};\dots;\boldsymbol{z}_{K^{\prime}}^{(u)};\boldsymbol{w}_{n_{i}^{j}}\big{]}\Big{)}$ (4)
where $\boldsymbol{W}_{L}\in\mathbb{R}^{1\times d_{L-1}}$, and the function $\phi_{k}(\boldsymbol{x})$ is defined:
$\phi_{k}(\boldsymbol{x})=\phi(\boldsymbol{W}_{k}\boldsymbol{x})\in\mathbb{R}^{d_{k}\times 1},\forall k\in[L-1].$ (5)
The element-wise activation function $\phi$ is Lipshitz continuous with a Lipshitz constant $c_{\phi}$, has the property $\sigma(0)=0$ and $\boldsymbol{W}_{k}\in\mathbb{R}^{d_{k}\times d_{k-1}}$ represents the weight matrix.
The preference model $\mathcal{M}$ is exactly the function space of $f_{Din}$. Given a user $u$, the preference score for the $i$-th node in the $j$-th level is denoted as $o_{i}^{j}(u)=f_{Din}(u,n_{i}^{j})$. For simplicity and unless otherwise specified, the user $u$ is omitted and the notation $o_{i}^{j}$ is adopted for the remainder of this paper.
3.2.3 Top-k Retrieval Process
As each item is mapped to a leaf node of the tree, the top-$k$ retrieval in recommendation is transformed to retrieve the $k$ leaf nodes with the largest preference scores for the given user $u$. If the user $u$’s real preference for any tree node is available, we can retrieve the top-$k$ leaf nodes by only searching all the top-$k$ nodes according to the preference at each level, proceeding from the top layer to the bottom layer. In the case where the tree is exactly a max-heap, this retrieval process ensures that the top-$k$ leaf nodes are retrieved finally because the top-$k$ nodes with the largest preference scores in the $j\text{+}1$-th layer must belong to the children of top-$k$ nodes with the largest preference scores in the $j$-th layer. Furthermore, the time complexity of the above retrieval process is logarithmic w.r.t. the corpus size, as only a constant number of nodes are searched at each layer, and the number of layers scales logarithmically with the corpus size.
The layer-wise retrieval process is presented in Algorithm 1, where $p(n|u)$ is the user $u$’s preference for node $n$ computed by the preference model $\mathcal{M}$. Essentially, this process is a layer-wise beam search with beam size $k$ along the tree $\mathcal{T}$, guided by the preference $o_{i}^{j}$ between the user $u$ and the node $n_{i}^{j}$, and can be formulated as follows:
$\mathcal{B}^{j}(u)\in\operatornamewithlimits{argTopk}\limits_{i\in\widetilde{\mathcal{B}}^{j}(u)}\ o_{i}^{j},$ (6)
where $\mathcal{B}^{j}(u)$ denotes the set of indices of selected nodes in the $j$-th level for the user $u$, and ~$\widetilde{\mathcal{B}}^{j}(u)=\left\{\delta(n)\mid n\in\bigcup_{i\in\mathcal{B}^{j-1}(u)}\mathcal{D}(n^{j-1}_{i})\right\}$ represents the index set of nodes to be expanded in the $j$-th level. It’s important to note that E.q. (6) employs the symbol $\in$” instead of “$=$” because ties in probabilities may occur, leading to multiple potential sets of top-$k$ nodes. If |~$\left|\widetilde{\mathcal{B}}^{j}(u)\right|<k$, then all nodes in $j$-th level will be selected. By recursively applying E.q. (6) until the $H$-th level, the top-$k$ leaf nodes $\mathcal{B}^{H}(u)$ is retrieved, and the final retrieved item set for the user $u$ is
$\widehat{\mathcal{Y}}(u)=\{\pi(n_{i}^{H})\mid i\in\mathcal{B}^{H}(u)\}.$ (7)
Input: User $u$, tree $\mathcal{T}$, beam size $k$, the preference model $\mathcal{M}$
Output: top-$k$ items
1:  Result set $A=\emptyset$, candidate set $Q=\{root\ node\ n_{1}^{0}\}$;
2:  while Q is not empty do
3:     Remove all the leaf nodes from $Q$ and add these leaf nodes into result set $A$ if Q contains leaf nodes;
4:     Compute preference $p(n|u)$ through $\mathcal{M}$ for each left node in set $Q$;
5:     Parents={the top-$k$ nodes of Q according to $p(n|u)$};
6:     Q={children of node $n\ |\ n$in Parents};
7:  end while
8:  return The top-$k$ items w.r.t. the top-$k$ leaf nodes according to $p(n|u)$, $n\in A$.
Algorithm 1 Beam Search: Layer-wise Retrieval
3.2.4 Learn the preference model and the tree
To improve the retrieval performance of the tree model, we need to learn the preference model (i.e. the neural network model to output the scores between users and nodes) and the tree (i.e. the mapping $\pi$ between items and the leaf nodes of the tree). The preference model and the tree can be learned alternately. Concretely, fix the mapping $\pi$ when learning the preference model and fix the preference model when learning the mapping $\pi$. This alternating learning process is repeated until both components converge and the tree model achieves optimal performance.
To learn the preference model, TDM and JTM regard the training task as a binary classification problem essentially. Supposing user $u$ has an interaction with an item that corresponds to a leaf node $n_{leaf}$, $n_{leaf}$ and its ancestors are positive nodes with label 1 but all the left nodes are negative nodes with label 0. Then the training loss for user $u$ is
$\mathcal{L}(u,S_{u}^{+},S_{u}^{-})=$ $-\sum_{n\in S_{u}^{+}}z_{u}(n)\log p(\hat{z}_{u}(n)=1|u)$ (8)
$-\sum_{n\in S_{u}^{-}}(1-z_{u}(n))\log p(\hat{z}_{u}(n)=0|u).$
$S_{u}^{+}$ and $S_{u}^{-}$ denote the positive node set and negative node set for user $u$, respectively. $z_{u}(n)$ denotes the label of node $n$ for user $u$. $p(\hat{z}_{u}(n)=1|u)$ and $p(\hat{z}_{u}(n)=0|u)$ denote the like and dislike probabilities for user $u$ to node $n$, which are exactly corresponding to the two outputs of preference model. Since using all negative nodes to train the model is unacceptable both in time-consuming and memory consumption, both TDM and JTM sample a set of negative nodes (denoted as $S_{u}^{{}^{\prime}-}$) uniformly from all the negative nodes. Then $\mathcal{L}(u,S_{u}^{+},S_{u}^{{}^{\prime}-})$ is used as training loss.
Treating the training task as a binary classification problem leads to three shortcomings: 1) The max-heap assumption is not well satisfied due to the lack of explicit horizontal competition among all tree nodes at the same level. 2) The retrieval is conducted in a layer-wise manner (e.g. Algorithm 1) but the training process doesn’t include layer-wise mode which can lead to a gap between training and prediction. 3) There are much more negative nodes than positive nodes which can make the negative nodes dominate the training loss. To overcome these shortcomings, we regard the preference model learning task as a multi-class classification problem which will be elaborated in later sections.
To update the tree structure, both TDM and JTM adopt a hierarchical clustering-like method. TDM clusters all the items recursively by k-means using the item embedding until each cluster only contains one item. The recursive clustering process can form a tree structure and the item in each final cluster is exactly mapped to a leaf node. JTM assigns all the items to each node from the top layer to the bottom layer so that the sum of the log-likelihoods for all users on each layer can be maximized. Interested readers can refer to the content of TDM  [11] and JTM  [12] for more details about how to learn the trees respectively. In our work, we propose a new tree learning method that is compatible with our proposed preference model learning mode.
4 Multi-class classification training mode
To address the shortcomings of the training mode in TDM and JTM, we develop a layer-wise training mode in this section. We regard the training task as a multi-class classification problem at each layer, employing the classic multi-class cross-entropy loss to optimize the preference model. Upon examining the Bayes optimality of this training scheme, we found that the traditional multi-class cross-entropy loss leads to suboptimality under beam search. To mitigate this issue, we introduce a rectification method to enhance the tree model’s retrieval performance under beam search.
4.1 Multi-class cross-entropy loss
TDM and JTM regard the training task as a binary classification problem and binary cross-entropy is used as the training loss. In this way, each node contributes to the training loss individually. However, the layer-wise retrieval is conducted from top layer to bottom layer and we need to compare the candidates with each other node at the same layer. There is a gap between training and prediction in TDM and JTM. To address this, we propose the layer-wise training mode which regards the training task as a multi-class classification problem.
In the context of this training mode, at layer $j$, which consists of $c(j)$ nodes, each node corresponds to a class, and multi-class classification with $c(j)$ classes is performed at that layer. Given item $y$, the positive node can be identified by backtracking from the item’s mapping leaf node to the root node, as demonstrated in TMD and JTM, and the corresponding label will be assigned as $1$, i.e., $z^{j}_{\delta^{j}(y)}=1$. The remaining $c(j)-1$ nodes are the negative nodes with the label 00. The preference model outputs a score for each node, resulting in $c(j)$ preference scores for the $j$-th layer, which is well-suited to the multi-class classification task with $c(j)$ classes. For the user $u$, $p_{i}^{j}$ denotes the softmax probability that the $i$-th node is positive node at layer $j$, where
$p_{i}^{j}=\frac{\exp{o_{i}^{j}}}{\sum_{k=1}^{c(j)}\exp{o_{k}^{j}}}.$ (9)
Given the instance $(u,y)$, the multi-class cross entropy is used as the training loss at layer $j$, i.e.,
$\mathcal{L}_{j}(u,y)=-\log p^{j}_{\delta^{j}(y)}=-\log\frac{\exp o_{\delta^{j}(y)}^{j}}{\sum_{k=1}^{c(j)}\exp o_{k}^{j}}.$ (10)
From E.q. (9) and E.q. (10), we can know that all the nodes contribute to the training loss simultaneously at layer $j$ which may relieve the gap between training and prediction. The training loss w.r.t. the whole tree can be written as
$\mathcal{L}(u,y)=\sum_{j=1}^{H}\mathcal{L}_{j}(u,y).$ (11)
4.2 Suboptimality of multi-class cross-entropy loss under beam search
Treating the training task as a multi-class classification problem at each layer can alleviate some of the shortcomings associated with binary classification. However, the use of multi-class cross-entropy loss still leads to suboptimality under beam search. In this subsection, we will demonstrate how multi-class cross-entropy loss results in suboptimal outcomes when combined with beam search.
Similar to Bayes optimality defined in [52, 54], we define the top-$k$ retrieval Bayes optimality under beam search to suit our hierarchical multi-class classification setting :
Definition 1 (Top-$k$ Retrieval Bayes Optimal).
Given the beam size $k$, the tree model consist of a tree $\mathcal{T}$ and a preference model $\mathcal{M}$ is top-$k$ retrieval Bayes optimal, if the following equation
$\widehat{\mathcal{Y}}(u)\in\operatornamewithlimits{argTopk}\limits_{y\in\mathcal{Y}}\eta_{y}(u)$ (12)
holds for any $u\in\mathcal{U}$.
The underlying idea of Definition 12 is straightforward: for each user $u$, a top-$k$ retrieval Bayes optimal tree model should return the $k$ items with the highest conditional probabilities.
Building on the Definition 12, we show that optimizing a tree model using multi-class cross-entropy loss can result in suboptimal outcomes under beam search. This demonstration leverages the concept of Bregman divergence and the fact the multi-class cross-entropy loss is a Bregman divergence. Let $M$ denote the number of classes, for a strictly convex and differentiable function $\psi:\mathbb{R}^{M}\to\mathbb{R}$, the Bregman divergence $D_{\psi}:\mathbb{R}^{M}\times\mathbb{R}^{M}\to\mathbb{R}$ induced by $\psi$ is defined as follows:
$D_{\psi}(\boldsymbol{z},\boldsymbol{o})=\psi(\boldsymbol{z})-\psi(\boldsymbol{o})-\nabla\psi(\boldsymbol{o})^{T}(\boldsymbol{z}-\boldsymbol{o}).$ (13)
By taking $\psi(\boldsymbol{s})=\sum_{m=1}^{M}\boldsymbol{s}_{m}\log\boldsymbol{s}_{m}$ and $g(\boldsymbol{o})_{m}=\exp\boldsymbol{o}_{m}/\sum_{k=1}^{M}\exp\boldsymbol{o}_{k}$, the multi-class cross-entropy loss
$\mathcal{L}(\boldsymbol{e}^{(i)},\boldsymbol{o})=-\log\left(\frac{\exp\boldsymbol{o}_{i}}{\sum_{m=1}^{M}\exp\boldsymbol{o}_{m}}\right)$ (14)
can be expressed as the form of Bregman divergence:
$\mathcal{L}(\boldsymbol{e}^{(i)},\boldsymbol{o})=D_{\psi}(\boldsymbol{e}^{(i)},g(\boldsymbol{o})).$ (15)
Here, $\boldsymbol{z},\boldsymbol{o},\boldsymbol{e}^{(i)}\in\mathbb{R}^{M}$, and the vector $\boldsymbol{e}^{(i)}$ is a one-hot vector where the $i$-th component is $1$ and other components are 00.
To make the demonstration more concise and clear, we define the concept of rank consistent as follows:
Definition 2 (Rank Consistent).
Given $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{M}$, we say that $\boldsymbol{x}$ is rank consistent w.r.t. $\boldsymbol{y}$, denoted $R(\boldsymbol{x},\boldsymbol{y})$, if and only if for all $i,j\in\left[M\right]$,
$(\boldsymbol{x}_{i}>\boldsymbol{x}_{j}\Rightarrow\boldsymbol{y}_{i}>\boldsymbol{y}_{j})\wedge(\boldsymbol{y}_{i}>\boldsymbol{y}_{j}\Rightarrow\boldsymbol{x}_{i}>\boldsymbol{x}_{j}).$
Given the user $u$, the conditional risk of multi-class cross-entropy loss in layer $j$ is
$\mathcal{R}_{j}(u)=\mathbb{E}_{y\sim\eta(u)}\left[\mathcal{L}_{j}(u,y)\right]=\mathbb{E}_{y\sim\eta(u)}\left[-\log\frac{\exp o_{\delta^{j}(y)}^{j}}{\sum_{k=1}^{c(j)}\exp o_{k}^{j}}\right].$ (16)
As previously mentioned, given the user $u$, a label $z_{i}^{j}$ is attached for the node $n_{i}^{j}$. In fact, $z_{i}^{j}$ can be viewed as a conditional random variable and the $j$-th layer’s variables composite the random vector $\boldsymbol{z}^{j}$. Let $\boldsymbol{o}^{j}$ be the preference score vector for the $j$-th layer, whose $i$-th component is exactly the user’s preference score $o^{j}_{i}$ for node $n_{i}^{j}$. With the fact that multi-class cross-entropy loss is a Bregman divergence and the softmax function $g$ satisfies $R(\boldsymbol{s},g(\boldsymbol{s}))$ for any $\boldsymbol{s}\in\mathbb{R}^{M}$, previous literature [52] has proven that optimizing the multi-class cross-entropy loss (i.e., E.q. (16)) would lead to each element in $\boldsymbol{o}^{j}$ matches the rank of the corresponding element in $\mathbb{E}\left[\boldsymbol{z}^{j}\right]$(i.e., $R(\boldsymbol{o}^{j},\mathbb{E}[\boldsymbol{z}^{j}])$). We state their theorem and make some modifications to be compatible with our problem:
Theorem 1 (Theorem 3.1 of [52]).
Given a convex, differentiable function $\psi:\mathbb{R}^{M}\mapsto\mathbb{R}$. Let $\boldsymbol{z}$ be an arbitrary random variable taking values in $\mathbb{R}^{M}$ for which both $\mathbb{E}[\boldsymbol{z}]$ and $\mathbb{E}[\psi(\boldsymbol{z})]$ are finite. If continuous function $g:\mathbb{R}^{M}\mapsto\mathbb{R}^{M}$ satisfies that $R(\boldsymbol{s},g(\boldsymbol{s}))$ holds for $\forall\boldsymbol{s}\in\operatorname{domain}(g)$, and $\mathbb{E}[\boldsymbol{z}]\subseteq\operatorname{range}(g)$, then
$\operatornamewithlimits{argmin}\limits_{\boldsymbol{s}\in\mathbb{R}^{M}}\mathbb{E}_{\boldsymbol{z}}\left[D_{\psi}(\boldsymbol{z},g(\boldsymbol{s})\right]\subseteq\left\{\boldsymbol{s}\in\mathbb{R}^{M}\mid R(\boldsymbol{s},\mathbb{E}[\boldsymbol{s}])\right\}.$
In the multi-class training mode, for a given instance $(u,y)$, we set the label $z_{\delta^{j}(y)}^{j}=1$, and other labels are set to 00 in the $j$-th layer. Then the expectation of $\boldsymbol{z}^{j}$ is as follows:
$\mathbb{E}[\boldsymbol{z}^{j}]=\left(\sum\limits_{n\in\mathfrak{L}(n_{1}^{j})}\eta_{\pi(n)},\sum\limits_{n\in\mathfrak{L}(n_{2}^{j})}\eta_{\pi(n)},\ \cdots,\sum\limits_{n\in\mathfrak{L}\left(n_{c(j)}^{j}\right)}\eta_{\pi(n)}\right).$ (17)
Notice that optimizing the multi-class cross-entropy loss (i.e. E.q. (16)) will lead to $R(\boldsymbol{o}^{j},\mathbb{E}[\boldsymbol{z}^{j}])$ from layer $1$ to $H$, we can identify that it is sub-optimal under beam search.
Proposition 1.
The multi-class cross-entropy loss will result in the tree model not being top-$k$ retrieval Bayes optimal.
Proof.
Since optimizing the multi-class cross-entropy loss leads to $R(\boldsymbol{o}^{j},\mathbb{E}[\boldsymbol{z}^{j}])$ for $1\leq j\leq H$, when performing the beam search along the tree, the result of $\operatornamewithlimits{argTopk}\limits_{i\in[c(j)]}\boldsymbol{o}^{j}_{i}$ is identical to the result of $\operatornamewithlimits{argTopk}\limits_{i\in[c(j)]}\mathbb{E}[\boldsymbol{z}^{j}_{i}]$.
[…]
4.3 Rectification under beam search
In this subsection, we introduce how to mitigate the above suboptimal results under beam search, which is achieved by modifying the original label assignment for the non-leaf nodes and the loss function. Besides, we provide the theoretical support for these modifications and propose effective solutions for their practical implementation.
4.3.1 Label Rectification
For a fixed user $u$, do not directly assign labels of the ancestors of the leaf node $\varphi(y)$ to 1 for an observed instance $(u,y)$. Only if the item $y$ is the one with the largest conditional probability among the items mapped from $\mathfrak{L}(n_{i}^{j})$, can the label $z_{i}^{j}$ of node $n_{i}^{j}$ be assigned as $1$, making the tree become the max-heap from the aspect of expectation. Denote the rectified label as $\bar{z}^{j}_{i}$, the formalization of assignment is as follows:
$\bar{z}_{i}^{j}(u)=\begin{cases}1,&n_{i}^{j}=\rho^{j}\left(\varphi(y)\right)\wedge y=\operatornamewithlimits{argmax}\limits_{y^{\prime}\in\pi\left(\mathfrak{L}(n_{i}^{j})\right)}\eta_{y^{\prime}}\\ 0,&\text{else}.\end{cases}$
With such a rectified assignment of the label of nodes, for each node $n_{i}^{j}$, we can calculate the expectation of the corresponding random variable:
$\mathbb{E}_{y\sim\eta(u)}[\bar{z}_{i}^{j}(u)]=\sum_{y\in\mathcal{Y}}\bar{z}_{i}^{j}(u)*\eta_{y}(u)=\max\limits_{n\in\mathfrak{L}(n_{i}^{j})}\eta_{\pi(n)},$
so the expectation of $j$-th layer’s random vector is
$\mathbb{E}[\bar{\boldsymbol{z}}^{j}]=\left(\max\limits_{n\in\mathfrak{L}(n_{1}^{j})}\eta_{\pi(n)},\max\limits_{n\in\mathfrak{L}(n_{2}^{j})}\eta_{\pi(n)},\dots,\max\limits_{n\in\mathfrak{L}(n_{c(j)}^{j})}\eta_{\pi(n)}\right).$ (18)
Since we regard the recommendation task as a multi-class classification problem, the conditional probabilities for a given user $u$ satisfy the normalization condition $\sum_{y\in\mathcal{Y}}\eta_{y}(u)=1$. The current expectations of each layer’s random variables corresponding to the rectified label $\bar{z}_{i}^{j}$ are not normalized; hence, a normalization term $\alpha^{j}(u)=\sum_{i=1}^{c(j)}\bar{z}_{i}^{j}(u)$ should be applied to the unnormalized rectified label, resulting in the following label assignment:
Definition 3 (Label Rectification).
For a fixed user $u$, with a given item $y$, the normalized rectified label $\tilde{z}_{i}^{j}(u)$ of node $n_{i}^{j}$ is defined as follows:
$\tilde{z}_{i}^{j}(u)=\begin{cases}\frac{1}{\alpha^{j}(u)},&n_{i}^{j}=\rho^{j}\left(\varphi(y)\right)\wedge y=\operatornamewithlimits{argmax}\limits_{y^{\prime}\in\varphi\left(\mathfrak{L}(n_{i}^{j})\right)}\eta_{y^{\prime}}\\ 0,&\text{else}.\end{cases}$
4.3.2 Loss Modification
We have shown that optimizing the multi-class cross-entropy loss will make each element of the preference vector match the rank of the corresponding element of the expectation vector. However, the whole analysis is based on the fact that the multi-class cross-entropy loss is a Bregman divergence and that each layer’s labels correspond to a one-hot vector. A subsequent question arises — if the labels are rectified as Definition 3, can we still conduct an analysis using the Bregman divergence? The answer is yes. All we need is to modify the original multi-class cross-entropy loss.
Proposition 2.
For any given $w\in[0,+\infty)$, construct a vector $\boldsymbol{e}^{(i)}(w)=(0,\dots,0,w,0,\dots,0)$ where the $i$-th component is $w$ and all other components are 00. For $\boldsymbol{o}\in\mathbb{R}^{M}$, the loss
~$\widetilde{\mathcal{L}}(\boldsymbol{e}^{(i)}(w),\boldsymbol{o})=-w\log\left(\frac{\exp(\boldsymbol{o}_{i})}{\sum_{m=1}^{M}\exp(\boldsymbol{o}_{m})}\right)+w\log w+w-1$ (19)
can be expressed in the form of Bregman divergence.
Proof.
See Appendix A.1. ∎
Compared to the original multi-class cross-entropy loss (i.e., E.q. (10)), the loss function in E.q. (19) includes a slight modification that involves multiplying by a weight and adding a constant when given $w$. This modification is fully compatible with the rectified label. By replacing the original multi-class cross-entropy loss with this modified version and assigning the labels of nodes as Definition 3, for a given instance (u,y) and at the $j$-th layer, the following loss function is optimized:
~$\widetilde{\mathcal{L}}_{j}(u,y)=\widetilde{\mathcal{L}}(\boldsymbol{e}^{(\delta^{j}(y))}(\tilde{z}^{j}_{\delta^{j}(y)}),\boldsymbol{o}^{j})$
$=-\tilde{z}^{j}_{\delta^{j}(y)}\log\left(\frac{\exp o^{j}_{\delta^{j}(y)}}{\sum_{k=1}^{c(j)}\exp o_{k}}\right)+\tilde{z}^{j}_{\delta^{j}(y)}\log\tilde{z}^{j}_{\delta^{j}(y)}+\tilde{z}^{j}_{\delta^{j}(y)}-1,$
and corresponding conditional risk for user $u$ of $j$-th layer is
~$\widetilde{\mathcal{R}}_{j}(u)=\mathbb{E}_{y\sim\eta_{y}(u)}\left[\widetilde{\mathcal{L}}(u,y)\right].$
4.3.3 Theoretical Support
As indicated in Proposition 2, the modified loss $\widetilde{\mathcal{L}}$ still owns the property of Bregman divergence. Notice with the normalization, $\mathbb{E}[\tilde{\boldsymbol{z}}^{j}]\subseteq\operatornamewithlimits{range}(g)$ holds for $1\leq j\leq H$, by the Theorem 1, we have
$\operatornamewithlimits{argmin}\limits_{\boldsymbol{o}^{j}\in\mathbb{R}^{c(j)}}\widetilde{\mathcal{R}}_{j}(u)\subseteq\left\{\boldsymbol{o}^{j}\in\mathbb{R}^{c(j)}|R(\boldsymbol{o}^{j},\mathbb{E}[\tilde{\boldsymbol{z}}^{j}])\right\}$
for any user $u\in\mathcal{U}$, which means optimizing the modified loss with rectified labels can lead the tree model to be top-$k$ retrieval Bayes optimal under beam search. Specifically, we have the following theoretical results:
Lemma 1.
Given a user $u\in\mathcal{U}$, if the tree model, which consists of a tree $\mathcal{T}$ and a preference model $\mathcal{M}$, satisfies $R(\boldsymbol{o}^{j},\mathbb{E}[\tilde{\boldsymbol{z}}^{j}])$ for all $1\leq j\leq H$, then
$\mathcal{B}^{j}(u)\in\operatornamewithlimits{argTopk}\limits_{i\in\left[c(j)\right]}\mathbb{E}[\tilde{\boldsymbol{z}}^{j}_{i}]\Longrightarrow\mathcal{B}^{j+1}(u)\in\operatornamewithlimits{argTopk}\limits_{i\in\left[c(j+1)\right]}\mathbb{E}[\tilde{\boldsymbol{z}}^{j+1}_{i}]$
holds for any beam size $k$ within the range $11\leq j\leq H-$ 1.
Proposition 3.
For any user $u\in\mathcal{U}$, if the tree model, which consists of a tree $\mathcal{T}$ and a preference model $\mathcal{M}$, satisfies that $R(\boldsymbol{o}^{j},\mathbb{E}[\tilde{\boldsymbol{z}}^{j}])$ for all $1\leq j\leq H$, then the tree model is top-$k$ retrieval Bayes optimal for any beam size $k$.
The proofs of Lemma 1 and Proposition 3 are provided in Appendix A.2 and A.3, respectively. The above analysis demonstrates that optimizing the modified loss function with the normalized rectified label can mitigate the issue of sub-optimality under beam search of original multi-class cross-entropy loss from a theoretical aspect.
4.3.4 Practical Implementation
The above analysis and methodology rely on the conditional probability given user $u$, which is unknown in practice. A direct way would be to estimate the conditional probability using the preference model $\mathcal{M}$. However, since $\mathcal{M}$ is a complicated neural network, estimating the conditional probabilities through $\mathcal{M}$ is computationally expensive. Additionally, because the preference model is randomly initialized, its estimations can be highly inaccurate, leading to difficulties in convergence or the risk of falling into local optima. Considering the above shortcomings, we employ a trained small model to estimate the conditional probability in our experiments; the specific details will be introduced in the experimental section.
In practice, we optimize the loss using stochastic gradient descent. Since the gradient of term $w\log w+w-1$ w.r.t. preference score is zero for a given $w$, it can be disregarded in the actual optimization process. Therefore, the loss function (i.e., E.q. (19)) simplifies to the original multi-class cross-entropy loss weighted by $w$. While $w$ should be set as the normalized rectified label $\tilde{z}$ based on our earlier analysis, we find that optimizing the loss with normalization is difficult and use the unnormalized rectified label $\bar{z}$ in practice. In fact, the unnormalized rectified label can already mitigate the suboptimality, as demonstrated by E.q (18). The role of normalization is to make the sum of expectations of random variables $\bar{z}^{j}_{i}$ of the $j$-th layer equal $1$, to satisfy the constraints in the theoretical analysis. Normalization actually influences the learning rate in the stochastic gradient descent compared to the unnormalized label $\bar{z}$. Note that the normalization coefficients across different layers vary dramatically for a given training instance, but we aggregate the loss of all layers together for optimization, so the normalization will make the optimization difficult in practice due to the difficulty of setting a suitable learning rate. Therefore, we optimize the modified loss with the original rectified label $\bar{z}^{j}_{i}$ without normalization, which can be interpreted as an adaptive adjustment of the learning rate in each layer, and the loss for $j$-th layer is
~$\widetilde{\mathcal{L}}_{j}(u,y)=-\bar{z}^{j}_{\delta^{j}(y)}\log\left(\frac{\exp o^{j}_{\delta^{j}(y)}}{\sum_{k=1}^{c(j)}\exp(o^{j}_{k})}\right),$ (20)
and the loss of the whole tree is
~$\widetilde{\mathcal{L}}(u,y)=\sum_{j=1}^{H}\widetilde{\mathcal{L}}_{j}(u,y).$ (21)
5 Negative Sampling to Estimate Softmax Loss Gradient
For the training instance $(u,y)$, calculating its training loss (i.e., E.q. (21)) requires computing the softmax probability of the positive sample at each layer. In practical applications, where the corpus size ranges from millions to billions or even tens of billions, computing softmax probabilities at such scales is extremely inefficient. This necessitates techniques to estimate the gradient of the softmax loss.
5.1 Unbiased estimation of softmax loss gradient
Sampled softmax is proposed to approximate full softmax during model training [59, 60]. It provides significant advantages in recommendation scenarios, such as mitigating popularity bias, discovering and utilizing difficult negative samples, and maximizing ranking metrics [61]. Therefore, we utilize it to alleviate the high computational complexity issue associated with softmax-based loss functions in our proposed method.
In sampled softmax, rather than calculating the training loss of layer $j$ (i.e., E.q. (10)) over all classes, only the positive class and $M$ negative classes are considered. The $M$ negative classes are sampled from all the negative classes according to certain probability distribution $Q$ with replacement. In the rest of this section, we use $s^{j}$ to denote the training samples at layer $j$ and use $q_{s^{j}_{i}}$ ($1\leq i\leq c(j)$) to denote the probability that the $i$-th element of $s^{j}$ can be sampled from the negative nodes of layer $j$. We suppose, without loss of generality, that $s_{1}^{j}$ is always the positive class (i.e. $s^{j}_{1}=\delta^{j}(y)$), while the remaining elements of $s^{j}$ are negative classes of layer $j$. For example, if $M=5$ and the sample set is $s^{j}=\{2,3,4,8,7,4\}$, then $s^{j}$ indicates that the $2$nd node of layer $j$ is the positive node, the $4$th node of layer $j$ is sampled twice, and the other nodes (indexed at $3,8,7$) are each sampled once.
However, we don’t use the outputs w.r.t. sampled nodes of $s^{j}$ to approximate the loss directly. A slight adjustment is conducted for each output by
$\hat{o}_{s_{i}^{j}}^{j}=\left\{\begin{aligned} &o^{j}_{s^{j}_{i}}-\ln(Mq^{j}_{s^{j}_{i}})\ \ \ \ if\ z^{j}_{s_{i}^{j}}=0\\ &o^{j}_{s^{j}_{i}}-\ln(1)\ \ \ \ \ \ \ \ \ \ if\ z^{j}_{s_{i}^{j}}=1\end{aligned}\right.$ (22)
We subtract the natural logarithm of the sampling probability times $M$ from their original preference scores for the sampled negative sample nodes, while the preference scores of the positive samples remain unchanged. This adjustment can guarantee the sampled softmax is unbiased when $M$ is infinite (i.e. $M\to\infty$) [60]. The training loss is calculated over the adjusted outputs and the modified training loss E.q. (20) at layer $j$ can be adjusted to
^$\widehat{\mathcal{L}}_{j}(u,y)=-\bar{z}^{j}_{\delta^{j}(y)}\log\hat{p}^{j}_{\delta^{j}(y)}=-\bar{z}^{j}_{\delta^{j}(y)}\log\frac{\exp\hat{o}_{\delta^{j}(y)}^{j}}{\sum_{k=1}^{M+1}\exp\hat{o}_{s^{j}_{k}}^{j}}.$ (23)
The training loss w.r.t. the whole tree for the training instance $(u,y)$ becomes
^$\widehat{\mathcal{L}}(u,y)=\sum_{j=1}^{H}\widehat{\mathcal{L}}_{j}(u,y).$ (24)
Through simple calculations, the gradient of the modified multi-class cross-entropy loss function (i.e., E.q. (20)) with respect to $o_{i}^{j}$ can be obtained as
[…]
Furthermore, the gradient of the approximate loss (i.e., E.q (23)) with respect to $o_{i}^{j}$ is as follows:
^$\frac{\partial\widehat{\mathcal{L}}_{j}(u,y)}{\partial o_{i}^{j}}=\bar{z}^{j}_{\delta^{j}(y)}\left(\sum_{k=1}^{M+1}\mathbb{I}(s_{k}^{j}=i)\hat{p}_{k}^{j}-z_{i}^{j}\right),$ (26)
where $\mathbb{I}(\cdot)$ is an indicator function that takes the value 0 or 1. Previous literature  [60, 47] has proved that proper specified sampling distribution $Q$ used in sampled softmax can lead to the unbiased estimation of the gradient for the original loss, i.e., $\mathbb{E}_{Q}[\frac{\partial\widehat{\mathcal{L}}^{j}(u,y)}{\partial o_{i}^{j}}]=\frac{\partial\mathcal{L}^{j}(u,y)}{\partial o_{i}^{j}}$. We state their theorem in a different way so that it can be compatible with our method:
Theorem 2 (Theorem 2.1 of  [47]).
The gradient of loss E.q. (23) (or E.q. (24)) w.r.t. sampled softmax is an unbiased estimator of the gradient of the loss E.q. (20) (or E.q. (21)) w.r.t. full softmax if and only if $q^{j}_{i}\propto\exp{o_{i}^{j}}$ holds where $1\leq i\leq c(j)$ and the $i$-th node isn’t a positive node at layer $j$.
This theorem indicates that if the sampling probability is proportional to the exponential of the preference w.r.t. the negative node, we can get an unbiased estimation of the training loss gradient. However, it still results in linear time complexity w.r.t. the corpus size, similar to calculating the full softmax, because it requires computing the preference score for every node. To address this problem, previous literature  [47, 62] develops kernel-based methods to approximate the full softmax for inner product models. Nonetheless, their kernel-based methods aren’t suitable for our approach, as our preference model is a complicated neural network. To reduce the computational complexity while maintaining compatibility with our approach, we explore alternative sampling strategies, including uniform sampling and top-down sampling.
5.2 Uniform Sampling and Top-down Sampling
Uniform sampling is straightforward to comprehend and can be implemented rapidly, making it a popular choice in numerous research and practical applications. It is also applicable to our methodology, where the negative nodes at each layer can be directly obtained through uniform sampling. Concretely, at layer $j$, each negative node is sampled uniformly with replacement until $m$ negative nodes are sampled. The $i$-th node of layer $j$ is sampled with probability $q_{i}^{j}=\frac{1}{c(j)-1}$ at each sampling operation. Besides its simplicity, uniform sampling offers another significant advantage in that it can lead to good exploration ability. As every negative node can be chosen with equal chance, the sampled nodes deserve good diversity among each other. Additionally, uniform sampling does not introduce any extra computational overhead. Unlike kernel-based sampling methods, which require computing the kernel to obtain the distribution $Q$, uniform sampling keeps the training process efficient and fast. In our experiments, we have explored uniform sampling to sample the negative nodes at each layer, i.e. $q_{i}^{j}=\frac{1}{c(j)-1}$ at E.q.(22).
While uniform sampling possesses many advantageous properties, the uniform distribution often diverges significantly from the true softmax distribution. To better estimate the gradient of the loss function, it is desirable to use a sampling distribution $Q$ that closely approximates the true softmax distribution. In response to this need, we propose a tree-based sampling method called top-down sampling. This method leverages both the tree index and the preference model to make the sample distribution $Q$ closer to the true softmax distribution than the uniform distribution.
We elaborate on the proposed top-down sampling method, which is implemented by modifying the beam search retrieval process (see Algorithm 1). Firstly, during sampling along the tree, only one node is retained at each layer (i.e., we set $k=1$ in Algorithm 1). Secondly, instead of directly choosing the child node with the highest preference score, we sample one child node for expansion. This sampling-based expansion is performed by local softmax probabilities, calculated based on the preference scores of all child nodes under the current parent node. Concretely, after expanding node $n_{i}^{j}$, the expanding probability for node $n_{c}^{j+1}\in\mathcal{D}(n_{i}^{j})$ is:
$\tilde{q}_{c}^{j+1}=\frac{\exp{o_{c}^{j+1}}}{\sum\limits_{k\in\left\{\delta(n)\mid n\in\mathcal{D}(n_{i}^{j})\right\}}\exp{o_{k}^{j+1}}}.$ (27)
Since only one parent node is retained (we have set $k=1$ in Algorithm 1) and each parent node has only a few child nodes, the local softmax probabilities can be efficiently computed at each layer. Then, by expanding one node at each layer, we obtain a sample path through the tree from the root to a leaf. These sampled nodes along this path serve as the negative nodes for their respective layers. Finally, by repeating this sampling process $M$ times, we can obtain $M$ negative nodes for each layer. Notice that performing beam search retrieval on the tree can be completed with a logarithmic complexity w.r.t. the corpus size (i.e., $O(\log|\mathcal{Y}|)$); thus, the additional training time introduced by our proposed sampling method remains acceptable.
Figure 1: Illustration of top-down sampling along the tree. The number beside the edge is the child’s expanding probability given the parent.
To elucidate the top-down sampling method, we illustrate the process in Figure 1. The sampling process starts from the root, i.e., node 00. The node 00’s child nodes are node $1$ and node $2$, with expanding probabilities of $0.3$ and $0.7$, respectively, in this case. Node $1$ is sampled for expansion, and the process continues from node $1$, repeating until the leaf node $8$ is sampled. Finally, the path $0\rightarrow 1\rightarrow 3\rightarrow 8$ is sampled from the tree, and the corresponding sampled negative nodes are $\{1,3,8\}$. For any node in the tree, the probability of being sampled via top-down sampling is the product of the expanding probabilities along the edges from the root to that node. For example, in Figure 1, the sampling probability for the node $4$ is $0.18=0.3\times 0.6$, and for the node $9$ is $0.144=0.3\times 0.6\times 0.8$. It’s easy to verify that the sum of the probabilities of each node being sampled is exactly $1$ at any layer, which affirms the correctness of our proposed top-down sampling.
To obtain an unbiased estimation of the gradient for the original training loss, Theorem 2 dictates that the sampling probability for each node should be proportional to the exponential of its preference score. Our proposed top-down sampling method is grounded in this principle. Specifically, if the preference scores for a parent node and its child nodes satisfy certain conditions, the sampling probability for any node in the top-down sampling aligns precisely with the softmax probability of that node. This is formalized in the following theorem:
Theorem 3.
For any node $n_{i}^{j}$ of the tree, if the preference score $o_{i}^{j}$ satisfies the following condition:
$\exp(o_{i}^{j})\propto\sum_{k\in\left\{\delta(n)\mid n\in\mathcal{D}(n_{i}^{j})\right\}}\exp(o_{k}^{j+1}),$ (28)
and the proportionality coefficients within the same layer are equal, then the probability of sampling the $i$-th node at the $j$-th layer through top-down sampling precisely equals the softmax probability of node $n_{i}^{j}$ at $j$-th layer, that is,
$q_{i}^{j}=\frac{\exp(o_{i}^{j})}{\sum_{k=1}^{c(j)}\exp(o_{k}^{j})}.$ (29)
Note that the sum of the exponentials of the nodes’ preference scores is shared among nodes within the same layer in E.q. (29). Consequently, $q_{i}^{j}$ is proportional to $\exp(o_{i}^{j})$ in layer $j$, satisfying the conditions outlined in Theorem 2. Although Theorem 29 imposes a relatively stringent condition that may not be fully satisfied in practical applications, it provides valuable insights for designing sampling schemes and offers justification for the rationality of our proposed top-down sampling method.
6 Tree Index Updating
The performance of tree-based retrieval is influenced by both the preference model and the tree index. Without sufficient domain knowledge, initializing a well-structured tree is challenging. Even if the preference model converges during training, it can only adapt to the current, potentially suboptimal tree. Consequently, it is necessary to update the tree index alongside the optimization of the preference model. In this section, we introduce how to update the tree index (i.e., updating the mapping $\pi$) while keeping the preference model fixed.
At first, we employ a method similar to TDM and JTM to construct the initial tree. The tree is used to represent the user interests’ hierarchical information so that similar items should organized in close positions on the last layer of the tree. It’s natural to rely on the category information of items to build the initial tree. Firstly, we shuffle all the categories and put the items together if they share the same category ID. If an item belongs to more than one category, a random one is assigned to the item for uniqueness. Secondly, the items within each category are recursively divided into two equal parts until each partition contains only one item. In this way, we can construct a near-complete binary tree from root node to leaf nodes and some simple balancing strategy can be used to adjust the tree. We use a binary tree in our experiments. In fact, any multi-way tree can be constructed in such a way if we partition each cluster into more nearly equal parts at each iteration.
Tree index updating is still based on the training set $S=\{(u_{i},y_{i})\}_{i=1}^{m}$. We partition $S$ into different user-item pair sets, denoted by $\mathcal{A}_{y}=\{(u_{i},y_{i})|y_{i}=y\}$, where $\mathcal{A}_{y}$ consists of all user-item pairs where the item is exactly $y$. Given the old mapping $\pi$ and each user-item pair set $\mathcal{A}_{y}$, we fix the preference model and assign each item to the tree nodes step by step, from the root node to the leaf nodes, to obtain the new mapping $\pi^{\prime}$.
All items are initially assigned to the root node, i.e., current level $l=0$. Then, we try to assign all the items to the nodes at level $min(H,l+d)$, where the hyperparameter $d$ determines the number of layers to skip during the allocation process along the tree. For the item $y$ assigned to the root node, it needs to be reassigned to one of the nodes at level $min(H,l+d)$. Since we are using a binary tree, there will be $c\leq 2^{d}$ candidate nodes. Without loss of generality, we denote these candidate nodes as $n_{1},n_{2},\dots,n_{c}$. The preference of user $u$ ($u\in\mathcal{A}_{y}$) for node $n_{i}$ ($1\leq i\leq c$) is denoted as $o_{i}(u)$ and is calculated by the fixed preference model. We define item $y$’s matching score for node $n_{i}$ ($1\leq i\leq c$) as follows:
$Score(y,n_{i})=\sum_{(u,y)\in\mathcal{A}_{y}}\log\hat{p}(n_{i}|u)$ (30)
where $\hat{p}(n_{i}|u)=\frac{\exp{o_{i}(u)}}{\sum_{k=1}^{c}\exp{o_{k}(u)}}$. We rank the $c$ scores and assign $y$ to the node with the highest matching score. All items are disjointly assigned to the nodes at level $min(H,l+d)$, where the number of items assigned to the node $n$ at this level equals the number of leaf nodes in the subtree rooted at $n$. If the number of assigned items exceeds the limitation, $y$ is reassigned to the node with the second-highest matching score, or to successive nodes, until the assignment satisfies the limit. In this way, we can assign each item belonging to the root to a node at level $min(H,l+d)$. Now, we regard $n_{1},\dots,n_{c}$ as root nodes of $c$ sub-trees and repeat the process recursively to assign the items to deeper layers until each leaf node is assigned an item. Finally, we can get the new mapping $\pi^{\prime}$. To compute the matching score (i.e., E.q. (30)), we need to compute the denominator $\sum_{k=1}^{c}\exp{o_{k}(u)}$, so the number $d$ need to be small to fit the time complexity and computational resource; in our experiments, we set $d=7$.
Figure 2: The process of updating the mapping between items and leaf nodes, where $d=2$ in this case.
To better understand the tree update process, Figure 2 provides an illustrative example. In this example, there are eight items labeled $a$ to $h$, and the step $d$ is set to $2$. Initially, all items are located in the root node. The task is to allocate these items to the layer that is at a distance of $2$ from the root. This layer contains four nodes, with each node receiving two items. Based on the matching scores, item $f$ should ideally be assigned to node $4$, which has the highest matching score. However, since node $4$ already contains two items (i.e. item $c$ and $d$), item $f$ is reallocated to node $5$, which has the next highest matching score. After the items are allocated to the second layer, the process continues to deeper layers. Since the leaf nodes are located at the third layer, which is only $1$ step away from the second layer (less than the step $d$), the items are directly assigned to the leaf layer. Ultimately, each leaf node is allocated one item, and the tree updating is completed.
In fact, our tree learning strategy also carries on the hierarchical clustering from coarse grain to fine grain, and $d$ controls the grain size. The main difference between ours and the one in JTM is the calculation of matching scores. The matching score in JTM is calculated by summing up the log-likelihood probability as the preference model outputs the probability directly while our strategy relies on the softmax probability when calculating the matching scores. Throughout the training process, model optimization and tree updating are alternated, leading to the gradual improvement of the model’s retrieval performance until convergence is finally achieved.
7 Generalization Analysis
In this section, we provide a generalization analysis of the deep tree-based retriever, to offer a comprehensive theoretical understanding of our proposed method. Our analysis employs Rademacher complexity to derive a data-depended generalization error bound of the tree model optimized with the modified loss with the rectified label in conjunction with the sampled softmax technique, demonstrating its great generalization capability.
The variation of DIN is used as the preference model, which is formalized in Section 3.2.2. Suppose the norm of the weight matrix is bounded, we can get such function space for the model:
=$\begin{split}\mathcal{M}=\Big{\{}f:(u,n)\mapsto f_{Din}(u,n)\mid\|\boldsymbol{w}_{n}\|_{2}\leq B_{0},\forall n\in\mathcal{N};\\ \|\boldsymbol{W}_{k}\|_{1}\leq B_{1},\forall k\in[L];\ \|\boldsymbol{W}_{w}^{(j)}\|\leq B_{2},\forall j\in{1,2}\Big{\}}.\end{split}$ (31)
Lemma 2 (Rademacher Complexity of DIN).
Suppose $\forall u\in\mathcal{U},\ \forall k\in[K],\|\boldsymbol{a}_{k}^{(u)}\|_{2}\leq B_{a}$, then the Rademacher complexity of ${\mathcal{M}}$ can be bounded as follows:
^$\hat{\mathcal{R}}_{m}(\mathcal{M})\leq\frac{2c_{\phi}^{L-1}B_{1}^{L}(B_{0}+KB_{w_{1}}+KB_{w_{2}}\mathcal{\tau})}{\sqrt{m}},$ (32)
where $B_{w_{1}}=c_{\phi}^{2}B_{2}^{2}B^{2}_{a}$, $B_{w_{2}}=c_{\phi}^{2}B_{0}B_{2}^{2}\left(B_{a}^{2}+B_{a}\right)$, and $\tau=\sqrt{(2B-1)|\mathcal{Y}|}/\sqrt{B-1}$.
The proof of Lemma 2 is provided in Appendix A.5. With the Rademacher complexity of the DIN model, we now derive the generalization bound for the modified loss with rectified labels (i.e., E.q. (21)) in Section 4.3. Since the overall loss for the tree is the sum of the losses at each layer, we begin by examining the generalization bound for the loss at the $j$-th layer. The result is presented as follows:
Lemma 3.
Suppose the function in $\mathcal{M}$ is bounded by a constant $B_{\mathcal{M}}$, the following inequality holds with a probability of at least $1-\delta$:
$\begin{split}\mathbb{E}_{(u,y)\sim\mathcal{D}}\left[\widetilde{\mathcal{L}}_{j}(u,y)\right]&\leq\frac{1}{m}\sum_{i=1}^{m}\widetilde{\mathcal{L}}_{j}(u_{i},y_{i})\\ &+8c(j)\frac{c_{\phi}^{L-1}B_{1}^{L}(B_{0}+KB_{w_{1}}+KB_{w_{2}}\mathcal{\tau})}{\sqrt{m}}\\ &+\Big{(}4\log c(j)+8B_{\mathcal{M}}\Big{)}\sqrt{\frac{2\log\left(4/\delta\right)}{m}}.\end{split}$ (33)
The theoretical analysis presented above demonstrates the generalization capacity of the proposed method of optimizing the preference model, where the objective is to minimize a modified softmax loss with rectified labels. In practice, however, the sampled softmax is employed to speed optimization as detailed in Section 5. To ensure theoretical analysis consistent with our proposed method, we incorporate the bias introduced by the sampled softmax loss w.r.t. the softmax loss into the theoretical result. For a given instance $(u,y)$, the model $\mathcal{M}$ outputs $c(j)$ preference scores for nodes in $j$-th layer, forming a logit vector $\boldsymbol{o}^{j}\in\mathbb{R}^{c(j)}$. Suppose the $i$-th node is positive (i.e., $i=\delta^{j}(y)$) while the others are negative. The softmax loss for layer $j$ is given by:
$\ell^{j}_{softmax}(u,y)=-\log\frac{\exp\boldsymbol{o}^{j}_{i}}{\sum_{k=1}^{c(j)}\exp\boldsymbol{o}^{j}_{k}}.$
A subset of $M$ indices of negative nodes $\mathcal{I}^{\prime}_{M}$ is sampled according to the sample distribution $Q^{j}(u,y)$, where $u$ and $y$ specify that the sampling distribution is dependent on the instance $(u,y)$ and can be omitted. The adjusted logit vector $\boldsymbol{\hat{o}}^{j}$ is then computed based on $\boldsymbol{o}^{j}$ and $Q^{j}$ as specified in E.q. (22). The sampled softmax loss for layer $j$ is calculated using $\boldsymbol{\hat{o}}^{j}$ and $\mathcal{I}^{\prime}_{M}$ as follows:
$\ell^{j}_{sampled\text{-}softmax}(u,y,\mathcal{I}^{\prime}_{M})=-\log\frac{\exp\boldsymbol{\hat{o}}^{j}_{i}}{\sum_{i^{\prime}\in\mathcal{I}^{\prime}_{M}\cup\{i\}}\exp\boldsymbol{\hat{o}}^{j}_{i^{\prime}}}.$
The bias of the sampled softmax loss w.r.t. the softmax loss is presented as follows:
Lemma 4.
For any given instance $(u,y)$ and for the $j$-th layer, the softmax loss can be bounded by the sum of two terms: the expectation of the sampled softmax loss over $\mathcal{I}^{\prime}_{M}$ and the KL divergence between $Q^{j}$ w.r.t. $P^{j}$. It is expressed as follows:
$\begin{split}\ell^{j}_{softmax}(u,y)\leq\mathbb{E}_{\mathcal{I}^{\prime}_{M}}\left[\ell^{j}_{sampled\text{-}softmax}(u,y,\mathcal{I}^{\prime}_{M})\right]+D_{\text{KL}}(Q^{j}\|P^{j}),\end{split}$
where $P^{j}(u,y)$ is the softmax distribution in the layer $j$ for the instance $(u,y)$, defined as $p^{j}_{i}\triangleq\exp{\boldsymbol{o}^{j}_{i}}/\sum_{k=1}^{c(j)}\exp{\boldsymbol{o}_{k}^{j}}$.
Then, combine Lemma 2 , Lemma 3 and Lemma 4, and aggregate from the $1$st layer to the $H$-th layer, we can get following result:
Theorem 4.
The following inequality holds with a probability of at least $1-\delta$:
$\begin{split}\mathbb{E}_{(u,y)\sim\mathcal{D}}\left[\widetilde{\mathcal{L}}(u,y)\right]&\leq\frac{1}{m}\sum_{i=1}^{m}\sum_{j=1}^{H}\widehat{{\mathcal{L}}}_{j}(u_{i},y_{i})\\ &+\sum_{j=1}^{H}\mathbb{E}_{(u,y)\sim\mathcal{D}}\left[D_{KL}(Q^{j}||P^{j})\right]\\ &+8\tau^{2}\frac{c_{\phi}^{L-1}B_{1}^{L}(B_{0}+KB_{w_{1}}+KB_{w_{2}}\mathcal{\tau})}{\sqrt{m}}\\ &+\Big{(}8\log(\tau)+8HB_{\mathcal{M}}\Big{)}\sqrt{\frac{2\log\left(4/\delta\right)}{m}}.\end{split}$ (34)
Theorem 4 provides a comprehensive theoretical analysis of our proposed method. As shown in E.q. (34), the left side of the inequality is the expected risk of the modified loss with rectified labels. Minimizing this risk results in optimal outcomes under beam search, as detailed in Section 4.3. On the inequality’s right side, the first term corresponds to the empirical loss optimized in practice. The second term is the expected KL divergence of the negative sampling distribution w.r.t. the softmax distribution at each layer. It can be observed that the closer the negative sampling distribution is to the softmax distribution, the lower the generalization risk. This precisely aligns with the core principle of our proposed top-down sampling, providing insight into why it performs well from a generalization perspective. The third term captures the influence of model parameters and tree structure on generalization risk. Notably, when the number of items (i.e., $|\mathcal{Y}|$) is fixed, increasing the number of branches $B$ results in a smaller $\tau$, leading to a tighter generalization bound. Moreover, when the sample size $m$ is sufficiently large, the last two terms tend to vanish, meaning that with enough data, the generalization risk can be effectively controlled by the empirical loss and the KL divergence.
8 Experiments
We conduct a series of experiments designed to answer the following research questions related to the joint optimization framework of preference models and tree indexes: RQ-1: Is the multi-class classification training mode more suitable than the binary classification training mode for the tree model? RQ-2: Is the modified loss with rectified labels effective in mitigating the sub-optimality of the multi-class cross-entropy loss under beam search in practice? RQ-3: Is the proposed sampling method, which uses sampled-softmax theory for negative sampling to accelerate training, effective? RQ-4: Can the DTR model learn a reasonable mapping relationship between items and leaf nodes? RQ-5: Does the performance of the proposed sampling method fluctuate with the number of negative samples?
8.1 Datasets
The experiments are conducted on the four real-world datasets [^1] https://drive.google.com/drive/folders/1ahiLmzU7cGRPXf5qGMqtAChte2eYp9gI: Movie Lens 10M (abbreviated as Movie), MIND Small Dev(abbreviated as MIND), Amazon Books(abbreviated as Amazon), Tmall Click (abbreviated as Tmall). Movie is a public film rating dataset, MIND is a news recommendation dataset collected by Microsoft, Amazon is a dataset documenting book purchases and ratings on the Amazon e-commerce site, Tmall records the shopping behaviors of users on Alibaba’s Tmall marketplace. We process all datasets into a form of implicit feedback, where interactions between users and items are marked as 1, and non-interactions as 0. Additionally, users with fewer than 15 interactions were filtered out, with the post-filtering data statistics presented in Table I. For each dataset, 10% of users are randomly selected as validation users, 10% as test users, and all remaining users as training users. For training users, each interaction sequence is constrained to the most recent 70 interactions, with shorter sequences padded with zeros. The first 69 interactions serve as input, while the 70th is used as the label. For validation and test users, the first half of their interaction history is used as input, with the second half serving as labels for prediction.
TABLE I: Statistics of datasets
Dataset #User #Item #Interaction Density
Movie 69,878 10,677 10,000,054 1.34%
MIND 36,281 7,129 5,610,960 2.16%
Amazon 29,980 67,402 2,218,926 0.11%
Tmall 139,234 135,293 10,487,585 0.05%
8.2 Baselines
We compared the proposed DTR model with other related algorithms, which fall into two broad categories: 1) models that employ brute-force search without indexing, and 2) tree-based models similar to our DTR approach. It should be noted that the former requires a significant amount of time to perform retrieval, comparing them with index-based retrieval models is not fair, and their results are only used as a reference. The introductions of these algorithms are as follows:
  • Item-CF [63] is a basic collaborative filtering algorithm and is widely used in recommendation tasks.
  • DIN [1] employs an attention mechanism to capture the user preference for items from the user’s interaction history. Since it cannot be used for efficient recommendation directly, the performance based on brute-force search is reported as a reference metric.
  • YoutubeDNN [64] utilizes inner products to compute user preference scores. Once the user and item embeddings are learned, the recommendation task is transformed into a Maximum Inner Product Search (MIPS). We use the implementation provided by DeepMatch [^2] https://github.com/shenweichen/DeepMatch and report the results obtained through brute-force search.
  • PLT [65] organizes items into a label tree, where each node represents a conditional probability estimator. The probability of an item is determined by multiplying the probabilities estimated along the path from the root to the corresponding leaf. Recommendations are then made based on the calculated probability.
  • TDM [11] is a tree-based recommendation model proposed by Alibaba Group. It simultaneously learns the preference model and the tree-index structure, using beam search during inference to perform efficient recommendation.
  • JTM [12] is an enhanced version of TDM. It introduces a method for jointly learning the index structure and the user preference prediction model.
  • OTM [54] mitigates the training-testing discrepancies of tree models under beam search by sampling negative nodes through beam search and correcting the pseudo-labels of nodes during training. It is a state-of-the-art tree model.
8.3 Metric
To evaluate the performance of the retrieval model, we use $Precision@K$, $Recall@K$, and $F\text{-}measure@K$ as the evaluation criteria. Suppose $\mathcal{P}(u)$ ($\left|\mathcal{P}(u)\right|=K$) denotes the set of items retrieved by the model for user $u$, and $\mathcal{G}(u)$ represents the set of items that the user has interacted with, i.e., the ground truth associated with user $u$. For user $u$, the evaluation metrics $Precision@K$ and $Recall@K$ are calculated as follows:
$\begin{split}Precision@K(u)=\frac{|\mathcal{P}(u)\cap\mathcal{G}(u)|}{K}\ ,\\ Recall@K(u)=\frac{|\mathcal{P}(u)\cap\mathcal{G}(u)|}{|\mathcal{G}(u)|}\ ,\ \ \ \end{split}$
and $F\text{-}measure@K$ is calculated as follows:
$F\text{-}measure@K(u)=\frac{2\cdot Precision@K(u)\cdot Recall@K(u)}{Precision@K(u)+Recall@K(u)}.$
8.4 Settings
We provide detailed settings of the baseline algorithm and our algorithm.
For Item-CF, the number of neighbors is set to $500$ for the Movie, MIND and $1000$ for Amazon, Tmall respectively. The cosine similarity coefficient is employed to measure the similarity between items. DIN comprises an embedding layer followed by an MLP with layer sizes [128,64,2], which outputs the probability of a user liking a given item. YouTubeDNN also includes an embedding layer and an MLP with layer sizes [128,64,24], and the number of negative samples is set to 1,000.
For all tree-based algorithms, a binary tree is selected as the tree index (i.e., $B$ is set to 2), and the DIN model or its variance is used as the preference score model. The first 69 interactions in the sequence are segmented using 10 sliding windows, with the number of interactions in each window being $\left[20,20,10,10,2,2,2,1,1,1\right]$ respectively. For all tree-based algorithms, except for OTM, the same tree structure initialized based on item categories is used, whereas OTM employs a tree structure learned from JTM as described in [54]. TDM, JTM, and DTR include a process for updating the tree index. Due to the instability associated with tree updating in TDM, it is configured to include only 4 trees, while JTM and DTR are set to include 12 trees each. Negative nodes are the sibling nodes of the positive node in PLT, while for other tree-based algorithms, the negative nodes are sampled from the same layer as the positive node in each layer. For TDM and JTM, the number of negative samples per layer is set to 6, 3, 5, and 6 for the Movie, MIND, Amazon, and Tmall datasets, respectively. In OTM, since sampled nodes are the candidates generated from the beam search process, the number of samples is set to $2k$, where $k$ is the beam size. For DTR, the number of negative samples is set to 70.
DTR employs two sampling methods: uniform sampling and top-down sampling, denoted as DTR(U) and DTR(T), respectively. DTR(T) further incorporates a modification that includes a modified loss with the rectified label, referred to as DTR(T-RL). For DTR(T-RL), we need to use a trained small model to estimate the conditional probability. SASRec is chosen to be the estimator. For a given user $u$, the historical sequence is input into SASRec to generate a semantic vector, which is then used to compute the inner product with each item embedding. Finally, the inner products across the entire item set are processed through a softmax function to obtain the conditional probabilities. We train SASRec with the same trainset as other algorithms. The number of encoder layers, the number of heads, and the model dimension are tuned on the validation dataset for each dataset respectively. When evaluated on the testset, the trained SASRec performs at most 0.8% lower than DTR(T) in terms of $F\text{-}measure@20$ across four datasets, indicating that it is a well-behaved estimator. Notice that SASRec is a two-tower model, so the additional computation required for probability estimation is relatively minimal. In our experiments, the time spent on estimating probabilities accounts for approximately 17.5% to 22.6% of the total time across four datasets, which is entirely acceptable.
In all algorithms, the dimension of the item embedding is set to 24, and the batch size is set to 100. We employ Adam as the optimizer [66], with a learning rate initialized at $1.0e$-$3$ and subjected to exponential decay. For the tree-based algorithms, the beam size $k$ is set to 150, i.e. we select at most 150 items to expand when we conduct a beam search along the tree. All experiments are conducted on a Linux server equipped with a $3.00$ GHz Intel CPU, $300$ GB of main memory, and NVIDIA $20/30$ series GPUs.
8.5 Comparation results
TABLE II: Experimental results for different algorithms across four datasets.
Dataset Algorithm top-K=20 top-K=40 top-K=60
Precision Recall F-measure Precision Recall F-measure Precision Recall F-measure
Movie Item-CF 0.1058 0.0508 0.0600 0.1211 0.1218 0.1037 0.1266 0.1886 0.1292
DIN 0.2132 0.1142 0.1309 0.1957 0.2013 0.1706 0.1807 0.2685 0.1848
YoutubeDNN 0.2043 0.1113 0.1270 0.1874 0.1924 0.1630 0.1746 0.2560 0.1770
PLT 0.1606 0.0859 0.0981 0.1478 0.1489 0.1264 0.1385 0.2006 0.1389
TDM 0.1901 0.1012 0.1159 0.1741 0.1770 0.1499 0.1624 0.2383 0.1641
JTM 0.2141 0.1169 0.1335 0.1922 0.1993 0.1682 0.1770 0.2650 0.1815
OTM 0.2321 0.1290 0.1470 0.2009 0.2086 0.1767 0.1822 0.2687 0.1862
DTR(U) 0.2367 0.1231 0.1431 0.2054 0.2032 0.1759 0.1842 0.2616 0.1850
DTR(T) 0.2511 0.1343 0.1546 0.2135 0.2169 0.1853 0.1893 0.2772 0.1924
DTR(T-RL) 0.2545 0.1374 0.1580 0.2176 0.2240 0.1905 0.1936 0.2876 0.1985
Mind Item-CF 0.2243 0.1218 0.1416 0.1962 0.1977 0.1740 0.1769 0.2518 0.1833
DIN 0.4174 0.1912 0.2320 0.3485 0.2930 0.2756 0.3042 0.3634 0.2861
YoutubeDNN 0.4025 0.1850 0.2243 0.3357 0.2851 0.2671 0.2936 0.3541 0.2775
PLT 0.2341 0.0927 0.1168 0.2024 0.1532 0.1494 0.1818 0.2011 0.1626
TDM 0.3156 0.1365 0.1687 0.2706 0.2236 0.2116 0.2375 0.2832 0.2225
JTM 0.3457 0.1547 0.1892 0.2806 0.2354 0.2213 0.2432 0.2925 0.2288
OTM 0.3724 0.1609 0.1978 0.3144 0.2523 0.2404 0.2785 0.3188 0.2544
DTR(U) 0.3904 0.1804 0.2188 0.3063 0.2621 0.2448 0.2612 0.3191 0.2482
DTR(T) 0.4044 0.1853 0.2250 0.3318 0.2808 0.2634 0.2889 0.3472 0.2726
DTR(T-RL) 0.4210 0.1938 0.2350 0.3401 0.2869 0.2695 0.2957 0.3544 0.2787
Amazon Item-CF 0.0175 0.0178 0.0169 0.0284 0.0554 0.0360 0.0314 0.0860 0.0440
DIN 0.0706 0.0495 0.0528 0.0605 0.0828 0.0625 0.0536 0.1080 0.0641
YoutubeDNN 0.0606 0.0399 0.0436 0.0522 0.0682 0.0527 0.0470 0.0910 0.0553
PLT 0.0238 0.0147 0.0162 0.0201 0.0245 0.0193 0.0181 0.0321 0.0202
TDM 0.0514 0.0333 0.0365 0.0419 0.0530 0.0416 0.0361 0.0674 0.0418
JTM 0.0606 0.0411 0.0444 0.0507 0.0674 0.0517 0.0452 0.0892 0.0537
OTM 0.0704 0.0440 0.0486 0.0584 0.0710 0.0566 0.0514 0.0923 0.0583
DTR(U) 0.0709 0.0502 0.0535 0.0594 0.0820 0.0619 0.0521 0.1067 0.0630
DTR(T) 0.0757 0.0517 0.0557 0.0613 0.0829 0.0631 0.0533 0.1073 0.0639
DTR(T-RL) 0.0777 0.0542 0.0580 0.0626 0.0847 0.0644 0.0544 0.1093 0.0651
Tmall Item-CF 0.0093 0.0084 0.0083 0.0181 0.0314 0.0216 0.0205 0.0498 0.0273
DIN 0.0488 0.0350 0.0376 0.0379 0.0531 0.0405 0.0322 0.0665 0.0397
YoutubeDNN 0.0331 0.0211 0.0236 0.0276 0.0349 0.0278 0.0243 0.0457 0.0287
PLT 0.0167 0.0111 0.0122 0.0133 0.0176 0.0137 0.0114 0.0224 0.0137
TDM 0.0310 0.0209 0.0229 0.0253 0.0335 0.0262 0.0220 0.0433 0.0266
JTM 0.0394 0.0275 0.0298 0.0315 0.0427 0.0331 0.0269 0.0538 0.0327
OTM 0.0459 0.0283 0.0318 0.0361 0.0433 0.0353 0.0307 0.0546 0.0353
DTR(U) 0.0442 0.0312 0.0337 0.0356 0.0489 0.0377 0.0308 0.0623 0.0376
DTR(T) 0.0464 0.0329 0.0355 0.0364 0.0503 0.0385 0.0312 0.0634 0.0382
DTR(T-RL) 0.0482 0.0342 0.0369 0.0373 0.0514 0.0395 0.0317 0.0644 0.0388
The experiment results are shown in Table II. In the table, the best performance of tree-based algorithms is highlighted in bold. Additionally, any results from brute-force search algorithms that exceed the best performance of tree-based algorithms are also emphasized in bold for reference. Based on the results, we can make the following findings:
  • The proposed DTR algorithm consistently outperforms both TDM and JTM across all four datasets. When compared to the enhanced version OTM, DTR remains highly effective: DTR(U) outperforms OTM in most cases, while DTR(T) and DTR(T-RL) consistently surpass OTM. For example, considering $F\text{-}measure@20$, DTR(T-RL) shows improvements of 7.48%, 18.81%, 19.34%, and 16.03% across the four datasets compared to OTM. This result demonstrates that the multi-class classification training mode is better suited for jointly optimizing the preference model and tree structure index than the binary classification training mode of TMD, JTM, and OTM, effectively answering RQ-1.
  • The superiority of DTR(T-RL) over DTR(U) is consistently observed across all datasets. For example, in terms of $F\text{-}measure@20$, DTR(T-RL) achieves improvements of 2.20%, 4.44%, 4.13%, and 3.94%, respectively, across the four datasets. This outcome is in line with the theoretical insights discussed in Sections 4.2 and 4.3. It confirms that the modified loss function with rectified label effectively mitigates the sub-optimality of the multi-class cross-entropy loss when applied under beam search conditions, thus providing an effective answer to RQ-2.
  • The performance of DTR(T) consistently surpasses that of DTR(U). For instance, when the top-K value is set to 20, DTR(T) shows an increase in $F\text{-}measure$ by 8.04%, 2.83%, 4.11%, and 4.41% across the four datasets, respectively. These improvements suggest that the proposed top-down sampling method more effectively aligns with the requirements of sampled softmax theory, as its sampling probabilities are more closely aligned with the actual softmax probabilities. This result validates the theoretical analysis presented in Section 5.2 (i.e., Theorem 29), and effectively addresses RQ-3.
  • Compared with brute-force search retrieval algorithms, DTR(T-RL) continues to demonstrate strong performance. It outperforms YouTubeDNN across all four datasets and surpasses DIN on the Movie and Amazon datasets. This observation suggests that joint optimization of the index and preference model can exhibit superior performance compared to optimizing them independently. Moreover, with the help of the tree index structure, the preference model can exclude some distracting items in certain scenarios, thereby enhancing retrieval performance.
8.6 Effectiveness of index updating
We demonstrate the effectiveness of the tree index updating strategy employed in DTR, i.e. the validity of updating the mapping relationships between the tree‘s leaf nodes and items. We illustrate how $F\text{-}measure@20$ changes with updates to the tree index across four datasets. Since PLT and OTM do not involve tree updates, and JTM is an enhanced version of TDM with a tree updating strategy that better aligns with the preference model, we mainly compare JTM and the three variations of DTR in this subsection.
(a) Movie
(b) MIND
(c) Amazon
(d) Tmall
Figure 3: The variation of $F\text{-}measure@20$ with tree updating across four algorithms.
The experimental results are presented in Figure 3. From these results, the following findings can be made:
  • The three variations of DTR consistently show an initial increase followed by a gradual convergence across four different datasets. This phenomenon suggests that the tree update strategy utilized by DTR is effective in establishing a reasonable mapping between items and the tree’s leaf nodes. Consequently, the index structure evolves alongside the preference model, leading to improved retrieval performance. This result convincingly addresses RQ-4.
  • The three variations of DTR consistently exhibit superior performance, significantly outperforming JTM across four datasets. This result further illustrates the superiority of the multi-class classification training mode employed by DTR over the binary classification training mode used by JTM, thereby answering RQ-1 once again.
  • Compared to DTR(T), DTR(T-RL) consistently offers advantages during the tree updating process. This phenomenon further demonstrates that the modified loss function, incorporating the rectified label, can effectively enhance the retrieval performance of the tree-based model, aligning with our theoretical analysis and also answering RQ-2.
  • The variation of DTR based on top-down sampling consistently outperforms the one based on uniform sampling. This result indicates that top-down sampling can estimate the softmax gradient with greater accuracy compared to uniform sampling. It validates the proposed sampling method’s rationality and echoes the theoretical conclusions underpinning top-down sampling, also addressing RQ-3.
8.7 Sensitivity w.r.t. number of negative samples
We analyze the impact of the number of negative samples on retrieval performance for our proposed DTR and its variations. We demonstrate how the $F$-$measure@20$ varies with the number of negative samples, which ranges from $10$ to $90$ in increments of $10$. All experiments in this subsection are conducted on the initial tree constructed based on the item categories.
The experimental results are shown in Figure 4. It can be observed that DTR(U) performs quite well compared to the other two variations based on top-down sampling when the number of negative samples is small. However, when the negative sample size becomes relatively large, the other two variations consistently outperform DTR(U). This indicates that while uniform sampling is effective due to its extensive exploration and lack of bias, top-down sampling becomes advantageous with larger sample sizes by leveraging a sampling distribution whose probabilities are closer to the true softmax probabilities, leading to a more accurate estimation of the gradient for the multi-class cross-entropy loss. Moreover, notice that DTR(T-RL) outperforms DTR(T) consistently across four datasets, it validates the effectiveness of the modified loss with the rectified label again. As the number of negative samples increases, the $F$-$measure@20$ values of the three variations of DTR all show a gradual improvement and eventually converge. When the number of negative samples reaches $50$ or more, the performance of all three DTR variations stabilizes, indicating that DTR’s performance is relatively unaffected by the number of negative samples. This effectively addresses RQ-5, confirming that our proposed DTR is insensitive to the number of negative samples once it reaches a suitable threshold.
(a) Movie
(b) MIND
(c) Amazon
(d) Tmall
Figure 4: The $F\text{-}measure@20$ of DTR with varying numbers of negative samples across four datasets.
9 Conclusion
The existing tree-based recommendation systems regard the preference model training task as a binary-class classification problem. Their training mode suffers from the gap between training and prediction due to the lack of competition among nodes at the same layer. To address this problem, We develop a layer-wise training mode that regrades the training task as a multi-class classification problem and we propose a tree learning method that is compatible with the multi-class training mode. Through further theoretical analysis, we identify the suboptimality of the multi-class classification training mode under beam search and propose a rectification method to address this issue. Given the significant time cost of calculating softmax probabilities with large-scale datasets, we developed a sampling strategy based on the sampled softmax theory to estimate the softmax gradient, thus accelerating the training process. Finally, we provide a generalization analysis of the deep tree-based retriever, demonstrating the great generalization capability of our proposed methods. The experimental results validate the effectiveness of our proposed DTR.
Appendix A Proofs of Lemma, proposition, and Theorem
A.1 Proof of Proposition 2
Proof.
By taking $\psi(\boldsymbol{x})=\sum_{m=1}^{M}\boldsymbol{x}_{m}\log\boldsymbol{x}_{m}$ and $g(\boldsymbol{o})_{m}=\exp\boldsymbol{o}_{m}/\sum_{k=1}^{M}\exp\boldsymbol{o}_{k}$, we have
$D_{\psi}$ $\left(\boldsymbol{e}^{(i)}(w),g(\boldsymbol{o})\right)$
$=\psi\left(\boldsymbol{e}^{(i)}(w)\right)-\psi\left(g(\boldsymbol{o})\right)-\nabla\psi(g(\boldsymbol{o}))^{T}\left(\boldsymbol{e}^{(i)}(w)-g(\boldsymbol{o})\right)$
$=w\log w-\sum_{m=1}^{M}g(\boldsymbol{o})_{m}\log g(\boldsymbol{o})_{m}$
$\quad-\Big{<}\left(\log g(\boldsymbol{o})_{1}+1,\dots,\log g(\boldsymbol{o})_{i}+1,\dots,\log g(\boldsymbol{o})_{m}+1\right),$
$\quad\quad\quad\left(-g(\boldsymbol{o})_{1},\dots,w-g(\boldsymbol{o})_{i},\dots,-g(\boldsymbol{o})_{m}\right)\Big{>}$
$=w\log w-\sum_{m=1}^{M}g(\boldsymbol{o})_{m}\log g(\boldsymbol{o})_{m}$
$\quad-(-\sum_{m=1}^{M}g(\boldsymbol{o})_{m}\log g(\boldsymbol{o})_{m}-1+w\log g(\boldsymbol{o})_{i}+w)$
$=-w\log g(\boldsymbol{o})_{i}+w\log w+w-1=\widetilde{\mathcal{L}}(\boldsymbol{e}^{(i)}(w),\boldsymbol{o}).$
A.2 Proof of Lemma 1
Proof.
We complete the proof by contradiction. When $c(j+1)\leq k$, $\operatornamewithlimits{argTopk}\limits_{i\in[c(j+1)]}\mathbb{E}[\tilde{\boldsymbol{z}}^{j+1}_{i}]=\{[c(j+1)]\}=\mathcal{B}^{j+1}(u)$, so we only need to consider the case where $c(j+1)>k$.
Denote $\operatornamewithlimits{argTopk}\limits_{i\in[c(j+1)]}\mathbb{E}[\tilde{\boldsymbol{z}}^{j+1}_{i}]$ as $\{\mathcal{C}^{j+1}_{1},\mathcal{C}^{j+1}_{2},\dots,\mathcal{C}^{j+1}_{r}\}$ without loss of generality, where $r$ is a natural number and each $\mathcal{C}_{i}^{j+1}$ is a set of $k$ elements. Suppose $\mathcal{B}^{j+1}(u)\notin\operatornamewithlimits{argTopk}\limits_{i\in[c(j+1)]}\mathbb{E}[\tilde{\boldsymbol{z}}^{j+1}_{i}]$, we can make the following derivation:
$\mathcal{B}$ (${}^{j+1}(u)\notin\operatornamewithlimits{argTopk}\limits_{i\in[c(j+1)]}\mathbb{E}[\tilde{\boldsymbol{z}}^{j+1}_{i}]$
$\Longrightarrow\mathcal{B}^{j+1}(u)\neq\mathcal{C}^{j+1}_{i}(1\leq i\leq r)$
$\Longrightarrow\exists i_{1},i_{2}\in\left[c(j+1)\right]\ s.t.\ i_{1}\in\mathcal{B}^{j+1}(u)\wedge i_{1}\notin\mathcal{C}^{j+1}_{i}$
$\quad\quad\quad\quad\quad\quad\quad\quad\wedge i_{2}\notin\mathcal{B}^{j+1}(u)\wedge i_{2}\in\mathcal{C}_{i_{0}}^{j+1}\left(i_{0}\in[r]\right)$
$\Longrightarrow\mathbb{E}[\tilde{\boldsymbol{z}}^{j+1}_{i_{1}}]<\mathbb{E}[\tilde{\boldsymbol{z}}^{j+1}_{i_{2}}]\wedge o^{j+1}_{i_{1}}\geq o^{j+1}_{i_{2}}$
$\Longrightarrow i_{1}\in\widetilde{\mathcal{B}}^{j+1}(u)\wedge i_{2}\notin\widetilde{\mathcal{B}}^{j+1}(u)$
$\Longrightarrow\delta(\rho^{j}(n_{i_{1}}^{j+1}))\in\mathcal{B}^{j}(u)\wedge\delta(\rho^{j}(n_{i_{2}}^{j+1}))\notin\mathcal{B}^{j}(u).$
Let’s denote $\mathcal{B}^{j}(u)$ as $\{{s_{1}},{s_{2}},\dots,{s_{k}}\}$, where the elements satisfy $o^{j}_{s_{1}}\geq o^{j}_{s_{2}}\geq\dots\geq o^{j}_{s_{k}}$. Notice that $R(\boldsymbol{o}^{j},\mathbb{E}[\widetilde{\boldsymbol{z}}^{j}])$ and $\mathbb{E}[\widetilde{\boldsymbol{z}}^{j}_{s_{i}}]=\max\limits_{n^{\prime}\in\mathfrak{L}(n_{s_{1}}^{j})}\eta_{\pi(n^{\prime})}$, we have:
$\max\limits_{n^{\prime}\in\mathfrak{L}(n_{s_{1}}^{j})}\eta_{\pi(n^{\prime})}\geq\max\limits_{n^{\prime}\in\mathfrak{L}(n_{s_{2}}^{j})}\eta_{\pi(n^{\prime})}\geq\dots\geq\max\limits_{n^{\prime}\in\mathfrak{L}(n_{s_{k}}^{j})}\eta_{\pi(n^{\prime})}.$ (35)
For each node $n_{s_{i}}^{j}$, we then consider its node child $n^{j+1}_{c_{i}}\in\mathcal{D}(n_{s_{i}}^{j})$ where $c_{i}=\operatornamewithlimits{argmax}\limits_{i\in\{\delta(n)\mid n\in\mathcal{D}(n^{j}_{s_{i}})\}}\mathbb{E}[\tilde{\boldsymbol{z}}^{j+1}_{i}]$, and the following equation
$\max\limits_{n^{\prime}\in\mathfrak{L}(n^{j+1}_{c_{i}})}\eta_{\pi(n^{\prime})}=\max\limits_{n^{\prime}\in\mathfrak{L}(n^{j}_{s_{i}})}\eta_{\pi(n^{\prime})}$ (36)
holds for $1\leq i\leq k$. Combine the E.q. (35) and E.q. (36), we can get:
$\max\limits_{n^{\prime}\in\mathfrak{L}(n^{j+1}_{c_{1}})}\eta_{\pi(n^{\prime})}\geq\max\limits_{n^{\prime}\in\mathfrak{L}(n^{j+1}_{c_{2}})}\eta_{\pi(n^{\prime})}\geq\cdots\geq\max\limits_{n^{\prime}\in\mathfrak{L}(n^{j+1}_{c_{k}})}\eta_{\pi(n^{\prime})}.$
Then, for the $(j\text{+}1)$-th layer, with $R(\boldsymbol{o}^{j+1},\mathbb{E}[\widetilde{\boldsymbol{z}}^{j+1}])$ and $\mathbb{E}[\widetilde{\boldsymbol{z}}^{j+1}_{c_{i}}]=\max\limits_{n^{\prime}\in\mathfrak{L}(n_{c_{i}}^{j+1})}\eta_{\pi(n^{\prime})}$, we have $o^{j+1}_{c_{1}}\geq o^{j+1}_{c_{2}}\geq\dots\geq o^{j+1}_{c_{k}}$, which means that the rank of $o^{j+1}_{c_{k}}$ in the elements of $\boldsymbol{o}^{j+1}$ is not lower than $k$. And because $i_{1}\in\mathcal{B}^{j+1}(u)$, we can obtain the followings:
$o^{j+1}_{i_{1}}\geq o^{j+1}_{c_{k}}$ $\Longrightarrow\mathbb{E}[\tilde{\boldsymbol{z}}^{j+1}_{i_{1}}]\geq\mathbb{E}[\tilde{\boldsymbol{z}}^{j+1}_{c_{k}}]\Longrightarrow\mathbb{E}[\tilde{\boldsymbol{z}}^{j+1}_{i_{2}}]>\mathbb{E}[\tilde{\boldsymbol{z}}^{j+1}_{c_{k}}]$
$\Longrightarrow\max\limits_{n^{\prime}\in\mathfrak{L}(n_{i_{2}}^{j+1})}\eta_{\varphi(n^{\prime})}>\max\limits_{n^{\prime}\in\mathfrak{L}(n_{c_{k}}^{j+1})}\eta_{\varphi(n^{\prime})}.$ (37)
Note that $i_{\rho}\triangleq\delta(\rho^{j}(n_{i_{2}}^{j+1}))\notin\mathcal{B}^{j}(u)$, $s_{k}\in\mathcal{B}^{j}(u)$, and $\mathcal{B}^{j}(u)\in\operatornamewithlimits{argTopk}\limits_{i\in[c(j)]}\mathbb{E}[\tilde{\boldsymbol{z}}^{j}_{i}]$, we can have $\mathbb{E}[\tilde{\boldsymbol{z}}^{j}_{s_{k}}]\geq\mathbb{E}[\tilde{\boldsymbol{z}}^{j}_{i_{\rho}}]$. Therefore, we have
$\begin{split}\max\limits_{n^{\prime}\in\mathfrak{L}(n^{j}_{s_{k}})}\eta_{\pi(n^{\prime})}\geq\max\limits_{n^{\prime}\in\mathfrak{L}(n^{j}_{i_{\rho}})}\eta_{\pi(n^{\prime})}\geq\max\limits_{n^{\prime}\in\mathfrak{L}(n^{j+1}_{i_{2}})}\eta_{\pi(n^{\prime})},\end{split}$ (38)
where the last inequality holds as the $\mathfrak{L}(n^{j+1}_{i_{2}})\subseteq\mathfrak{L}(n^{j}_{i_{\rho}})$. Combine the E.q. (36) and E.q. (38), we can get:
$\max\limits_{n^{\prime}\in\mathfrak{L}(n^{j+1}_{c_{k}})}\eta_{\pi(n^{\prime})}\geq\max\limits_{n^{\prime}\in\mathfrak{L}(n^{j+1}_{i_{2}})}\eta_{\pi(n^{\prime})}.$ (39)
E.q. (37) and E.q. (39) induces the contradiction, so the proposition holds. ∎
A.3 Proof of Proposition 3
Proof.
Let’s consider the retrieval process along the tree for a given user $u$ and a given beam size $k$.
Let $j^{*}=\operatornamewithlimits{argmin}\limits_{\{j:c(j)>k\}}c(j)$. When $11\leq j\leq j^{*}-$ 1, $\mathcal{B}^{j}(u)=\{[c(j)]\}\in\operatornamewithlimits{argTopk}\limits_{i\in[c(j)]}\mathbb{E}[\tilde{\boldsymbol{z}}^{j}_{i}]$. When $j\geq j^{*}$, by recursively applying Lemma 1 from $j^{*}$ to $H-1$, we ultimately obtain $\mathcal{B}^{H}(u)\in\operatornamewithlimits{argTopk}\limits_{i\in[c(H)]}\mathbb{E}[\tilde{\boldsymbol{z}}^{H}_{i}]$. Notice that $\mathbb{E}[\tilde{\boldsymbol{z}}_{i}^{H}]=\eta_{\pi(n_{i}^{H})}$, that is to say $\mathcal{B}^{H}(u)\in\operatornamewithlimits{argTopk}\limits_{i\in[c(H)]}\eta_{\pi(n_{i}^{H})}$, meaning the final retrieved item set $\widehat{\mathcal{Y}}=\{\pi(n_{i}^{H})|i\in\mathcal{B}^{H}(u)\}\in\operatornamewithlimits{argTopk}\limits_{y\in\mathcal{Y}}\eta_{y}(u)$. The above analysis can be performed on any $u$ and $k$, so the tree model is top-$k$ retrieval Bayes optimal. ∎
A.4 Proof of Theorem 2
Proof.
We use induction to prove this theorem. Assuming the sampling probability of node $n_{i}^{j}$ according to top-down sampling is $p(n_{i}^{j})=\frac{\exp{o_{i}^{j}}}{\sum_{k=1}^{c(j)}\exp{o_{k}^{j}}}$. When $j=0$, there is only root node $n_{1}^{0}$ in the 0th layer, and it is sampled with probability $1$, equaling to its softmax probability; when $j=1$, notice that $\mathcal{D}(n_{1}^{0})=\mathcal{N}^{j}$, so the assumption holds due to the expanding probability of top-down sampling (i.e., E.q.(27)). For any $n_{i}^{j}$’s child node $n_{c}^{j+1}$ in layer $j+1$, the sampling probability of node $n_{c}^{j+1}$ is calculated as:
$p(n_{c}^{j+1})$ $=p(n_{i}^{j})\cdot\frac{\exp{o_{c}^{j+1}}}{\sum\limits_{k\in\left\{\delta(n)\mid n\in\mathcal{D}(n_{i}^{j})\right\}}\exp{o_{k}^{j+1}}}$ (40)
$=\frac{\exp{o_{i}^{j}}}{\sum_{k=1}^{c(j)}\exp{o_{k}^{j}}}\cdot\frac{\exp{o_{c}^{j+1}}}{\sum\limits_{k\in\left\{\delta(n)\mid n\in\mathcal{D}(n_{i}^{j})\right\}}\exp{o_{k}^{j+1}}}$
According to E.q. (28), and with the assumption that the proportionality coefficients within the same layer are equal, we have
$\sum_{k=1}^{c(j)}\exp{o_{k}^{j}}$ $\propto\sum_{k=1}^{c(j)}\sum\limits_{k^{\prime}\in\left\{\delta(n)\mid n\in\mathcal{D}(n_{k}^{j})\right\}}\exp{o_{k^{\prime}}^{j+1}}$ (41)
$=\sum_{k=1}^{c(j+1)}\exp{o_{k}^{j+1}}.$
Therefore, by substituting equations E.q. (28) and E.q. (41) into E.q. (40), we can obtain:
$p(n_{c}^{j+1})=\frac{\exp{o_{c}^{j+1}}}{\sum_{k=1}^{c(j+1)}\exp{n_{k}^{j+1}}}.$
That is, when sampling to the $(j\text{+}1)$-th level in the top-down sampling, the probability of sampling the corresponding node is equivalent to the softmax probability in the $(j\text{+}1)$-th level. As $j$ increases, when reaching the layer where the leaf nodes are located, this conclusion also holds. ∎
A.5 Proof of Lemma  2
Proof.
For any vectors $\boldsymbol{v},\boldsymbol{u}_{i}\in\mathbb{R}^{d},\|\boldsymbol{v}\|_{1}\leq B_{v}$, notice the following inequality:
$\sup_{\boldsymbol{v}}\sum_{i}\boldsymbol{v}^{\mathsf{T}}\boldsymbol{u}_{i}\leq B_{v}\max_{j\in[d]}\left|\sum_{i}\boldsymbol{e}_{j}\boldsymbol{u}_{i}\right|\leq\sum_{i}B_{v}\max_{\begin{subarray}{c}j\in[d]\\ s\in\{-1,1\}\end{subarray}}s\boldsymbol{e}_{j}\boldsymbol{u}_{i}.$ (42)
By applying the E.q. (42), we can get:
^$\hat{\mathcal{R}}_{m}$ $(\mathcal{M})=\mathbb{E}_{\sigma}\left[\sup\limits_{f\in\mathcal{M}}\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}f(u_{i},n_{i})\right]$
$=\mathbb{E}_{\sigma}\Bigg{[}\sup\limits_{\boldsymbol{z},\boldsymbol{w}_{n},\{\boldsymbol{W}_{k}\}^{L}_{k=1}}\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\Big{<}\boldsymbol{W}_{L},\phi_{L-1}\circ$
$\quad\quad\phi_{L-2}\circ\dots\circ\phi_{1}\left(\boldsymbol{z}^{(u_{i})}_{1};\boldsymbol{z}^{(u_{i})}_{2};\dots;\boldsymbol{z}^{(u_{i})}_{K^{\prime}};\boldsymbol{w}_{n_{i}}\right)\Big{>}\Bigg{]}$
$\underset{(a)}{\leq}\|\boldsymbol{W}_{L}\|_{1}\mathbb{E}_{\sigma}\Bigg{[}\sup\limits_{\begin{subarray}{c}s\in\{-1,1\},j\in[d_{L-1}]\\ \boldsymbol{z},\boldsymbol{w}_{n},\{\boldsymbol{W}_{k}\}_{k=1}^{L-1}\end{subarray}}\frac{1}{m}\sum_{i=1}^{m}s\sigma_{i}\Big{<}\boldsymbol{e}_{j},\phi_{L-1}\circ$
$\quad\quad\phi_{L-2}\circ\dots\circ\phi_{1}\left(\boldsymbol{z}^{(u_{i})}_{1};\boldsymbol{z}^{(u_{i})}_{2};\dots;\boldsymbol{z}^{(u_{i})}_{K^{\prime}};\boldsymbol{w}_{n_{i}}\right)\Big{>}\Bigg{]}$
$\underset{(b)}{\leq}c_{\phi}\|\boldsymbol{W}_{L}\|_{1}\mathbb{E}_{\sigma}\Bigg{[}\sup\limits_{\begin{subarray}{c}s\in\{-1,1\},j\in[d_{L-1}]\\ \boldsymbol{z},\boldsymbol{w}_{n},\{\boldsymbol{W}_{k}\}_{k=1}^{L-1}\end{subarray}}\frac{1}{m}\sum_{i=1}^{m}s\sigma_{i}\Big{<}\boldsymbol{e}_{j},\boldsymbol{W}_{L-1}\cdot$
$\quad\quad\phi_{L-2}\circ\dots\circ\phi_{1}\left(\boldsymbol{z}^{(u_{i})}_{1};\boldsymbol{z}^{(u_{i})}_{2};\dots;\boldsymbol{z}^{(u_{i})}_{K^{\prime}};\boldsymbol{w}_{n_{i}}\right)\Big{>}\Bigg{]}$
$\underset{(c)}{\leq}c_{\phi}\|\boldsymbol{W}_{L}\|_{1}\|\boldsymbol{W}_{L-2}\|_{1}\mathbb{E}_{\sigma}\Bigg{[}\sup\limits_{\begin{subarray}{c}s\in\{-1,1\},j\in[d_{L-2}]\\ \boldsymbol{z},\boldsymbol{w}_{n},\{\boldsymbol{W}_{k}\}_{k=1}^{L-1}\end{subarray}}\frac{1}{m}\sum_{i=1}^{m}s\sigma_{i}$
$\quad\quad\Big{<}\boldsymbol{e}_{j},\phi_{L-2}\circ\dots\circ\phi_{1}\left(\boldsymbol{z}^{(u_{i})}_{1};\boldsymbol{z}^{(u_{i})}_{2};\dots;\boldsymbol{z}^{(u_{i})}_{K^{\prime}};\boldsymbol{w}_{n_{i}}\right)\Big{>}\Bigg{]}$
$\underset{(d)}{\leq}2c_{\phi}^{L-1}\prod_{k=1}^{L}\|\boldsymbol{W}_{k}\|_{1}\mathbb{E}_{\sigma}\Bigg{[}\sup\limits_{\begin{subarray}{c}\boldsymbol{z},\boldsymbol{w}_{n},\\ j\in[(K^{\prime}+1)d]\end{subarray}}\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\Big{<}\boldsymbol{e}_{j},$
$\quad\quad\left(\boldsymbol{z}^{(u_{i})}_{1};\boldsymbol{z}^{(u_{i})}_{2};\dots;\boldsymbol{z}^{(u_{i})}_{K^{\prime}};\boldsymbol{w}_{n_{i}}\right)\Big{>}\Bigg{]}$
$\leq 2c_{\phi}^{L-1}\prod_{k=1}^{L}\|\boldsymbol{W}_{k}\|_{1}\Bigg{(}\underbrace{\mathbb{E}_{\sigma}\Big{[}\sup_{j\in[d],\boldsymbol{w}_{n}}\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\left<\boldsymbol{e}_{j},\boldsymbol{w}_{n_{i}}\right>\Big{]}}_{I_{1}}$
$\quad+\sum_{t=1}^{K^{\prime}}\ \underbrace{\mathbb{E}_{\sigma}\Big{[}\sup_{j\in[d],\boldsymbol{z}_{t}}\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\left<\boldsymbol{e}_{j},\boldsymbol{z}_{t}^{(u_{i})}\right>\Big{]}}_{I_{2}}\Bigg{)}$ (43)
where (a) use E.q. (42) as $\boldsymbol{W}_{L}\in\mathbb{R}^{1\times d_{L-1}}$ is actually a vector; (b) holds since $\sigma$ is applied element wise, we can bring $\boldsymbol{e}^{T}_{j}$ inside the function and the use of contraction inequality [67]; (c) use E.q (42) again as $\boldsymbol{e}_{j}^{\mathsf{T}}\boldsymbol{W}_{L-1}$ is still a vector; (d) holds by applying E.q. (42) recursively and utilizing the fact that ^$\hat{\mathcal{R}}_{m}(\mathcal{F}\cup-\mathcal{F})\leq 2\hat{\mathcal{R}}_{m}(\mathcal{F})$.
As the term $I_{1}$, using Cauchy-Schwartz inequality and Jensen inequality, we have:
$I_{1}$ $\leq\frac{1}{m}\mathbb{E}_{\sigma}\left[\left\|\sum_{i=1}^{m}\sigma_{i}\boldsymbol{w}_{n_{i}}\right\|_{2}\right]\leq\frac{1}{m}\left(\mathbb{E}_{\sigma}\left[\left\|\sum_{i=1}^{m}\sigma_{i}\boldsymbol{w}_{n_{i}}\right\|_{2}^{2}\right]\right)^{1/2}$
$=\frac{1}{m}\left(\sum_{i=1}^{m}\left\|\boldsymbol{w}_{n_{i}}\right\|^{2}\right)^{1/2}\leq\frac{B_{0}}{\sqrt{m}}.$ (44)
As the term $I_{2}$, we have:
$I_{2}$ $=\frac{1}{m}\mathbb{E}_{\sigma}\left[\sup_{j\in[d],\boldsymbol{w}}\sum_{i=1}^{m}\sigma_{i}\left\langle\boldsymbol{e}_{j},\sum_{k\in T_{t}}w_{k}^{(u_{i})}\boldsymbol{a}_{k}^{(u_{i})}\right\rangle\right]$
$\leq\frac{1}{m}\sum_{k\in T_{t}}\mathbb{E}_{\sigma}\left[\sup_{j\in[d],\boldsymbol{w}}\sum_{i=1}^{m}\sigma_{i}\left\langle\boldsymbol{e}_{j},w_{k}^{(u_{i})}\boldsymbol{a}_{k}^{(u_{i})}\right\rangle\right]$
$=\frac{1}{m}\sum_{k\in T_{t}}\mathbb{E}_{\sigma}\Bigg{[}\sup_{j\in[d],\boldsymbol{w}}\sum_{i=1}^{m}\sigma_{i}\phi\Big{(}\boldsymbol{W}_{w}^{(2)}\phi\Big{(}\boldsymbol{W}_{w}^{(1)}$
$\quad\quad\quad\left[\boldsymbol{a}_{k}^{(u_{i})};\boldsymbol{a}_{k}^{(u_{i})}\odot\boldsymbol{w}_{n_{i}};\boldsymbol{w}_{n_{i}}\right]\Big{)}\Big{)}\left\langle\boldsymbol{e}_{j},\boldsymbol{a}_{k}^{(u_{i})}\right\rangle\Bigg{]}$
$\leq\frac{1}{m}c_{\phi}^{2}\left\|\boldsymbol{W}_{w}^{(1)}\right\|_{1}\left\|\boldsymbol{W}_{w}^{(2)}\right\|_{1}\sum_{k\in T_{t}}\mathbb{E}_{\sigma}\Bigg{[}\sup_{\begin{subarray}{c}\boldsymbol{w},j\in[d]\\ j^{\prime}\in[3d]\end{subarray}}\sum_{i=1}^{m}\sigma_{i}$
$\quad\quad\left\langle\boldsymbol{e}_{j^{\prime}},\left[\boldsymbol{a}_{k}^{(u_{i})};\boldsymbol{a}_{k}^{(u_{i})}\odot\boldsymbol{w}_{n_{i}};\boldsymbol{w}_{n_{i}}\right]\right\rangle\left\langle\boldsymbol{e}_{j},\boldsymbol{a}_{k}^{(u_{i})}\right\rangle\Bigg{]}.$ (45)
Furthermore, we can get
$\mathbb{E}_{\sigma}\Bigg{[}\sup_{\begin{subarray}{c}\boldsymbol{w},j\in[d]\\ j^{\prime}\in[3d]\end{subarray}}\sum_{i=1}^{m}\sigma_{i}\left\langle\boldsymbol{e}_{j^{\prime}},\left[\boldsymbol{a}_{k}^{(u_{i})};\boldsymbol{a}_{k}^{(u_{i})}\odot\boldsymbol{w}_{n_{i}};\boldsymbol{w}_{n_{i}}\right]\right\rangle\left\langle\boldsymbol{e}_{j},\boldsymbol{a}_{k}^{(u_{i})}\right\rangle\Bigg{]}$ (46)
$\leq\mathbb{E}_{\sigma}\left[\sup_{j\in[d],j^{\prime}\in[d]}\sum_{i=1}^{m}\sigma_{i}\left\langle\boldsymbol{e}_{j^{\prime}},\boldsymbol{a}_{k}^{(u_{i})}\right\rangle\left\langle\boldsymbol{e}_{j},\boldsymbol{a}_{k}^{(u_{i})}\right\rangle\right]$
$\quad+\mathbb{E}_{\sigma}\left[\sup_{j\in[d],j^{\prime}\in[d],\boldsymbol{w}}\sum_{i=1}^{m}\sigma_{i}\left\langle\boldsymbol{e}_{j^{\prime}},\boldsymbol{a}_{k}^{(u_{i})}\odot\boldsymbol{w}_{n_{i}}\right\rangle\left\langle\boldsymbol{e}_{j},\boldsymbol{a}_{k}^{(u_{i})}\right\rangle\right]$
$\quad+\mathbb{E}_{\sigma}\left[\sup_{j\in[d],j^{\prime}\in[d],\boldsymbol{w}}\sum_{i=1}^{m}\sigma_{i}\left<\boldsymbol{e}_{j^{\prime}},\boldsymbol{w}_{n_{i}}\right>\left<\boldsymbol{e}_{j},\boldsymbol{a}_{k}^{(u_{i})}\right>\right]$
$=I_{3}+I_{4}+I_{5}.$
As the term $I_{3}$, notice that
$\sum_{i=1}^{m}\sigma_{i}\left<\boldsymbol{e}_{j^{\prime}},\boldsymbol{a}_{k}^{(u_{i})}\right>\left<\boldsymbol{e}_{j},\boldsymbol{a}_{k}^{(u_{i})}\right>=\sum_{i=1}^{m}\sigma_{i}\boldsymbol{e}_{j^{\prime}}^{\mathsf{T}}\boldsymbol{P}_{a}^{(u_{i})}\boldsymbol{e}_{j}$
$=\sum_{i=1}^{m}\sigma_{i}\mathrm{Tr}\left(\boldsymbol{e}_{j}\boldsymbol{e}_{j^{\prime}}^{\mathsf{T}}\boldsymbol{P}_{a}^{(u_{i})}\right)=\mathrm{Tr}\left(\boldsymbol{e}_{j}\boldsymbol{e}_{j^{\prime}}^{\mathsf{T}}\left(\sum_{i=1}^{m}\sigma_{i}\boldsymbol{P}_{a}^{(u_{i})}\right)\right)$
$=\left\langle\boldsymbol{e}_{j}\boldsymbol{e}_{j^{\prime}}^{\mathsf{T}},\sum_{i=1}^{m}\sigma_{i}\boldsymbol{P}_{a}^{(u_{i})}\right\rangle_{F},$
where $\boldsymbol{P}_{a}^{(u_{i})}=\boldsymbol{a}_{k}^{(u_{i})}{\boldsymbol{a}_{k}^{(u_{i})}}^{\mathsf{T}}$. Then, we can get
$I_{3}$ $=\mathbb{E}_{\sigma}\left[\sup_{j\in[d],j^{\prime}\in[d]}\sum_{i=1}^{m}\sigma_{i}\left\langle\boldsymbol{e}_{j^{\prime}},\boldsymbol{a}_{k}^{(u_{i})}\right\rangle\left\langle\boldsymbol{e}_{j},\boldsymbol{a}_{k}^{(i)}\right\rangle\right]$ (47)
$=\mathbb{E}_{\sigma}\left[\sup_{j\in[d],j^{\prime}\in[d]}\left\langle\boldsymbol{e}_{j}\boldsymbol{e}_{j^{\prime}}^{\mathsf{T}},\sum_{i=1}^{m}\sigma_{i}\boldsymbol{P}_{a}^{(u_{i})}\right\rangle_{F}\right]$
$\leq\mathbb{E}_{\sigma}\left[\left\|\sum_{i=1}^{m}\sigma_{i}\boldsymbol{P}_{a}^{(u_{i})}\right\|_{F}\right]=\sqrt{\sum_{i=1}^{m}\left\|\boldsymbol{P}_{a}^{(u_{i})}\right\|_{F}^{2}}\leq\sqrt{mB_{a}^{4}}.$
As the term $I_{4}$, use the same analysis technique as for $I_{3}$, we can get
$I_{4}=\mathbb{E}_{\sigma}\left[\sup_{j\in[d],j^{\prime}\in[d],\boldsymbol{w}}\sum_{i=1}^{m}\sigma_{i}\left\langle\boldsymbol{e}_{j^{\prime}},\boldsymbol{a}_{k}^{(u_{i})}\odot\boldsymbol{w}_{n_{i}}\right\rangle\left\langle\boldsymbol{e}_{j},\boldsymbol{a}_{k}^{(u_{i})}\right\rangle\right]$ (48)
$=\mathbb{E}_{\sigma}\left[\sup_{j\in[d],j^{\prime}\in[d],\boldsymbol{w}}\sum_{i=1}^{m}\sigma_{i}\left\langle\boldsymbol{e}_{j^{\prime}}\odot\boldsymbol{w}_{n_{i}},\boldsymbol{a}_{k}^{(u_{i})}\right\rangle\left\langle\boldsymbol{e}_{j},\boldsymbol{a}_{k}^{(u_{i})}\right\rangle\right]$
$=\mathbb{E}_{\sigma}\left[\sup_{j\in[d],j^{\prime}\in[d],\boldsymbol{w}}\sum_{i=1}^{m}\sigma_{i}\left\langle\boldsymbol{e}_{j}\boldsymbol{e}_{j^{\prime}}^{\mathsf{T}}\odot\boldsymbol{w}_{n_{i}}^{\mathsf{T}},\boldsymbol{P}_{a}^{(u_{i})}\right\rangle_{F}\right]$
$=\mathbb{E}_{\sigma}\left[\sup_{j\in[d],j^{\prime}\in[d],\boldsymbol{w}}\sum_{n\in\mathcal{N}}\sum_{i:n_{i}=n}\sigma_{i}\left\langle\boldsymbol{e}_{j}\boldsymbol{e}_{j^{\prime}}^{\mathsf{T}}\odot\boldsymbol{w}_{n}^{\mathsf{T}},\boldsymbol{P}_{a}^{(u_{i})}\right\rangle_{F}\right]$
$\leq\sum_{n\in\mathcal{N}}\mathbb{E}_{\sigma}\left[\sup_{j\in[d],j^{\prime}\in[d]}\left\langle\boldsymbol{e}_{j}\boldsymbol{e}_{j^{\prime}}^{\mathsf{T}}\odot\boldsymbol{w}_{n}^{\mathsf{T}},\sum_{i:n_{i}=n}\sigma_{i}\boldsymbol{P}_{a}^{(u_{i})}\right\rangle_{F}\right]$
$\leq B_{0}\sum_{n\in\mathcal{N}}\mathbb{E}_{\sigma}\left[\left\|\sum_{i:n_{i}=n}\sigma_{i}\boldsymbol{P}_{a}^{(u_{i})}\right\|_{F}\right]$
$\leq B_{0}B_{a}^{2}\sum_{n\in\mathcal{N}}\sqrt{\left|\{i:n_{i}=n\}\right|}\leq B_{0}B_{a}^{2}\sqrt{\frac{2B-1}{B-1}|\mathcal{Y}|m},$
where the last inequality holds as $\sum_{n\in\mathcal{N}}\left|\{i:n_{i}=n\}\right|=m$, $|\mathcal{N}|=\sum_{j=0}^{\lceil\log_{B}|\mathcal{Y}|\rceil-1}B^{j}+|\mathcal{Y}|\leq\frac{2B-1}{B-1}|\mathcal{Y}|$ for the $B$-ary tree, and use the Cauchy-Schwartz inequality.
As the term $I_{5}$, use the same technique as for $I_{4}$, we have
$I_{5}$ $=\mathbb{E}_{\sigma}\left[\sup_{j\in[d],j^{\prime}\in[d],\boldsymbol{w}}\sum_{i=1}^{m}\sigma_{i}\left<\boldsymbol{e}_{j^{\prime}},\boldsymbol{w}_{n_{i}}\right>\left<\boldsymbol{e}_{j},\boldsymbol{a}_{k}^{(u_{i})}\right>\right]$
$=\mathbb{E}_{\sigma}\left[\sup_{j\in[d],j^{\prime}\in[d],\boldsymbol{w}}\sum_{n\in\mathcal{N}}\sum_{i:n_{i}=n}\sigma_{i}\left<\boldsymbol{e}_{j^{\prime}},\boldsymbol{w}_{n_{i}}\right>\left<\boldsymbol{e}_{j},\boldsymbol{a}_{k}^{(u_{i})}\right>\right]$
$\leq\sum_{n\in\mathcal{N}}\mathbb{E}_{\sigma}\left[\sup_{j\in[d],j^{\prime}\in[d]}\left<\boldsymbol{e}_{j^{\prime}},\boldsymbol{w}_{n}\right>\left<\boldsymbol{e}_{j},\sum_{i:n_{i}=n}\sigma_{i}\boldsymbol{a}_{k}^{(u_{i})}\right>\right]$
$\leq B_{0}\sum_{n\in\mathcal{N}}\mathbb{E}_{\sigma}\left[\left\|\sum_{i:n_{i}=n}\sigma_{i}\boldsymbol{a}_{k}^{(u_{i})}\right\|_{2}\right]\leq B_{0}B_{a}\sqrt{\frac{2B-1}{B-1}|\mathcal{Y}|m}.$ (49)
[…]
^$\begin{split}&\hat{\mathcal{R}}_{m}(\mathcal{M})=\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{M}}\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}f\left(u_{i},n_{i}\right)\right]\\ &\leq 2c_{\phi}^{L-1}\prod_{k=1}^{L}\|\boldsymbol{W}_{k}\|_{1}\Bigg{[}\frac{B_{0}}{\sqrt{m}}+c_{\phi}^{2}\left\|\boldsymbol{W}_{w}^{(1)}\right\|_{1}\left\|\boldsymbol{W}_{w}^{(2)}\right\|_{1}K\cdot\\ &\quad\left(\frac{B_{a}^{2}}{\sqrt{m}}+\frac{B_{0}B^{2}_{a}}{\sqrt{m}}\sqrt{\frac{2B-1}{B-1}|\mathcal{Y}|}+\frac{B_{0}B_{a}}{\sqrt{m}}\sqrt{\frac{2B-1}{B-1}|\mathcal{Y}|}\right)\Bigg{]}\\ &\leq\frac{2c_{\phi}^{L-1}B_{1}^{L}(B_{0}+KB_{w_{1}}+KB_{w_{2}}\mathcal{\tau})}{\sqrt{m}},\end{split}$
where $B_{w_{1}}=c_{\phi}^{2}B_{2}^{2}B^{2}_{a},B_{w_{2}}=c_{\phi}^{2}B_{0}B_{2}^{2}\left(B_{a}^{2}+B_{a}\right),\tau=\sqrt{(2B-1)|\mathcal{Y}|}/\sqrt{B-1}$. ∎
A.6 Proof of Lemma  3
Proof.
Given the sample set $S=\{u_{i},y_{i}\}_{i=1}^{m}$, we consider the loss function space $\mathcal{F}_{\ell}^{j}=\{(u,y)\rightarrow f_{\ell}^{j}(u,y)\}$, where function $f_{\ell}^{j}(u,y)=\widetilde{\mathcal{L}}_{j}(u,y)=-\bar{z}^{j}_{\delta^{j}(y)}\log\frac{\exp o^{j}_{\delta^{j}(y)}(u)}{\sum_{k=1}^{c(j)}\exp o^{j}_{k}(u)}$ and $o_{k}^{j}(u)=f_{Din}(u,n_{k}^{j})$. Due to $\left|f_{Din}\right|\leq B_{\mathcal{M}}$, we can get $\left|f_{\ell}^{j}\right|\leq 2B_{\mathcal{M}}+\log c(j)$. Then, by the Lemma 5, we have:
$\begin{split}\mathbb{E}_{(u,y)\sim\mathcal{D}}\left[\widetilde{\mathcal{L}}_{j}(u,y)\right]&\leq\frac{1}{m}\sum_{i=1}^{m}\widetilde{\mathcal{L}}_{j}(u_{i},y_{i})+2\hat{\mathcal{R}}_{m}(\mathcal{F}_{\ell}^{j},S)\\ &+\Big{(}4\log c(j)+8B_{\mathcal{M}}\Big{)}\sqrt{\frac{2\log\left(4/\delta\right)}{m}}.\end{split}$ (50)
For the empirical Rachemader complexity of $\mathcal{F}_{\ell}^{j}$, we have
^$\begin{split}&\hat{\mathcal{R}}_{m}(\mathcal{F}_{\ell}^{j},S)=\mathbb{E}_{\sigma}\left[\sup_{f_{\ell}^{j}\in\mathcal{F}_{\ell}^{j}}\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}f_{\ell}^{j}(u_{i},y_{i})\right]\\ &=\mathbb{E}_{\sigma}\left[\sup_{\begin{subarray}{c}f_{Din}\in\mathcal{M}\\ \bar{z}\in\{0,1\}\end{subarray}}\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\bar{z}^{j}_{\delta^{j}(y_{i})}\log\frac{\sum_{k=1}^{c(j)}\exp o^{j}_{k}(u_{i})}{\exp o_{\delta^{j}(y_{i})}^{j}(u_{i})}\right]\\ &=\mathbb{E}_{\sigma}\left[\sup_{\begin{subarray}{c}f_{Din}\in\mathcal{M}\\ z^{\prime}\triangleq 2\bar{z}-1\in\{-1,1\}\end{subarray}}\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\frac{z^{\prime}+1}{2}\log\frac{\sum_{k=1}^{c(j)}\exp o^{j}_{k}(u_{i})}{\exp o_{\delta^{j}(y_{i})}^{j}(u_{i})}\right]\\ &\leq\mathbb{E}_{\sigma}\left[\sup_{\begin{subarray}{c}f_{Din}\in\mathcal{M}\end{subarray}}\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}*-\log\frac{\exp o_{\delta^{j}(y_{i})}^{j}(u_{i})}{\sum_{k=1}^{c(j)}\exp o^{j}_{k}(u_{i})}\right]\\ &=\hat{\mathcal{R}}_{m}(\ell\circ\mathcal{M},S),\end{split}$ (51)
where the last equality holds because
$-\log\frac{\exp o_{\delta^{j}(y_{i})}^{j}(u_{i})}{\sum\limits_{k=1}\limits^{c(j)}\exp o^{j}_{k}(u_{i})}=-\sum_{m=1}^{c(j)}\mathbb{I}(m=\delta^{j}(y_{i}))\cdot\log\frac{\exp o^{j}_{m}(u_{i})}{\sum\limits_{k=1}\limits^{c(j)}\exp o_{k}^{j}(u_{i})}$
is a composition of logistic loss $\ell_{y}(\boldsymbol{o})=-\sum_{i}y_{i}\log\frac{\exp\boldsymbol{o}_{i}}{\sum_{j}\exp\boldsymbol{o}_{j}}$ and $f_{Din}$. Since the partial derivative of $\ell_{y}(\boldsymbol{o})$ w.r.t. each component is bounded by $1$, the logistic loss function $\ell$ is $1$-Lipschitz [68]. For the $j$-th layer, where the number of classes is $c(j)$, by Lemma 6, we have
^$\hat{\mathcal{R}}_{m}(\ell\circ\mathcal{M},S)\leq 2c(j)\hat{\mathcal{R}}_{m}(\mathcal{M},S).$ (52)
Combine the E.q. (50), E.q. (51) and E.q. (52), and utilize the Lemma 2, we can obtain the desired result. ∎
A.7 Proof of Lemma 4
Proof.
The $i$-th node is positive among $c(j)$ nodes in the $j$-th layer. The sample distribution is $Q^{j}$, and the corresponding sampling probability for node $n_{i^{\prime}}^{j}$ is $q_{i^{\prime}}^{j}$. Therefore, according to E.q. (22), the adjusted logits are calculated as follows:
$\boldsymbol{\hat{o}}^{j}_{i^{\prime}}=\left\{\begin{aligned} &\boldsymbol{{o}}^{j}_{i}-\ln(Mq^{j}_{i^{\prime}})\ \ \ \ if\ i^{\prime}\neq i\\ &\boldsymbol{{o}}^{j}_{i}-\ln(1)\quad\quad\ \ \ if\ i^{\prime}=i\end{aligned}\right.$ (53)
Then, we can have:
$\begin{split}&\ell^{j}_{softmax}(u,y)-\mathbb{E}_{\mathcal{I}^{\prime}_{M}}\left[\ell^{j}_{sampled\text{-}softmax}(u,y,\mathcal{I}^{\prime}_{M})\right]\\ &=-\log\frac{\exp\boldsymbol{o}^{j}_{i}}{\sum_{k=1}^{c(j)}\exp\boldsymbol{o}^{j}_{k}}-{\mathbb{E}}_{\mathcal{I}^{\prime}_{M}}\Big{[}-\log\frac{\exp\boldsymbol{\hat{o}}_{i}^{j}}{\sum\limits_{i^{\prime}\in\mathcal{I}^{\prime}_{M}\cup\{i\}}\exp\boldsymbol{\hat{o}}_{i^{\prime}}^{j}}\Big{]}\\ &=\mathbb{E}_{\mathcal{I}^{\prime}_{M}}\left[\log\Big{(}\sum_{k=1}^{c(j)}\exp\boldsymbol{o}^{j}_{k}\Big{)}-\log\left(\exp\boldsymbol{o}^{j}_{i}+\sum_{i^{\prime}\in\mathcal{I}^{\prime}_{M}}\frac{\exp\boldsymbol{o}^{j}_{i^{\prime}}}{Mq_{i^{\prime}}^{j}}\right)\right]\\ &=\mathbb{E}_{\mathcal{I}^{\prime}_{M}}\left[-\log\left(p_{i}^{j}+\frac{1}{M}\sum_{i^{\prime}\in\mathcal{I}^{\prime}_{M}}\frac{p_{i^{\prime}}^{j}}{q_{i^{\prime}}^{j}}\right)\right]\\ &=\mathbb{E}_{\mathcal{I}^{\prime}_{M}}\left[\log\frac{M}{\sum_{i^{\prime}\in\mathcal{I}^{\prime}_{M}}\Big{(}\frac{q_{i^{\prime}}^{j}}{p_{i^{\prime}}^{j}+p_{i}^{j}q_{i^{\prime}}^{j}}\Big{)}^{-1}}\right]\\ &\underset{(a)}{\leq}\mathbb{E}_{\mathcal{I}^{\prime}_{M}}\left[\log\left({\prod_{i^{\prime}\in\mathcal{I}^{\prime}_{M}}\frac{q_{i^{\prime}}^{j}}{p_{i^{\prime}}^{j}+p_{i}^{j}q_{i^{\prime}}^{j}}}\right)^{\frac{1}{M}}\right]\\ &=\frac{1}{M}\sum_{i^{\prime}\in\mathcal{I}^{\prime}_{M}}\mathbb{E}_{i^{\prime}\sim Q}\left[\log\frac{q_{i^{\prime}}^{j}}{p_{i^{\prime}}^{j}+p_{i}^{j}q_{i^{\prime}}^{j}}\right]\\ &\leq\frac{1}{M}\sum_{i^{\prime}\in\mathcal{I}^{\prime}_{M}}\mathbb{E}_{i^{\prime}\sim Q}\left[\log\left(\frac{q_{i^{\prime}}^{j}}{p_{i^{\prime}}^{j}}\right)\right]=D_{\text{KL}}(Q^{j}\|P^{j}),\end{split}$
where (a) uses Harmonic-Geometric inequality. ∎
A.8 Proof of Theorem 4
Proof.
By Lemma 4, we have
$\mathbb{E}_{(u,y)\sim\mathcal{D}}\left[\widetilde{\mathcal{L}}_{j}(u,y)\right]-\mathbb{E}_{(u,y),I_{M}^{\prime}}\left[\widehat{\mathcal{L}}_{j}(u,y)\right]$ (54)
$\leq\mathbb{E}_{(u,y)\sim\mathcal{D}}\left[D_{KL}(Q^{j}||P^{j})\right],$
For convenience, we omit the subscripts of the expectations and use $\mathbb{E}$ to denote the expectation taken over $u$, $y$, and $I_{M}^{\prime}$. From Lemma 5, we have:
[…]
where
[…]
$=\mathbb{E}_{\sigma}\left[\sup_{\begin{subarray}{c}f_{Din}\in\mathcal{M}\\ \bar{z}\in\{0,1\}\end{subarray}}\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\bar{z}^{j}_{\delta^{j}(y_{i})}\log\frac{\sum_{k\in\mathcal{I}^{\prime}_{M}\cup\{{\delta^{j}(y_{i})}\}}\exp o^{j}_{k}(u_{i})}{\exp o_{\delta^{j}(y_{i})}^{j}(u_{i})}\right].$
Using the same analysis method with E.q. (51) and E.q. (52) in the proof of Lemma 3, we can get
[…]
Appendix B Auxiliary Lemmas
Lemma 5 (Theorem 26.5(2) of [69]).
If the magnitude of loss function $l$ is bounded above by $c$, with probability greater than $1-\delta$ for all $h\in\mathcal{H}$, we have
$\begin{split}\mathbb{E}_{(x,y)\sim\mathcal{D}}[\ell(h(x),y)]&\leq\frac{1}{m}\sum_{i=1}^{m}\ell\left(h\left(x_{i}\right),y_{i}\right)\\ &+2\hat{\mathcal{R}}_{m}(\ell\circ\mathcal{H},S)+4c\sqrt{\frac{2\ln(4/\delta)}{m}}\end{split}$
where =$\ell\circ\mathcal{H}=\{\ell(h(x),y)\mid(x,y)\in\mathcal{X}\times\mathcal{Y},h\in\mathcal{H}\}$.
Lemma 6 (Lemma 2 of [68]).
Let $\mathcal{F}$ be class of real functions and $\mathcal{H}=[\mathcal{F}_{j}]^{k}_{j=1}$ be a k-dimensional function class. If $\mathcal{A}:\mathbb{R}^{k}\rightarrow\mathbb{R}$ is a Lipschitz function with constant $L$ and satisfies $\mathcal{A}(0)=0$, then
^$\hat{\mathcal{R}}_{m}(\mathcal{A}\circ\mathcal{H})\leq 2kL\hat{\mathcal{R}}_{m}(\mathcal{F}).$
Lemma 7 (Theorem 1 of [70]).
Let $\psi:\mathbb{R}^{M}\mapsto\mathbb{R}$ be a strictly convex differentiable function, and $D_{\psi}:\mathbb{R}^{M}\times\mathbb{R}^{M}\mapsto\mathbb{R}$ is the Bregman divergence induced by $\psi$. Let $Y$ be an arbitrary random variable taking values in $\mathbb{R}^{M}$ for which both $\mathbb{E}[Y]$ and $\mathbb{E}[\psi(Y)]$ are finite, we have
$\operatornamewithlimits{argmin}\limits_{s\in\mathbb{R}^{M}}\mathbb{E}_{Y}[D_{\psi}(Y,s)]=\mathbb{E}[Y]$
Acknowledgments
This should be a simple paragraph before the References to thank those individuals and institutions who have supported your work on this article.